A Complete Bibliography of Publications in IEEE Spectrum (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

11 December 2024 Version 1.38

Title word cross-reference

$ [Ulr20b]. $1 [Pot20b]. $100 [Sch22a, Sch22d]. 3 [And20f]. $74 [GA20]. > [Coo22, Gen22]. ^{2} [Cas23c].

000-Qubit [Cho23].

1 [Ano20-51, Koz23d, Ros21]. 1-Trillion-Transistor [LW24]. 100 [Cas22f]. 100th [Moo22c]. 118 [Smi20h]. 143-Year-Old [Zor22c]. 1800s [Ano20-78]. 19 [And20b, Cha20a, Gib21, Ham20a, Has20a, Hut20, NS20, Per20a, Pes20e, Pes21a, Pre20, SHD20, SZ20, Wal22, YS20]. 1935 [Bro22b]. 1950s [Hoc22]. 1970 [McC20].

2 [Sch21e]. 2.0 [Moo24b, WB21].

200-Year-Old [Smi21f]. 2020 [Ano20t]. 2021 [Ano21-115]. 20th [Moo20c]. 22 [Bro22b].

3D [BF24, Moo22a, Moo24d, SC23]. 3D-Chip [Moo24d].

5 [Ros20a]. 536 [Bre20]. 5G

[Hig20i, Koz21b].

6G’s [Pou22].

70-Year-Old [Cly24]. 75-Year-Old [Abb23].

= [BSAH21, Pal22].

A.A.C. [Smi20h]. Aakhilesh [Gen24d].

Aalyria [Koz23b]. Able [Smi21e]. Accelerating [Ulr21, Gar20]. Accelerator [Boy24]. Accelerators [Gen24l, Gar20]. Access [Smi22a, Gol22b, Koz21a, Ano21-57, Ano21-58, Ano22-59]. Accident [Har21, Ros22a]. Accordingly [KK21]. Accurately [Sla21]. Act [Ack22a, Moo24e, Pat23a]. Act-Funded [Pat23a]. actions [RS23]. Activity [Sla21]. Actually [SG24, Smi21k]. AD [Mar24k]. AD-X2 [Mar24k]. Adam [Der22a]. adapt [ALRW21]. adaptations [Gol24e]. add [Zel20]. adds [Gol22f]. ADHD [And20e]. Adieu [Gol24c]. Adjusted [Per22c]. Administers [KK21, Ano20h]. adopting [Ras20b]. Adrian [Ano21a]. Ads [Ano21-61]. Advanced [Kon24, Arm20]. Advances [And20c, MCG24, WAC^{+}20, Moo20b, Pes21a]. advancing [Nor20a]. Advantage [Jag22]. Advertisement [Ano20-103, Ano20-104, Ano21-27, Ano21-29, Ano21-30, Ano21-28, Ano21-31, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-63, Ano21-111, Ano21-57, Ano21-58, Ano21-128, Ano21-59, Ano21-137, Ano21-141, Ano21-145, Ano22-46, Ano22-47, Ano22-49, Ano22-48, Ano22-50, Ano22-51, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w]. Advertisement [Ano22x, Ano22y, Ano22z, Ano22-59, Ano22-131, Ano23b, Ano23c, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23-49, Ano23-50, Ano23-51, Ano23-52, Ano23-53, Ano23d, Ano23-54, Ano23-55, Ano23-56, Ano23-57, Ano23-58, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23-59, Ano24a, Ano24b, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano24-41, Ano24-42, Ano24-43, Ano24-44, Ano24-45, Ano24-46, Ano24-47, Ano24-48, Ano24-49, Ano24-50]. Advertisement [Ano24-51, Ano24-52, Ano24c, Ano24-53, Ano24-54, Ano24-55, Ano24-56, Ano24-57, Ano24-58, Ano24-59, Ano24-60, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano20p, Ano21-120, Ano21-144, Ano22-55]. Advertisements [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano21-32]. advocates [Gen23c]. Aerial [Mie24]. aerosol [Ano20h]. Aerospace [Gen24e, Pot24, JAR24]. Affair [Mar24k]. affects [Smi22d]. Affordable [WPBS23]. Africa [Smi22a]. African [Fai21, Moo20c]. After [Zor20, Zor22d, Hoc22, Ros20a, Smi20h, Zor23b]. Afterlife [WC23]. Aftermarket [Usk21]. Again [Gol24i, Sch21e, Str22b, Cas20i, DM20, Smi22j]. Against [Mar23j, Ros23b]. Age [Ack22a, Baj24, Bea22, Pes21b, Lee20]. ago [Pat24b, Tho21b]. agricultural [Fai21]. agricultures [Per20n]. Ahead [Ano21-55, Gol23d, Gol23h, And20g]. AI [Abb23, Ano20f, Ano21-113, Ano23a, BSG22, Baj22, Baj24, Bau24, BBS23, Bro23c, Bro21, BSAH21, Cal22, Cas20d, Chi21, CHM21, Fis24, Gen24a, Gen24h, Gen24j, Gen24k, GH23, Gol22a, Gol22b, Gol23a, Gol24h, Gol24f, Hsu20a, Koz23a, Koz24c, Lau24a, LW24, MMM24, Mar23k, MSS21, Moo22b, PN20, Per20f, SHP24a, Sch21a, Scu20b, Smi24, Str21, Str22a, SZ23, Str24a, VA23, Wal20a]. AI-assisted [Hsu20a]. AI-Generate [VA23]. AI-Generated [Koz23a]. AI-Powered [Lau24a]. Aibo [Ack24b]. aided [Hoc22]. aides [Bar20]. aids [Sch20b]. Aim [AZPC22, Smi23a]. Aims [Gen23e, Dum20]. Air [AZPC22, NG23, Smi21n, DM20, Gal21b, Smi20h]. Air-Quality [Smi21n]. Air-Taxi [AZPC22]. aircraft [Ros21]. airliner [JAR24]. Airliners [Sch21e, Smi21e]. airplanes [Cri20]. Airship [Koz23d]. AIs [Cho21, Usk21]. alchitry [Cas20f]. Alerts [Fit21]. Alexa [Has20b]. Algebra [Smi22c]. Algorithm [Cal24, Wal22]. Algorithmic [Lau24b]. Algorithms [RZ21, Per20e, Per21b, Str20b]. Alien [BMRW20]. All-Digital [Bow21]. All-Electric [Ulr21]. All-Metal [Cha21a]. All-Purpose [Chi21]. All-Seeing [Baj24, GAG20]. All-Sky [Sch23c]. Allaire [Cas22d]. allowed [Tho21b]. allows [Bra21]. Almost [Jon24a, Per21c, Svr23, Sch21b]. Alone [MMM24, Rae20]. along [Pat22]. already [Pat24a]. Also [Pat23b, AFH22]. Alt [Shi22]. Alt-Tech [Shi22]. Altarock [Gal20a]. alternative [MMP22, Moo21]. Alternatives [Der22a]. Alto [Bro23a]. Always [Pot24, SM24]. Am [Bro22b]. amazing [Cas22e]. Amazon [Has20b, Jon23b]. ambition [Zor22a]. AMD [Moo23a]. America [Ano20t]. American [Moo20c]. Amethyst [Sar20]. Ammonia [Gal21a]. amounts [Jaf22]. Amp [Pau22]. Amplifiers [Shi22]. Analog [BSAH21, CAS24b, Cas20b]. Andrew [Jon23a, Str22a]. angel [Per20k]. Anguilla [Koz24c]. Angus [Der22g]. Animatronic [Mar24e]. Ankur [Der21h]. anniversary [Moo22c]. Annual [Cas21d, Per22c]. anodes [And20c]. Anonymous [Hig20a]. Antenna [Cha21a, Sch22b]. antennas [Cas22b]. Anthrax [Mar21h]. antibody [And20b]. Antigenic [Wal22]. antiracist [Hsu20a]. antiviral [Ano20h]. antivirals [Scu20b]. Anton [Koz23a]. any [And20f, RK22, Svr23]. Anyone [Cha23]. Anything [Lev21]. Anytime [Bro21]. anywhere [YS20]. Aperture [Van23]. Apocalypse [SZ23]. Apollo [Cas22g]. Apollo-Era [Cas22g]. App [Fai21, BBS23, Per20l]. app-controlled [Per20l]. appear [And20c, WB21]. Appetite [Gol24h]. Apple [Smi22h, Per21a]. Applying [Ano20f]. approach [Cly22]. apps [Hsu20b]. AR/VR [Per20j]. Arc [Mar21i]. architect [Der20]. Architecture [Kas23]. Arctic [Rab22]. Arduino [Sch23b]. Arduino-based [Sch23b]. Argotec [BD22]. Argumentative [Ano23a]. Arjun [Goo23a]. Arm [Ack24c, Ano23-133, FBMI23]. Arms [You22, Ros20a]. Army [Ack21b]. Arpanet [Cro20]. Arrays [BF24]. Art [Fra21, Jon22d, Koz23a, SM24, McC20]. Arthritis [Dol23a]. Arthur [Gen24e]. Artificial [Gol23a, KK21, Lau24a, Zim21, Fun20, Hut21]. Artist [Gol23f, Gol24g, WB21]. Asad [Per22a]. Ascendant [Smi21j]. Ashore [Fai24a]. Asia [Fru20]. aspect [Smi22d]. Assisted [Sch24b, Hsu20a]. Assistive [Jon23a, Gol22e]. Asteroid [GO22]. astray [Hut20]. Astronaut [Ste21]. Astronomy [Sch24b]. athletes [Pat22]. Atom [Rad20]. Atom-thick [Rad20]. Atomic [CFPP22, Gen24c]. Atomically [WB21]. Attack [Ano20g]. attain [And20i]. Attorneys [Gol24i]. attracting [Per20f]. Audio [Gan24, LDO22, Pau22, Sch20b]. Augmented [Ros22a, HVW20, Per21a]. augmented-reality [HVW20, Per21a].

August [GO22]. Australia [Fai23b]. Authenticators [TB21]. author [Ano20p]. autism [Bar20]. autistic [HVW20]. Autodidacts [Jon23c]. Autoimmune [Dol23a]. Automakers [Cha23]. Automate [Lau24a]. Automated [Aro24, Zor23c]. Automatic [Mar24f, Ros22c]. Automating [Scu20b]. Automation [Coo21, Mar23k, Has20b, Scu20b]. Automaton [Ano20h, Fru20]. Autonomous [Ack22c, Ack23c, Ano20i, Der22g, LG20, Ros23a, Smi23a, AG20, And20a, Ano20-90, Der22h, Per20l, Wea20]. Autonomy [Jon24c, And20g]. AV [And20g]. available [Pou22]. Avatar [Ack23a]. Avicena [Moo23b]. away [Per22a]. AWS [Ano20-95]. Axial [GP22]. Axial-Flux [GP22]. AxiDraw [Cas20g]. Ayar [Moo23b].

B [Der21g]. Back [Ano20g, Ano20j, Ano20k, Ano20r, Ano20q, Ano20-69, Ano20-47, Ano20-48, Ano20-81, Ano20-92, Ano20s, Ano20-80, Ano20-102, Ano21-49, Ano21-50, Ano21-47, Ano21-48, Ano21-62, Ano22-46, Ano22-47, Ano22-50, Ano22-122, Ano22-132, Cas23c, Gol23d, Jon24a, LDO22, Luc20a, Rab22, WC23, Ano20-74, Ano20-75, Ano20-85, Red20, Ros20a, Sch21b, Ano20l, Ano20m, Ano20n, Ano20o, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano22-52, Ano22-53, Ano22-54]. Backbone [Tow21]. backers [AS22]. backpack [Ham20d]. backup [Sch20c]. Bad [Cho21, Smi21l]. baggage [GAG20]. Baidu [Smi23a]. Bajcsy [Ack24a]. Balance [AFH22]. Balloons [Sch22b]. ban [SG22]. Bandgap [Mis23]. bands [RHBM22]. Bandwidth [BD22, FBMI23, Pou22]. Bang [Sch21b]. Bangladesh [Fai20a]. Banish [Poo24b]. Bank [Pau22]. bare [Cas20h]. bargain [Ano20-45]. Barometer [Sch22a]. barrier [Moo20e, Pan20]. Based [Sch23a, Tei23, Ack23d, Cas23d, Cas24d, Gol24a, Moo23b, Sch23b, Zor20]. Basic [Gen24c]. Batteries [Cha24, Smi21e, And20c, Ano20i, Cri20, Fai21, Kum21, MW20, Moo21, Pat20, Ulr20b]. Battery [Bow21, GH23, Kum21, Mur23, Kum20]. Battle [Cas20b]. Battlefield [Ack21b]. battling [YS20]. Be [Cas22c, Pot24, Ack23d, And20b, Ano20-68, BS22, Cas22g, Cas23d, Cas24e, Cho21, CHM21, DM20, FBMI23, GOL24d, Gui20, Hsu20a, Hsu20b, MMM24, Moo22c, Pat22, Pot20b, SG24, Sch21d, Sch22e, Smi21d, Smi22l, Smi23c, VH20, WPBS23]. Beam [Smi21d]. beamed [Jaf22]. Beaming [Bar24]. Beams [Mos22]. Beat [Ack23c, Cha24]. Beatles [Arm20]. Became [Deg24, Ano20-68]. Because [Ano21-27, Ano22-55]. Become [Ano20p, Ano21a, MS22, Ros22c]. Becomes [SH22b]. Becoming [TGLM21]. Bee [Jon24c]. Been [Smi21m, Der23c, GOL24d, Koz24c, Per21b, Smi22g]. Before [Gen22, Jon23k, Mul24, NG23, Pes21a, SG24, Sch21b, BMRW20, Dut22]. began [Tem20]. begin [Koz23d]. beginning [Smi22d]. Begun [Gol24i, Usk21, Ulr20c]. Behemoths [Fai23a]. Behind [AZPC22, Koz23a, She24, HP20, Hut20]. Behold [Mar24a]. Being [GEN24g, Pes20a, SM24]. Bekeryte [GEN24g]. Bell [Der21e]. bellwether [Ano20s]. Below [CPBZ21]. bend [ZL20b]. bendable [Som21]. bespoke [Dum20]. Best [Ack23c, Ano21-144, Ano22-131, Ano22-132, Der21g, GP22, Sto23, Wat24, Ano20-103, Ano20-104, Ano21-145, Pot20b, Wal20a, Rao24]. Bet [Ano21-53, Koz24d, MCG24]. bets [Rak24a, Sch21a, Ulr20b]. Better [Baj22, Bro23c, BSAH21, Gol22a, Ano21-91, Arm20]. betting [Scu20b]. Between [Gol24a, Ano20-79]. Beware [Bro22d, Ros22b]. Beyond [Ano21-124, CAS24b, PSD22, BS22, Ros20a]. bias [Str20b]. biased [Hut21]. Bids [Gol24c]. Big [Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-51, Ano21-52, Ano21-114, Ano21-136, Ano22-56, Ano22-57, Ano22-58, Cly20, Der22a, Koz21a, Koz24d, LH24, MCG24, McC20, Poo24b, Ras20b, Sch24c, Str22a, Ano20-71, BS22, DM20, Ulr20b, Sch21b]. Biggest [Hub24]. Bike [Sch21d]. bikes [MMP22]. Billion [Per24b, Ulr20b]. Billion-Dollar [Per24b]. Billions [Fis24]. Biofuels [MW20]. Biology [Ano20q]. Bionic [Str22b, SH22b, You22]. Bionic-Hand [You22]. Biosensors [Per20a]. Birding [Sch24a]. Birds [Sch24a, Smi20c]. Bit [Coo21]. Bitcoin [Per22b]. Bits [Cas22g, Sar20]. Black [Mar21d, Per20b, Poo24b, Bou20, Moo20c]. BlackFly [Zor23c]. Blazed [Gol24c]. Blessin [Sac20]. blimp [Rae20]. Blind [Jon23a, SH22b, Tho21b]. Bling [Mar24j]. Bloat [Hub24]. blockchain [Per20j, Sto20]. blocks [And20b]. Blood [KK21, Wal24b]. Blue [Moo23b]. board [Ack21a, Lew20, Per21a, RG20]. Boat [Smi21d]. Bob [Gol24f]. Body [Smi21e, SH22b, Ano20r, SMD20, Som21]. body-conforming [Som21]. Boeing [JAR24]. Bold [RS23]. bolsters [CFPP22]. Bomb [Zor24a]. Bombs [Smi22i]. Bon [And20a]. Bonding [Moo24d]. Boom [Bro22c, CFPP22, Fis24, Jag22, LW24, Koz24c, Str21, Ros21]. boom-and-bust [Str21]. Boost [Bow21, Bro23c, Poo24a, Cha24, Fun20, Moo20b]. Boosters [Bro22a]. Boot [Ack21b, Koz24b]. boozy [Ano20-100]. boredom [Wal20b]. Borkowski [Gen24i]. Born [Bro22b, Mar20a, Mar24c]. Bosch [Gen24d]. Boston [Ack22b]. Both [Gan24, Gol24j, VA23, Hsu20b]. Box [Cas24f, Jon23h, Mar21d, Sch20d]. boxes [Ack22b]. Brain [AS22, Bou21, Cha21b, Dut22, Jon23c, Wal24b, Ack20, Dol23b, Gen24h, SG22, WAC^{+}20]. Brain-Controlled [Jon23c]. Brain-Implant [Wal24b]. brain-inspired [Gen24h]. Brains [Mar24d]. Brakes [Ros22c]. braking [Ros22c]. Brand [CAS24b]. Brass [Mar24d]. Break [Pau22]. Breaking [Pan20, Hec20]. breakout [Lew20]. breakthrough [And20b, Pri21]. Breeder [Pat23b]. Brew [Szk21]. Brewed [Ack23d, Cas23d]. Bricked [Pes21b]. Brief [Ack24b]. Briefcase [Gen24c]. Brightest [NYI24]. Brij [Der23a]. brilliance [Zor22a]. Bring [WC23]. Bringing [Red20, Lee20]. Brings [Cas24d, Gen23b]. briny [Pat24a]. Brittle [Cho21]. broadband [Hig20h]. broadcast [Ano20-46]. Brooklyn [Ano20-100]. Brutal [Cha21a]. Bugs [Sca22]. Build [BF24, Cas20e, GK24, Jon23c, Sch20c, Szk21, Cro20, Per22b]. Building [Bro22c, CHM21, Der22e, Pat23a, Smi20a, Wal24a]. buildings [AF23]. Builds [Der22g, Gen24e, Kum21]. Built [Cha21b, Coh22, Sch24c]. Buolamwini [Str24a]. buoyancy [Rae20]. Buoyant [Fai23a]. Burandt [Gol24i]. Buried [CPBZ21, Sch22d]. Bursts [Sch22a]. Buses [Ano22-56]. business [Pes20d, Rak24a]. bust [Str21]. Butler [GK24]. Buy [Zor23c, GA20].

C [Cas23c]. C64s [Ake23]. cab [Ulr20a]. cable [Koz21a]. calculator [Cas20h]. California [Ano20-101]. Called [GO22]. Calling [MZR21, Mar20b]. Calls [Smi22c, Ano20-79]. Calories [Mar21b, Sla21]. Camera [Sch23c, PK20]. camera-town [PK20]. Cameras [Gob23, Per21a]. Camp [Ack21b]. Campaign [SG22]. Can [Ack22a, Ack23b, AF23, BVKR23, Bau24, Bra21, Car24, Cas22g, Cas23c, Cha21a, Cha24, Cly23, DM20, FBMI23, FK21, Gol22a, Hsu20a, Hsu20b, Lau24a, MZR21, NG23, ROK24, Sch21d, Sch23a, Sha23, Smi21d, Szk21, Wal22, Zor23c, Ack22b, AFH22, BF24, BMRW20, Cas20a, Cas20f, Cas24e, Cho21, Der20, FOSMD23, Fra21, Gib21, Gol22b, GA20, Hec20, LA22, PN20, RHBM22, Sch21a, Scu20b, SHD20, Svr23, Wal20a, ZL20b]. canceling [Ros20b]. cancellation [Bra21]. cancer [Arm20]. candidate [Per20f]. can’t [RK22, Smi22k, Wal20a]. Capable [LH24, GA20, Sar20]. capacity [Fai23b]. Capture [Pat24a, Sch23c, Ano20-100, Gal21b]. car [Ano20-69, GA20, Per20k, Smi23a]. Carbide [MZR21]. Carbon [Gal21b, Pat24a, Smi21j, Smi22b, Smi22c, Ano20-100, Gal21a]. carbon-capture [Ano20-100]. Carbon-Sucking [Gal21b]. card [Per22b]. care [Str20b]. Career [Gen24a]. Careers [Ano20-95, Ano21a, Ano21-60, Ash23, Bar20, Cas20i, Der20, Der21a, Der21b, Der21c, Der21d, Der22a, Der22g, Der23a, Der23c, Der23b, Der23d, GEN24g, Gen23a, Gen23b, Gen23c, Gen23d, Gen24e, Gen24d, Gen24f, Gen24h, Gen24i, Gen24j, Gen24k, Gol23f, Per20j, Per20m, Per20i, Per24a, Sac20, SG24, Zor23a]. cargo [DM20, Gal21a, Pat24b, Ros22d]. carry [Hec20]. Cars [Ack21a, Gen24d, Mar23k, WPBS23, And20g, Hsu20c, Pan20, RHBM22, Ulr20c]. Carves [Cas22d]. Case [Abb23, Ano21-140, Ano21-141, BSG22, Gol24i, BCM^{+}19]. Castles [Bar24]. Cat [Cas24f, Per21b]. Cate [Koz24c]. Catfish [Mar21a]. cathode [Cas20e]. cathode-ray [Cas20e]. cathodes [And20c]. CDs [Nor20a]. Celebrate [Gol22d, SM24]. Celebration [Ano21-142]. cell [WAC^{+}20]. Cellphones [Lau23b]. cells [Gal21a, Ulr20b, WAC^{+}20]. Center [HH24, Jon23f, Gol22e, ZL20a, Gen24i]. Central [Gol21]. Centric [MCG24]. Century [Mar23j, Lee20, LDO22, Moo20c, Tho21b]. century-old [Lee20, LDO22]. CEO [Koz23b, Zor23a]. CEOs [NS20]. Cerf [Per23]. CERN [Gen24l]. Certifiably [Pes20b]. chain [RG20]. Chakraverty [Gen23d]. Challenge [Ack22c, Smi22d]. Challengers [Smi24]. Challenges [Gol23h, NG23]. champion [Cas20b]. Change [Bro22c, FK21, Tei23, MW20]. Changed [Gol22c, NS20]. Changing [SC23, Ano20-45]. channel [SMD20]. characteristic [Smi22j]. Characterization [Ano21-138, Ano21-137]. characters [Cas20a]. Charge [Ano22-56]. charging [Ulr20a]. Chasing [Sch22b]. Chasm [Gol24a]. Chatbot [Gol23g]. Cheap [Sch22b, Sch24a, Sch24b]. Cheaper [MD23, Ros23b]. Checkered [Eis23]. Chenming [Per20c]. Chess [Gol22b, Mar23a]. chicken [Smi20g]. Chief [CD21, Zor24b]. Chile [SG22]. China [And20i, Pat22, Pat23b, Sto20]. Chinese [Mul24, Smi23a]. Chip [Aro24, Gen24b, Gol24h, Kim22, LA22, LW24, MMM24, Moo24d, Moo24a, WPBS23, Moo20e]. Chiplets [Moo20b, Moo23a]. chipmaker [Rak24a]. CHIPS [Moo24e, Gol24b, Moo22b, Moo24c, Ano20-87, Cho20, Sar20, Pat23a]. Choong [Kim22]. Choong-Ki [Kim22]. Chores [GK24]. Chris [Koz23b]. Christian [Bar20]. Christiansen [Zor24b]. Chronicles [Gol23f]. Chutzpah [SM24]. Circa [Mar23a]. Circuit [Bow21, RG20]. circuitry [Cho20]. Circuits [BSAH21, Der22d, MZR21]. cities [Ong20, Sto20]. Citizen [Gol23e]. city [MMP22]. civilization [Gol22c]. claimed [Der22f, Pot20b]. Claiming [Smi20b]. Clapping [Mar24l]. Clark [Zor23a]. Clash [Koz21b]. clashed [McC20]. Class [Jon22f, Zor23c]. classic [Cas20h]. Classical [MMM24, MCG24, CD21]. Classical-Quantum [MCG24]. Classified [Ano21-61]. Cleaning [GEN24g]. clear [Smi22d]. Clever [Ash23, Hoc22]. click [Cas24e]. clickbait [BSG22]. Clients [Per24a]. Climate [Bro22c, Fis24, FK21, Tei23, Pat22]. climate-friendly [Pat22]. clinics [Der20]. Clippers [Bro22b]. Clock [Gen24c, Wat24, Cas20e]. close [Koz23b]. cloud [Pes21a, RS23]. Clueless [Cha23]. Clustering [Cas21c]. CMOS [Moo24b, PK20]. Coal [JJS22]. Cob [Gol24i]. cockpit [Mar20b]. cockpits [Der20]. code [Cas20a]. Coding [VA23]. Cognitive [BMRW20]. coil [Sch20b]. Cold [Abb23, Zor22e, BCM^{+}19, Arm20]. Coldness [AF23]. collaborated [McC20]. collaborating [JAR24]. Collaboration [Gol24g]. collaborative [Cro20]. Collapse [Smi22m]. colleagues [Per20b]. collect [Sch22c]. Color [De 22, HF22, SC23]. Colossus [Mar20j]. combat [Str24b]. Combating [Ano21-62]. Combination [MMM24, SZ20]. Combined [Bro23a]. Combines [GP22]. combustible [Zor22a]. come [LSM23, Pal22]. Comes [Dum23, Moo22b, Wat24, Zor22e, And20e, Ano21-113, Per20l]. Coming [GOL24d, Moo24c, Smi24, Tow21, Lau24b, Sch20a]. command [Ano20-50]. command-print [Ano20-50]. commercial [Ros21]. commercialize [Cly20]. Commercializes [Der22c]. Commodordion [Ake23]. common [Koz24d]. Comms [Cas23c]. communicate [Ong20]. Communication [Per24a, Ano20-79, Mit20]. Communications [Ano21-115, Zor22b]. Companies [WC23, Ack24d, Cly20, Per20a, Per20f, Ras20b]. Company [Cho23, Koz23a, Lau24a, Smi21l, Wal24b, Has20b]. comparable [SB22a]. Compelling [Gen23d]. Competitors [Rak24a]. Compiling [CG23]. complacency [Pes20c]. complaint [Koz24d]. complement [MMP22]. completely [Pan20]. complexity [Gol22f]. Complicated [You22]. components [Koz21b]. Compression [Per21b]. Compute [Gol24h, Moo24c]. Computer [CAS24b, Cal22, Kas23, Mar23a, Mar24j, Sch23b, Sch24c, Per20o, Sar20, VH20]. Computer-vision [Cal22]. Computers [HH24, AMK^{+}20, BS22, Gen24h, VH20]. Computing [Bro23a, CD21, Moo22b, Moo23b, Ham21, Pes21a, Ras20b, Sch21a]. COMSOL [Ano20t, Ano21-29, Ano21-30, Ano21-28, Ano21-63, Ano23-59]. Concept [Smi22l]. Concerning [CD21]. Conclusion [Mar23k]. Concrete [Ano20u, Fis24]. conditioning [Smi20h]. Conditions [Rab22]. Conference [Ano21-115, Ano20t]. confidential [Ras20b]. Configurations [ROK24]. Confinement [BBFS22]. conforming [Som21]. Congress [Fai23c]. Conjurer [Per21b]. connected [BMRW20, Ong20]. connected-vehicle [Ong20]. Connection [Moo24d]. Connections [Gol24f]. Connnected [FOSMD23]. Conquer [Ack22c]. Cons [SG24]. consensus [Cro20]. Consent [WC23, Hig20c]. Consequential [Gol23f]. conserve [Gar20]. consortium [NG23]. Constructing [VH20]. Consumer [Gob23, Per20g, Der22f]. Consumers [Cha23]. contact [Hsu20b, Pes21a]. contact-tracing [Hsu20b]. contacts [Cas22e]. Content [Str24b]. Contents [Ano20v, Ano20w, Ano20x, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano21-64, Ano21-75, Ano21-65, Ano21-66, Ano21-67, Ano21-68, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano21-73, Ano21-74, Ano22-60, Ano22-68, Ano22-69, Ano22-70, Ano22-71, Ano22-61, Ano22-62, Ano22-63, Ano22-64, Ano22-65, Ano22-66, Ano22-67, Ano23-60, Ano23-65, Ano23-66, Ano23-67, Ano23-68, Ano23-69, Ano23-70, Ano23-61, Ano23-62, Ano23-63, Ano23-64, Ano24-61, Ano24-67, Ano24-68, Ano24-69, Ano24-70, Ano24-71, Ano24-72, Ano24-62, Ano24-63, Ano24-64, Ano24-65, Ano24-66, Ano20-98, Ano21-113]. Continent [Smi22a]. Contrera [Gen24j]. Contributions [Gol24f]. Contributors [Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano21-76, Ano21-87, Ano21-77, Ano21-78, Ano21-79, Ano21-80, Ano21-81, Ano21-82, Ano21-83, Ano21-84, Ano21-85, Ano21-86, Ano22-72, Ano22-80, Ano22-81, Ano22-82, Ano22-83, Ano22-73, Ano22-74, Ano22-75, Ano22-76, Ano22-77, Ano22-78, Ano22-79, Ano23-76, Ano23-77, Ano23-78, Ano23-79, Ano23-80, Ano23-81, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75, Ano24-79, Ano24-80, Ano24-81, Ano24-82, Ano24-83, Ano24-73, Ano24-84, Ano24-74, Ano24-75, Ano24-76, Ano24-77, Ano24-78]. Control [Ano21-138, Ano21-137, Bor24, FBMI23, SHD20]. Controlled [Jon23c, Per20l]. controversial [Sch22c]. conventional [Ong20]. Conversation [Gol23g]. Convolutional [Fra21]. Cooker [Mar24f]. Cool [Gen24i, AF23]. COP26 [Smi22c]. Copper [Moo24d]. Copy [Cha21b, Per20o]. Core [Cas22g]. Corners [RHBM22]. coronavirus [And20b, Ham20a, Per20e, Wal20a]. Corporate [HS23]. Correction [BS22]. Cost [Poo24b, TGLM21, Ver24]. Costliest [Smi22d]. Costly [You22]. costs [Gar20, Moo20b]. Could [BD22, Boy24, CP21, GK24, Moo24e, Pat23b, Poo24b, Smi21j, TB21, Wal21, WC23, AMK^{+}20, Cho20, CHM21, Gal20a, Gal21a, Gar20, Gui20, Jes21, LCBH20, Moo24c, Ong20, Pat20, Pat24b, Per21a, Raf20b, RS23, WB21]. Count [Sla21, Smi24, Hig20e]. Counterfeiters [TB21]. Counting [Mar21b]. Countries [Smi21k]. Country [Kim22, Fai23b, JJS22]. Courts [Smi23c]. Cover [Ano21-88, Ano21-89, Ano21-90, Ano21-92, Ano21-93, Ano21-96, Ano21-103, Ano22-95, Ano22-103, Ano22-88, Ano22-96, Ano22-104, Ano22-89, Ano22-97, Ano22-105, Ano22-90, Ano22-98, Ano22-106, Ano22-91, Ano22-99, Ano22-107, Ano22-84, Ano22-92, Ano22-100, Ano22-85, Ano22-93, Ano22-101, Ano22-86, Ano22-94, Ano22-102, Ano22-87, Ano22-112, Ano22-113, Ano22-114, Ano22-115, Ano22-116, Ano22-117, Ano22-118, Ano22-119, Ano22-120, Ano23-82, Ano23-96, Ano23-105, Ano23-88, Ano23-97, Ano23-106, Ano23-89, Ano23-98, Ano23-107, Ano23-90, Ano23-99, Ano23-108, Ano23-91, Ano23-100, Ano23-83, Ano23-92, Ano23-101, Ano23-84, Ano23-93, Ano23-102, Ano23-85, Ano23-94, Ano23-103, Ano23-86, Ano23-95, Ano23-104, Ano23-87, Ano23-112, Ano23-113, Ano23-114, Ano23-115, Ano23-116, Ano23-117, Ano23-118, Ano23-119, Ano23-120, Ano23-121, Ano23-122, Ano24-100, Ano24-111, Ano24-90, Ano24-101, Ano24-112, Ano24-91, Ano24-102, Ano24-113, Ano24-92, Ano24-103, Ano24-114]. Cover [Ano24-93, Ano24-104, Ano24-115, Ano24-94, Ano24-105, Ano24-116, Ano24-95, Ano24-106, Ano24-117, Ano24-85, Ano24-96, Ano24-107, Ano24-86, Ano24-97, Ano24-108, Ano24-87, Ano24-98, Ano24-109, Ano24-88, Ano24-99, Ano24-110, Ano24-89, Ano24-119, Ano24-120, Ano24-121, Ano24-122, Ano24-123, Ano24-124, Ano24-125, Ano24-126, Ano24-127, Ano24-128, Ano24-129, Ano24-130, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20-52, Ano20-53, Ano20-54, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano20-60, Ano20-61, Ano20-62, Ano20-63, Ano20-64, Ano20-65, Ano20-74, Ano20-75, Ano20-76, Ano20-77, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-94, Ano21-95, Ano21-97, Ano21-98, Ano21-99, Ano21-100, Ano21-101, Ano21-102, Ano21-104, Ano21-111, Ano21-105, Ano21-106, Ano21-107, Ano21-108, Ano21-109]. cover [Ano21-110, Ano22-46, Ano22-47, Ano22-49, Ano22-48, Ano22-50, Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano22-109, Ano22-110, Ano22-111, Ano22-121, Ano22-122, Ano22-132]. Cover2 [Ano23-109, Ano24-118]. Cover3 [Ano23-110]. Cover4 [Ano23-111]. Coverage [BD22]. COVID [And20b, Cha20a, Gib21, Ham20a, Has20a, Hut20, NS20, Per20a, Pes20e, Pes21a, Pre20, SB22a, SZ20, Wal22, YS20, SHD20, Str22b]. COVID-19 [And20b, Cha20a, Gib21, Ham20a, Has20a, Hut20, NS20, Per20a, Pes20e, Pre20, SZ20, Wal22, YS20, SHD20]. CPU [Kas23, Reg23, Szk21]. Cracked [Abb23]. crafted [LCBH20]. Craig [Der22b]. Cranes [Smi20c]. cranks [Ulr20b]. crashes [Per22a]. cravings [Dol23b]. Craze [AZPC22]. creaky [Cha20b]. Cream [Gen24i]. Create [Der21e, Gol22d, Mit20]. Created [Gen22]. Creating [KK21, Sac20, Zor24a]. Creation [Koz23a]. Creativity [Bro23c, SM24, Gol22b]. Creator [Koz23c]. Creators [Per24b]. Credentials [Str24b]. credit [Per22b]. credit-card [Per22b]. Creepy [WC23]. crewed [Jon24a]. crisis [Cha20a]. Critical [VA23]. Critters [Ros22e]. Crop [Gen24e]. Crop-Spraying [Gen24e]. CrossTalk [Ano20u, Hig20a, Hig20c, Hig20g, Pes20d, Pes20f, Smi20c, Smi20e, Smi20f, Smi20h, Ano20-45, Ano20-94, Hig20j, Ano20-44, Ano20-88, HP20, Hig21, Smi20a, Smi20b, Smi20d, Smi20g, Smi21a]. Crowdfunded [Ste21]. Crown [Smi24]. crucial [Hsu20b]. Cruise [Ros23b, Per20l, Wea20]. Cruz [Gen24f]. Crypto [Ano21-53]. Cryptocurrency [Gen23e, Per22b]. cryptographic [Lew20]. Cryptography [Rao24]. Cryptokitties [Smi22m]. Crystal [Mar21e]. Crystals [NYI24]. CubeSats [Mit20]. Cummings [Mar23k, Ros23a]. Cup [Sto23]. Curbside [Ulr20a]. Currencies [HS23]. Current [BSAH21]. cut [Moo20b, Per20h, Per20o, Scu20b]. Cuts [Smi22c]. cybersecurity [Der23c]. cycle [Str21].

D [And20f]. Dangerous [VA23]. Dare [SHP24a]. dared [Ack23d, Cas23d]. Dark [Ano21-54, Moo23b, Sch20c]. DARPA [Ack22c]. dashboard [Sch23b]. Data [Cal24, Car24, Dum20, Har22, HH24, Hig20b, Koz21a, LH24, Moo24c, Rab22, AS22, Nor20a, Per21b, RS23, Ras20b, SMD20, SG22, ZL20a]. Data-driven [Hig20b]. databases [Dum20]. David [Gol23e, Koz24a]. Davis [Ano20-101]. Day [Aro24, Bro22b, Tow21, Gen24k]. Days [Ano21-55]. DC [Fai24a]. dead [Red20]. death [Der20, SB22a]. debut [Per20a]. Decade [Kon24]. Decades [Bro22a, HF22, Der23c, Gol23d, Pat24b, Per21b, Per23, Ros20a, Zor20, Smi23b]. Decarbonization [Smi22c, Smi22d]. Decarbonize [Cab24, Gol24e, MMP22]. December [Ano21-115]. Decent [BD22]. Decide [SG24]. decision [Hut21]. decision-making [Hut21]. Declining [Smi21k]. Deco [Jon22d]. Decoupling [RS23]. Dedicated [Tow21]. Deep [Ack21b, Ano21-113, Ham21, LCBH20, LDO22, MSS21, Sch21a, Sch22c, TGLM21, BMRW20, Jon24a, Moo20e, Pat24a, Per20n, Zor22d, Dum20]. deep-learning [Per20n]. deep-ocean [Zor22d]. Deep-Sea [Sch22c]. deep-space [BMRW20, Jon24a]. Deep-speare [LCBH20]. Deep-time [Dum20]. deeper [Ano20-84, Gal20a]. deepfakes [BSG22, Str20a, Str24b]. DeepMind [Chi21]. Deere [Der23a, Per20n]. Defender [Ulr22]. Defending [Cla24, Gol24b]. Defense [Cha21b, Lew23, Per22a]. Defined [Kas23, Cas22b, Lee20]. defunct [Red20]. Degl’Innocenti [Gen24l]. Delight [Cas23b]. Deliver [BD22, Per20e, Zor20]. delivered [Ano20-68, Wal24b]. delivering [LDO22]. Delivers [Gol24j]. Delivery [Cal24, Jon23b, Mar20h, Rak24a, Sch20a]. Dell [Smi22l]. Demand [AFH22, Str22b, Per20j]. Demands [LW24]. demonstrate [Cly22]. Demos [PRR23]. Dennisa [Der21d]. density [Cha24, Smi22k]. Departments [Per20d]. Depend [Fit21, Per20h]. describe [NS20]. Desert [Jon22e]. deserve [Hig20h]. Design [Ano21-91, Cha21a, Cly24, Kas23, MMM24, Gol22e, Hig20a]. designed [Per20l]. Designer [Gen23d, Zor24a]. Designing [Sto23]. Designs [GP22, Smi22e]. Despite [Fun20, Rab22]. destroy [And20h]. Destroyed [Mar23j, Moo20c]. destructiveness [Smi22i]. Detailed [Sch24b]. Detect [Sch22a, Sch23a, Sch24a, WB21]. Detecting [Mar23j]. Detection [Baj22]. Detector [Cas24f, Jon23i, Kuh24, Wal20b]. Deterrence [Baj22]. develop [Per20e]. developers [Per20j, Wea20]. Development [Ash23, Hsu20a, Zor20]. Developments [JPWB21, MJH^{+}21, SAS^{+}21, CKG^{+}21]. Device [Gen24m, Gol22c, KK21, Moo22c]. devices [Ano21-91, Arm20, MW20]. diagnostic [YS20]. Dialogue [CD21]. Did [Bro23c]. Didn’t [Bro23c, Cas20g]. difference [Pat24b]. Different [Dol23b, Ack23d, Cas23d]. dig [Gal20a]. Digital [Bow21, CAS24b, Der23d, Dum20, HS23, Smi23c, WC23, Koz23b, Mar20a, Str24b, Zel20]. dilemma [Hsu20b, Bro22d]. dimension [Per20c]. Diminishing [TGLM21]. dioxide [Ano20-100]. direct [Gal21b]. direct-air [Gal21b]. direction [RK22]. disabilities [Sac20]. Disability [Ano21-121]. Disappear [Poo24b, Per21b]. disappeared [Fru20]. disaster [Cha20a]. Disastrously [Cho21]. Discounts [Ano21-117, Ano21-118, Ano21-119, Ano20-71, Ano21-120, Ano22-123]. discovery [Scu20b]. discrimination [Per20i]. Dish [Sch24a]. disinfected [Ano20-68]. disinformation [Str24b]. Disks [Ake23]. Disney [Gen23b]. Disorders [Dol23a]. Disparate [Der21e]. Display [Mie24, Som21, ZL20b, Cas20e, Cas22f, Mie22]. Displays [Gan24, Pal22, Per21a]. Disrupt [FK21]. disruptors [Bro22d]. disrupts [Ham20a]. Distance [Jaf22]. Distanced [Ano21-142]. distances [Jaf22]. Distant [Sch23a]. distillery [Str23]. distinguish [LCBH20]. distribute [Gib21]. Dive [Cal22]. divide [Koz23b]. divided [Koz21b]. Divya [Der21g]. DIY [Bor24, Cas20d, Cas22a, Gol23e, Lew23, Pau22, Ros22b, Sch20b, Sch22d, Sla21, Tem20, Wat24]. DNA [Abb23, BF24, Car24]. DNS [Ras20a]. Do [GK24, PRR23, Pow24, SHP24a, SG24, Cas23a, Gal20b, Wal20a]. Do-or-Die [PRR23]. Docker [Cas21c]. Does [Pre20]. Doesn’t [Hus24]. dog [GA20]. Dollar [Per24b]. Domestic [GK24]. dominance [Ulr20c]. Don [Zor24b]. Done [BSAH21, Per20h]. Don’t [Gen24a, Smi21d, Smi21e, Smi21f, Smi21b, Smi21c, Smi21g, Smi21h, Smi21i, Smi21j, Smi21k, Smi22a, Smi22b, Smi22c, Smi22e, Smi22j, Smi24, BS22, Smi22i]. Doom [Ano21-62]. door [Lew20]. Doorstep [RZ21]. doses [Gib21]. dot [Scu20a]. Dots [Pal22, PK20]. Double [CP21, Moo22b, Sch20c]. double-backup [Sch20c]. Doubles [Gol24b]. Down [Gol24b, Gol24i, Som21, Rae20, Fai23b]. Downed [Sch22b]. Downs [Moo21]. Doyen [Gol24c]. drag [Per20m]. Dragon [Cha21b]. Dragonfly [Cha21b]. Dream [Hoc22, Smi21f]. Dressed [CP21]. drills [Gal20a]. drinking [Smi20e]. Drive [Str22b, Ack21a, Ano20-69]. driven [Hig20b]. driver [Har22, Wea20]. driverless [Per20l]. Drivers [Cal24]. Drives [Ros23b]. Driving [Mar23k, WPBS23, And20g, LG20, Pan20, Smi23a]. Drone [Tei23, DM20, Pot20a, Ros22d]. Drone-Based [Tei23]. Drones [Ack23c, Bau24, Cla24, Gen24e, Gol24b, Ros23b, DM20, Gen24h, Sch20a]. Dropout [Bow21]. Dropped [Cal24]. Drops [Per22c]. Drove [Moo23a]. drug [Scu20b]. Dual [Kas23]. Dual-CPU [Kas23]. Dug [Cal24]. Dumpster [Cal22]. Duo [SH22a]. Duolingo [BBS23]. during [Arm20, Cha20a, NS20]. Duty [Der23a]. DVDs [Nor20a]. Dynamic [SH22a]. Dynamics [Ack22b].

E-Bike [Sch21d]. E-Paper [HF22]. Each [Gol23g]. Ear [Lew23]. Early [PRR23, Koz20c]. Earn [Ano20-93]. ears [Hoc22]. Earth [AF23, Bar24, ROK24, BMRW20, Ano20-85, Ano24-145, Dum20, WB21, Zor22b]. Earth-like [WB21]. Earthquake [Mar23j]. Earthquakes [Mar23j, Sch23a]. Easier [Cas21c, Coo21, Gen24k]. Easy [Sch21d, Smi21n, Tho21a, Cas20d]. Eat [Smi22h]. Eating [Smi22h]. Eavesdropping [Dut22, Hoc22]. Echo [Has20b]. economic [Fun20]. Economics [Ros23b]. Economy [Sch23b]. ED [Wea20]. edge [Hig20d]. Editor [Zor24b]. Editorial [Zor23b]. Edwin [Zor22c]. EE [Der23b]. Effect [Zor22c]. effective [Hsu20b]. efficiency [Sch21a]. Efficient [Cha21b, PSD22]. eight [Ack24d]. Eighth [Cas21d]. Eileen [Der23b]. Either [Cas23b, Cha23]. Election [Str24b]. Electric [Ano22-56, MMP22, Mar21k, Mar22i, Smi21e, Smi21j, Ulr21, Ulr22, Wal21, And20c, Ano20r, Ano20-69, Cri20, Gol22f, JAR24, Ulr20b, Ulr20c]. electric-vehicle [Gol22f]. electrical [Ano20-86]. Electricity [AFH22, Ano20-45, Mar21i, Smi22a, Smi22j, Fai23b, Smi22g]. Electrified [GP22]. Electrifying [Der23a, Sch21d]. Electrifyingly [Sch21d]. electrodes [Dut22, Wal24b]. Electrolysis [MD23]. electrolytes [And20c]. Electromechanical [Pau21, TB21]. Electronic [CP21, Mar24a]. Electronically [Sch24b]. electronics [AMK^{+}20, Som21]. Electrons [Boy24]. Electrostatic [Mar23c]. Eliminates [BBFS22]. Elon [Smi21f]. Else [Lev21]. elusive [Cly22, Luc20a]. embed [Cas20d]. embedded [Ano20-90]. Embeds [Gen24m]. Embracing [Has20a]. emerged [Zor22a]. Emergency [Ano20-46, Ros22c]. Emerges [Moo23b]. Emissions [Smi22c, Gal21a, Gal21b]. emotional [CHM21]. emotions [HVW20]. EMP [Wal24a]. EMP-Proof [Wal24a]. employer [Ano20-95]. Empower [Ack23b]. emulate [BBS23]. Emulated [Reg23]. Enable [Gan24, NYI24]. Enabled [Pau22]. Enabling [Ong20]. Encrypted [Moo24c, Ras20a]. Encryption [Pau21, Koz20c, Moo24c]. End [Pes21a, Per20a, Smi22d]. ended [Ros20a]. Ending [Has20a]. Energiewende [Smi20d]. Energy [BSAH21, Cly24, Gar20, GP22, KH24, Moo21, SC23, Smi22b, Smi22f, Cha24, Fai23c, Ham21, Jaf22, MMP22, Smi22k, Cly22]. enforcer [Zor23a]. Engineer [Der21g, Gen24a, Gen24e, Gen24d, Gen24j, Gen24l, Gol23h, Kim22, Ano20-47, CD21, Gen23a]. Engineered [And20b, Cas20h]. Engineering [Ash23, Der21d, HPK22, JPWB21, MJH^{+}21, NS20, Per24b, SAS^{+}21, Ano20-86, Cas20b, CKG^{+}21, CFPP22, Koz24b]. Engineers [FK21, Gol22d, Per22c, Gen23c, NS20]. engines [Gal21a]. England [Jon23d]. Enhanced [Gol21]. Enhancing [Ack22a]. Enough [MMP22, Cha24, Gal21b, Gib21]. ensure [Der22f]. ensures [Der21d]. Entangled [Ros22e]. Enters [WPBS23]. enthralls [Smi23b]. entire [Sha23]. entrepreneur [Ano21a, Goo23a]. Entry [Dum23]. Environment [Cha21a]. Environmental [Cab24]. environments [BMRW20]. Envoy [Mar22k]. Equals [Ake23]. Equiano [Koz21a]. Equipment [Hec20]. equipped [Sch20b]. Era [Cas22g, Shi22, Smi22a, Wal21, Bro22b, Sch22e]. Erickson [Gen24e]. Erika [Gen24f]. Error [BS22]. Escape [Bor24, ROK24]. Essential [Cor21a, Ano20-68]. Eugene [Der23c]. Europa [Cha21a]. Europe [Dum23, Fai24a, Ros22c, Ros22d]. EUV [Boy24, Van23]. EV [And20c, Cha23, GEN24g, GH23, Gol23h, Smi23a]. evangelizing [Per23]. Even [Cas23c, Per20g, Sch23a, Szk21, BMRW20]. Event [Lau24b, Ano20t, Bou20]. Ever [NYI24, Dut22, Pri21]. Every [Har22, Smi22d]. Everything [Gol22c, Hig20b, Hig20d, Hig20e, Hig20f, Hig20h, Hig20i, HP21a, HP21b, HP21e, HP21f, HP21c, HP21d]. Everywhere [Gol23a, Per21c]. Evolution [Cha21b, Cas22e, Fun20]. Evolve [Wal22]. Evolved [Smi22h]. EVs [Cha24, Der22e, Dum22, Gol22f, MMP22, ROK24, Ulr20a]. eVTOL [AZPC22, Zor23a, Zor23c]. eVTOLs [ACDG21, NG23]. Exascale [Dum23, And20i, Sch22e]. exceed [Fai23b]. Exceptional [Pau21]. Excess [SB22a]. Exclusive [Ano20-71]. exoplanet [WB21]. exoskeleton [Ack20]. Exotic [ROK24]. Expect [Fun20]. expectations [And20g, Per20d]. Experience [Der22d, Gol21]. experiences [Der20]. Experimental [ROK24]. experiments [Cas20g, Ham20a]. Expert [Der23d, Mar23k]. Experts [Gal21b]. Exploding [Jag22]. Explorers [Der22g]. Explores [HPK22]. expressions [HVW20]. extension [Red20]. Extra [FBMI23]. Extreme [Gol24h, Smi22e]. Extremely [ACDG21]. Eye [Ano24-144, Zor22e, Sch20d]. Eyeballs [Gen23e]. Eyewear [Ano21-114]. EZ [Mie22].

Fab [Ano22-108, Aro24, Moo24e]. Fabs [Pat23a]. Facebook [HS23, Str20a]. faced [BMRW20]. Facial [Abb23, HVW20]. Factor [Gen23d]. facts [Ano20u]. faculty [Ano20-86]. fades [Pes21a]. Fail [Cho21, Cha20b, Raf20b]. failed [HS23, Pat24b]. Fair [McC20]. Fake [Fra21]. fallback [Raf20b]. False [Eis23]. famous [Bou20]. Fans [Gal21b]. fantasy [Ano20-51]. Far [Fai23a, Ano20-49, PSD22]. Farewell [Mar24e, Zor23b]. Farmers [Gen24e]. Farms [Pat24a, Usk21]. fascinating [Gol22d]. Fast [Cha21b, Cly24, WPBS23, YS20, Ano20-49, Ano20-79, Smi22j]. Faster [Tow21, Ano21-91, CHM21, Gal20a, Per20e, Ros21]. faster-than-sound [Ros21]. father [Per20c]. Feasibility [Gol24a]. Features [Gol23g]. Feedback [Gan24]. Feel [Ack22a]. Feels [CHM21]. Fell [Ros22a]. Ferrite [Cas22g]. Ferrite-Core [Cas22g]. fetch [GA20]. few [LCBH20, ZL20b]. fewer [Sca22]. Fi [Gol24j, Koz24d, Lew23, Mie24, Cas22a, Gol24a, Hig20j, Zel20]. fibers [Hec20]. Fiction [Der21g, Jon22a, Zim21]. fieldwork [Ham20a]. Fight [Gol24i, Ras20a]. Fighter [Ros23a]. fighting [Der23c]. Figure [Jon22a]. Figured [Pau21]. figures [Mar20j]. File [PSD22]. film [Zel20]. filter [Ros20b]. Final [Ano20-48]. finalist [Ano20-100]. Finally [Cha24, Gob23, GK24, Zor23c, Gol23d, Zor20, Zor22d]. Find [MD23, Pow24, Pri21]. Finding [Abb23, Zor20]. Finds [Gen23d]. FinFets [Per20c]. Finish [Gen24d]. Finished [Aro24]. Fires [Fit21]. First [Aro24, Gen24b, Koz21b, Mar22l, Mar24g, Per24b, Ros22a, Ste21, SG22, Zor22a, Zor24a, And20b, And20e, Fru20, Jon24a, Mar20c, Mar20d, Pot20a, Sch22e, Zor23c]. first-person [And20e]. Firsthand [Mar23k]. Fishing [Her20]. Fission [Jon23d, And20d]. Fit [Smi22l]. Fitbit [Per24b]. Fitness [Per24b]. fits [Mar20f]. five [Per23, Scu20b]. Fix [Smi21n, Str24a].

Fixed [Smi22l]. Flares [Sch22a]. Fleet [Der22e, Raf20b]. flesh [SMD20]. Flexible [Der21e]. flier [DM20]. fliers [Zor23c]. flies [Rae20]. Flight [Jon23j, Smi21e, Smi23b, Ano20-89, CD21, Fou22, Ros21]. Flights [Smi21e, Koz23d]. Flipper [Cas23b]. Floats [Mie24]. floor [Sch22c]. Floppy [Ake23]. flourished [Hoc22]. Flow [Cha24]. flu [SB22a]. fluctuations [Koz20c]. flurry [Pes21a]. Flux [GP22]. Fly [Cha21b, JAR24, Pot20a]. flyer [Pot20b]. Flying [Ano20-49, Ros20a, Ros22d, Zor23a]. focus [And20g]. focuses [Gol22d]. fold [ZL20b]. followed [Smi22j]. Following [GOL24d]. Food [Smi22f]. Foosball [Coh22]. Footprint [Smi22b]. foreshadows [Fai24a]. foresight [SZ20]. Forest [Fit21, PN20]. forestry [Ano20f]. Forgetful [Cho21]. Forgotten [Aro24, Mul24, Shi22, Smi22g]. fork [Per22a]. form [And20e]. Former [Ros23a, Zor24b, Gen23c]. Formerly [Pal22]. Forming [KH24]. Formula [Ano20-51]. Forward [Ano23-134, Bro23b, Cor21b, Goo23b, MMM24, Mar20a, Mar20b, Mar20c, Mar20f, Mar20g, Mar20i, Mar21d, Mar21e, Mar21f, Mar21h, Mar21k, Mar21i, Mar21j, Mar21c, Mar21g, Mar22i, Mar22j, Mar22l, Mar22k, Mar22a, Mar22b, Mar22c, Mar22d, Mar22e, Mar22f, Mar22g, Mar22h, Mar23b, Mar23d, Mar23e, Mar23f, Mar23g, Mar23h, Mar23i, Mar24b, Mar24f, Mar24d, Mar24g, Mar24e, Mar24h, Mar24c, Mar24i, Mar24j, Bre20]. Forword [Mar24k]. Fostered [Gol24g]. Fought [Cal24]. Found [SH22b]. Foundation [Ano20-70, Ano21-31]. Founder [Ack24a]. foundry [Rak24a]. FPGA [Cas20f]. Fractional [AMK^{+}20]. Fraden [Koz24b]. Frankel [Koz24a]. Free [Tho21a, Per20f]. freight [DM20]. frequency [Bra21]. Fried [Cas23a]. friendly [Fru20, Pat22]. Friends [Mar24e]. Front [Ano20-53, Ano20-54, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano21-92, Ano21-93, Ano21-96, Ano21-103, Ano22-49, Ano22-48, Ano22-51, Ano22-112, Ano22-113, Ano22-114, Ano22-115, Ano22-116, Ano22-117, Ano22-118, Ano22-119, Ano22-120, Ano23-112, Ano23-113, Ano23-114, Ano23-115, Ano23-116, Ano23-117, Ano23-118, Ano23-119, Ano23-120, Ano23-121, Ano23-122, Ano24-119, Ano24-120, Ano24-121, Ano24-122, Ano24-123, Ano24-124, Ano24-125, Ano24-126, Ano24-127, Ano24-128, Ano24-129, Ano24-130, Ano20-76, Ano20-77, Ano20-52, Ano20-60, Ano20-61, Ano20-62, Ano20-63, Ano20-64, Ano20-65, Ano21-94, Ano21-95, Ano21-97, Ano21-98, Ano21-99, Ano21-100, Ano21-101, Ano21-102, Ano21-104, Ano21-111, Ano21-105, Ano21-106, Ano21-107, Ano21-108, Ano21-109, Ano21-110, Ano22-109, Ano22-110, Ano22-111, Ano22-121]. frontier [Som21, Sch22e]. Frontiers [Kon24]. Fuel [Sch23b, Smi22f, Svr23, Gal21a, Smi22k]. Fuel-Economy [Sch23b]. fuelless [Ano20-49]. full [Ano20-70]. Fully [Moo24c, Per20l, Smi23a]. Fun [Gen23d, Mar20e]. Functional [Sca22]. Funded [Pat23a]. Fureka [Lau24a]. Fusion [BBFS22, Cly22, Cly24, Der21e, Gol23b, Jon22h, Kon24, BCM^{+}19, Cly20, Cly23]. Fuss [Kuh24]. Future [Ack22c, Ack23a, Ano21-122, Boy24, GOL24d, Ham21, Str21, Tei23, And20c, And20d, Ano20-72, Ano20-73, Ano22-130, Luc20a, MW20, Moo20b, Ulr20b].

Gadget [Wal21]. Gadgeteering

[Ano21-112]. gains [Per20i]. Gakutensoku [Fru20]. Galaksija [Ack23d, Cas23d]. Galaxy [Pow24]. Gallium [Jes21]. Game [Pes21a, Reg23, Wea20, And20e, Smi23b, ZL20a, Smi24]. Games [Gen23d]. gaming [And20e, Pes21a]. Gamma [Ros22b, Sch22a].

Gamma-Ray [Ros22b, Sch22a]. Ganguli [Der21h]. garage [Lew20]. Garwin [Zor24a]. Gas [Jon22c, Gal21b]. gateways [Gol23c]. gauge [Sch23b]. GE [Has20b, JAR24]. gear [And20c]. gears [Jon24a]. Geek [Cas20a, Cas20b]. Geely [Smi23a]. Gen [Pau22]. General [Der21h]. Generate [VA23]. Generated [Koz23a]. generating [Fai23b]. generation [Fai24a, Koz24d]. Generations [Sto23]. Generative [Bro23c, SHP24a]. Generator [Svr23]. Generators [Mar23c]. Geostationary [Koz20a]. geothermal [Gal20a]. Germany [Dum23]. Germicidal [And20h]. gestures [Wea20]. Get [HS23, BS22, Fai23c, Gol23a, Gol23d, Hut20, Per20f, RK22]. Gets [Dol23a, Rab22, Ulr22, Hig20i, Per20h]. Getting [Ash23, LDO22]. Ghost [Koz20c, Cas23a]. Giant [Zor22b, Fru20, Gal20b, HS23]. giants [Smi23a]. GIFs [Per21b]. Gig [Gen24i, Cal24]. Gill [Ack22a]. Give [Bro23c, FBMI23, KH24, Red20, RHBM22, WAC^{+}20]. Gives [Coo21]. Gizmo [SP21a, SP21b, SP21c, SB22b, SB22c, Smi22l]. Glass [Smi22b, Nor20a, HVW20]. Glasses [Per21a]. Gliders [Jon23e]. Glitch [Ros22b]. Global [Ano21-115, Fai23a, Fai24a, Raf20b, RG20]. Globalism [Bro22b]. Gloom [Ano21-62]. Glucose [KK21]. GM [Ulr20b]. Go [CAS24b, Wal21, Zel20, LG20, Ulr20a]. Goal [Lau24a, Cly22, Koz21b, Kum21]. Godfather [Kim22]. Goes [Ack21b, MD23, Smi21k]. GoFly [Pot20b]. Going [Koz23c, Ros20a, Wal24b, Kum21]. Gone [Jon23k, BMRW20]. Gooaall [Coh22]. Good [Ano20-66, Mar23j]. goodbye [Hig20j]. Google [HVW20, Has20b, Kon24, Koz21a]. government [Cha20a, Raf20b]. GPS [Har21, Raf20b]. GPU [LW24]. Gracey [Gen23c]. Grade [Pat23b]. Grain [Smi22f, Usk21]. graphics [Cas20e].

Grapples [Gol23h, Fai20b]. Grappling [Sch21e]. Gravity [Moo21]. Great [Ano21-113, KP24, Koz23c, Per21c, Smi21l, BBS23, Smi20a]. greatest [Cro20]. Greedy [Cha23]. Green [Fai23b, Str23, Pat22]. greener [Hus24]. greenhouse [Gal21b]. Greening [Gol24e]. Greta [GEN24g]. Grid [AFH22, Dum22, Eis23, Fai21, Gol22f, KH24, Poo24a, Fai23c, Fai24a, Lee20, Moo21, Pat20]. Grid-Forming [KH24]. grids [Ano20-66]. Grit [Abb23]. Groceries [RZ21]. Groove [LDO22]. Gropper [Ano21a]. Gross [Der20]. Grosser [Der22a]. Groundbreaking [HPK22]. Grounding [Jon23g]. Group [Ano21-121]. grow [Fai20a]. grown [Gui20]. Growth [Bro22c, Pat24a, Smi21j, Smi21k, Ano20-44, Per20j]. guard [Arm20]. guardian [Per20k]. Guide [Ano21-144, Ano21-145, Ano22-131, Ano22-132, CFPP22, RZ21, Ano20-103, Ano20-104, Gol23c]. Guided [Jon23e, Smi23c]. Gun [Ros23b]. guy [Ano20-81].

H [Der23c]. Hack

[Zel20, AF23, Lew20, RG20]. hack-proof [Lew20]. Hackable [Cas24e]. Hacker [Cas23b]. hacks [Hoc22]. Hall [Zor22c]. Ham [Cas24e, Gol24i]. HamMessenger [Tho21a]. Hand [Coo21, You22]. Handful [PRR23]. Handheld [Cas24e, Wal21]. Handie [Bre20]. Handie-Talkie [Bre20]. handle [Moo20a]. Hands [Cas20d, Cas20e, Cas20g, Cas20f, Cas20h, Cas21a, Cas21b, Cas22g, Cas24c, Ehn23, Gob23, Gol23e, Laz21, Lew20, Mar24l, Pom21, Sar20, Sch20b, Sch20c, Sch20d, Sch21c, Tem20, You22, Zel20, Pre20]. Hands-On [Gol23e]. Happen [Har21]. Happens [SH22b]. happier [AS22]. hard [Hut21, Sca22, VH20]. Harder [Cha23]. Hardest [Chi21]. Hardware [Gen24b, Per24b, Ros22a, Sch23c, Smi21l, LA22, Per22b]. Hardy [Ros22e]. harming [And20h]. harnesses [Koz20c]. Harsh [Rab22]. Harvard [Der23d]. Harvest [Jon22e]. Harvesting [Usk21]. hasten [JJS22]. Hate [Cas23b]. Having [Gol23b]. head [Per21a]. head-up [Per21a]. headphones [Hsu20c]. healing [Ano20-67]. Health [SHP24a, Str20b]. healthcare [Ano20-96, Ano20-97]. heap [Smi22g]. Hear [Bra21, Koz20b, Tho21b]. Hearing [Mar22i, Mos22, Sch20b]. Heat [MZR21]. Heating [Ano23-137]. Heavens [Sch23c, Sch24b]. Heavy [Der23a, Ack22b]. Heavy-Duty [Der23a]. Heights [Jon23b, RK22, Gol24g]. Helena [Koz21a]. helicopter [Pot20a]. Hell [Cha21a]. Hellish [MZR21]. Hello [AG20]. Help [Ack22a, Bau24, CPBZ21, Gol22a, Jon22b, Gal20a, Pre20, SHD20]. Helped [Bro23c, Kim22, Ros22a, Fai23c, Per22a]. helpful [CHM21]. Helping [Koz20b]. Helps [Gol22b, Bar20, Der22c, HVW20]. Her [Ack24a, Der22d, Der22g, Der23b, GEN24g, Cas22d]. Here [Ano21-144, Ano22-131, Ano22-132, Bou20, Car24, GK24, Per20l, Sch20b, Ano20-103, Ano20-104, Ano21-145, YS20, Hut20, Jag22]. here-and-now [YS20]. Hi [Cas22a, Gol24j, Lew23]. Hi-Fi [Gol24j, Lew23, Cas22a]. Hidden [GH23, TB21, Mar20j]. Hide [Baj24, LA22]. Hide-and-Seek [Baj24]. High [Ano21-114, Coo22, Fai24a, Gan24, Van23, You22, And20c, Cly23, Has20a, Has20b, YS20, Zor23a]. high-flying [Zor23a]. High-Numerical-Aperture [Van23]. High-Quality [Gan24, Has20b]. High-Resolution [Coo22]. High-Tech [Ano21-114, You22, Has20a, YS20]. high-temperature [Cly23]. High-Voltage [Fai24a]. Highway [Mar23k]. hiring [Hsu20a]. Historic [Gol24f]. History [Ack24b, Eis23, Jon23a, Mar21e, Mar23b, Mul24, Cro20, Scu20a]. hitch [Pot20a]. Hits [Gen24b, Nor20b]. Hoard [Har22]. Hoards [GH23]. hockey [Zor23a]. Hold [Pow24]. hole [Bou20]. Home [Ack23d, Ano23-137, Bor24, Cas21c, Cas23d, Jon23b, MZR21, Szk21, Wat24, AG20, Has20b, Per20g, Sar20, Sch20c]. home-automation [Has20b]. Home-Brew [Szk21]. Home-Brewed [Ack23d, Cas23d]. Homemade [Sch22b]. homes [Ano20-96, Ano20-97]. homomorphic [Moo24c]. Hopes [Per20d, Rak24a]. Horizon [Zim21, Bou20]. Horse [Moo23b]. hospital [Ano20-68]. Host [Dum23, Lau24b]. Hot [Cas21d, Szk21, Per20j]. hour [Ack22b]. Hsiang [Der22c]. Hu [Per20c]. Hubble [Pot24]. Huge [Coo21, GOL24d, Moo20e]. Human [Ack23a, Ack23c, Bro21, Smi22h, BBS23]. Humanity [Ack23b, Gib21]. Humanoid [Ack24d, GK24]. Humans [Ack21b, Gol22b, Gol23g, And20h, Gol22a]. Hummer [Smi22e]. Humming [Gen24j]. Hundreds [Usk21, Dum20]. Hunt [Pow24, Sch22b]. Hunts [GH23]. Hurdles [Pat20]. Hybrid [Gen24d, JAR24, Moo24d]. hybrid-electric [JAR24]. Hybrids [MCG24]. Hydrogen [Fai23b, MD23, Zor24a, Str23]. Hydropower [Smi22g]. hype [Fun20, Hig20i]. Hyper [Smi21f]. Hyperloop [Smi21f]. Hypersonic [Ros20a].

IBM [Aro24, MCG24]. IBM’s [Cho23]. Ice [Rab22, Tei23, Gen24i]. Iconic [Mar23c]. Icy [Cha21a]. ID [Gen23e]. Idea [Per24b]. ideal [Zor20]. ideas [Cly20]. Identifies [Koz23a]. identities [RS23]. identity [Per20b]. IEEE [Ano20-70, Ano20-71, Ano21-31, Ano21-115, Ano21-117, Ano21-118, Ano21-120, Ano21-119, Ano21-121, Ano21-116, Ano22-122, Ano22-123, Per22c, Ano21-57, Ano21-58, Ano21-59, Ano22-59, Gol21, HPK22, SM24, Zor23b, Zor24b]. If [BS22, Mar20f, Sch24c, Sha23, Pou22, Raf20b]. II [Hoc22]. image [Bou20, PK20]. Images [Koz23a, Sch23c, Sch24b]. Imagination [Gol24a]. Imagine [Bro23c, CD21]. Imaging [Jag22]. IMEC [HH24]. immersive [WAC^{+}20]. Immortality [Mar23c]. Impacts [Cab24]. Implant [Wal24b, Dut22]. Implants [Bou21, SH22b]. important [Ano22-130]. impossible [VH20]. Impossibly [Smi22c]. Impractical [You22]. impressive [Smi22i]. Improve [Per24a, Cho20, Gol22b]. Improvement [TGLM21]. improving [Per23]. included [Hig20f]. Income [Ano21-121, Per22c]. increases [Per20i]. increasingly [Lau23b, SB22a]. Indeed [Smi21d]. Indefinitely [Pes21b]. Indelible [Zor24b]. Independence [Ack22a]. Indeterminate [Ano20-44]. India [Fai20b, JJS22, Jon23f]. Indoor [Smi21n]. induction [Sch20b]. Industrial [Gen23a]. Industry [GEN24g, Jag22, Kim22, Moo24a, Pat24a, WC23, And20g, Koz21b]. inertial [Ano21-134, Per21a]. Inexpensive [Kuh24, Rab22]. Inflation [Per22c]. Inflation-Adjusted [Per22c]. information [Der20]. Infrastructure [Fit21, Gen24j, Gol22c]. Ingenious [Mul24, ROK24]. ingenuity [Luc20b]. inhaler [And20b]. Ink [HF22]. Inkjets [BF24]. Innovate [Ano20-72, Ano20-73, Ano21-122]. Innovation [Gen24i, Moo20c]. Innovator [Bro22d]. input [Cas20f]. input/output [Cas20f]. insde [Ano22-48]. Insight [Zor22c]. Insights [Per22c]. Inspire [Cla24]. Inspired [Sto23, Gen24h]. Instacart [RZ21]. Instant [Sch21d]. Institute [Gol22d, Raf20a]. Instrument [Ake23]. Instruments [CP21, Ano22-51]. Insulin [KK21]. Insurance [Ano21-121]. Integrated [Pau22]. Intel [Rak24a]. Intelligence [Bro21, Gol23a, Lau24a, CHM21, Fun20, Hut21]. intelligent [Ano20-90, Pou22, Pri21]. intended [Der21d]. Intentional [Jon23g]. Interactive [Ros22a]. Interconnector [KP24]. Interconnects [CPBZ21]. interface [Per20o]. interference [Bra21]. Internet [AFH22, Der22b, Hig20b, Hig20d, Hig20e, Hig20f, Hig20h, Hig20i, HP21a, HP21b, HP21e, HP21f, HP21c, HP21d, Mit20, PRR23, Per23, PSD22, Sha23, Shi22, Tow21]. interoperable [Koz21b]. InterPlanetary [PSD22]. interrupted [Ham20a]. Introduction [SM24]. Invented [Ano20-78, Cro20]. Invention [Gen24c, Lau24a, SM24, Ano22-130]. inventions [Gol24e]. Inventor [Mar24f]. inventors [Moo20c]. Inverters [KH24]. Invisible [Kuh24, Gol22c]. ion [Kum21, Mur23]. IoT [Ano20-79, Hig20c, Pes20b]. Irene [Gen24l]. Irresistible [VA23]. Island [Koz21a, Koz24c]. isn’t [Cha23, Koz23c, MMM24]. Italian [WA24]. items [CKG^{+}21]. Itself [Bra21, Smi21l]. Iverson [Der21e].

Jacob [Per21b]. Jam [Har21]. James

[Mar23c, Sch21b]. Janelle [Ano23a]. Japan [Fru20, McC20]. Java [Ano20-95]. Jenkins [Fai23c]. Jersey [Sha23]. Jesse [Fai23c]. Jetson [Cas20d]. Jim [Gen23a]. Job [Ack22b, Gen24a, Per20f]. Jobs [Der21e, Ack24d, Koz20a, Per20j, Per20m]. John [Der23a, Per20n]. Jolt [Dol23a]. Jonathan [Der23d]. Journal [And20f, Ano20-91, Ham20d, Ham20b, Hsu20c, Wal20b]. Journalists [Gol24c]. Journey [Der23b, GEN24g, Moo24e]. Joy [Str24a]. JPL [BD22, Cha21a]. Juan [Bro22b]. Jugular [Wal24b]. Juice [Jon23h]. Julie [Der22g]. Julien [Moo24b]. Jump [Ros22a]. Jump-Start [Ros22a]. JUPITER [Dum23]. Just [BF24, Gol24i, Koz23c, Pat23b, MS22].

K5 [Cas24e]. Kahn [Gol24f]. Keep [Ano21-123, Pes21b, BMRW20]. Keeps [Gen24f, Gen24j, Gen24l, AG20, Ulr20a]. Kelp [Pat24a]. Kenya [Der22e]. Kepler [Mit20]. key [YS20]. Keyboards [Mul24]. Ki [Kim22]. kicked [Per23]. Kicking [Gol23c]. kid [PK20]. kids [HVW20]. Killer [Fai21]. kilometer [Jaf22]. Kilometers [Smi21e, Cly23]. Kim [Kim22]. kind [RHBM22]. kinds [AMK^{+}20]. Kirchhoff [BSAH21]. kit [Cas20f]. kits [Per20e]. Know [Cas24f, Cas20g, Per23]. Knowledge [Mar23k]. Known [Mar23c]. Korea [Kim22]. KrakenSDR [Cas22b]. Kubernetes [Cas21c]. Kyle [Zor23a].

L [Der23d]. Lab [Moo24e, Ham20a, Hec20, Tem20, Moo23b]. Lab-to-Fab [Moo24e]. Lags [Smi22a]. Lake [Gol24i]. laminography [LA22]. lamps [And20h]. Land [Bau24, DM20, Ulr22]. landing [Fou21]. Lane [WPBS23]. Language [Cas22c, BBS23]. language-learning [BBS23]. Languages [Cas21d, CG23, Cas20i]. Lapse [Sch23c]. Laptop [Smi22l, Mar20c]. large [Per20f, VH20]. Larry [Per20o]. Laser [BVKR23, Gen24m, Mar22l, NYI24, Moo23b]. laser-based [Moo23b]. Lasers [BVKR23, BBFS22, NYI24, Ros23b, Cho20, Koz23b]. last [Som21]. Lasted [Shi22]. lasting [Ano20i]. Later [SB22a, Smi20d, Smi23b]. Latest [JPWB21, MJH^{+}21, SAS^{+}21, Cas20i, CKG^{+}21]. Lattice [BBFS22]. Launch [Fou22, GO22, Pot20a, Sto20]. launched [Bro22b]. Law [BSAH21, Der23d, Smi22i, Boy24, BSAH21, CPBZ21, Moo24d, Moo24b, Per20c, RK22, Van23]. Law-for [Smi22i]. layer [Gol22f]. lead [AMK^{+}20]. Leader [Der21f]. Leaders [Per24a]. leads [Der22h, Dum22]. league [Ano20-51]. leaks [MJS20]. lean [Hub24]. leaner [Hus24]. Leap [Cho23, Gen24c, Zor22b]. Leapfrog [Rak24a]. Learn [BBS23, Cas21c, Kas23, Per24b, CHM21, Sca22]. Learned [ALRW21, Goo23a]. Learning [Ack21b, Ano21-125, Ano21-126, GH23, Ham21, MMM24, MSS21, RZ21, Sch21a, Str22a, TGLM21, Ano21-113, BBS23, Cas20d, DHOW20, Hig20d, LDO22, Moo20e, Per20n]. Learns [BBS23]. lease [Red20]. Leaves [Bea22]. led [Hut20, Per22a]. Left [Zor24b]. Legacy [Ack24a, Ano20g]. legal [Hoc22]. Legged [Ack24d]. Lego [Tem20]. Leland [Jon23a]. Length [Smi21d]. Length-to-Beam [Smi21d]. Less [Sch22a, Per20b, Sch22d, Smi22e]. Lessons [CG23, Cha21b, Gen22, Mar22j, Goo23a]. Let [Cas20b, Der21e, Sch21b, WAC^{+}20]. Lets [Der22d]. Letters [Tho21b]. Letting [Cas24a, RHBM22]. Level [Gol22b, Kas23, And20g]. Leveraging [AS22]. LGBTQ [Gen23c]. Libi [Cas24a]. Lidar [WPBS23, Per21a]. Lie [Smi21d, Smi21e, Smi21f, Smi21c, Smi21g, Smi21h, Smi21i, Smi21j, Smi21k, Smi22a, Smi22b, Smi22c, Smi22e, Smi22j, Smi22i, Smi21b]. Life [Ano20-81, Bow21, Cas20c, Gen22, Cas20a, Der20, Fru20, Gen23b, Har22, Pri21, Red20, Smi22d, Cas20b]. Lifesaving [Per22a]. Lift [Fou22]. Light [Ano20-82, Gen23a, Mar21j, Mos22, SC23, Sch21a, Ano20-67, Ano21-50, AF23, Ham21, Koz20c, Mar20i, Raf20a]. lightens [Ham20d]. Lightest [Jon22c]. lighting [Has20b]. Lightmatter [Sch21a]. like [Moo22c, WB21, ZL20b]. Limbs [FBMI23]. Limited [And20d]. Limits [Smi21d, Ano21-113, Ano21-123]. Limor [Cas23a]. Line [Gen24d, Poo24a, Ulr20b]. Linear [Svr23, Gar20]. Lines [Cha20a, Has20a, Has20b, Luc20a, Per20b, Per20l, Per20n, Per20e, Per20h, Per20o, Pre20]. link [Dum20]. liquid [Smi22k]. Listen [Pri21]. lists [Ano20-95]. lite [LG20]. Literally [Ros22a]. Lithium [Cha24, Kum21, Mur23, Kum20]. Lithium-ion [Kum21, Mur23]. lithium-metal [Kum20]. Lithography [Boy24, Van23]. Little [Coo21, Mar23c, Mar22j]. Liu [Wal24a]. Live [Mar23c]. lively [Ano20-82]. lives [Ong20]. Living [De 22]. load [DM20, Ham20d]. Location [Ano20-83]. logjams [DM20]. Long [Bro22a, Mur23, Smi21d, Smi21m, And20d, Hig20j, Ros22d]. long-range [Ros22d]. long-term [And20d]. Look [Ano20-84, PK20, Per21a, Sch20d, Str24b, VA23]. Looking [Ano21-124, Gol23d]. Looks [Tei23, Ano21-27, Ano22-55]. Loop [Ack23a, Bro21, Sch20b]. loopholes [Hoc22]. LoRa [Cas24d]. LoRa-based [Cas24d]. lose [Per20i]. Lot [Coo21, Per24b, Cas20d, Cas20f]. Love [Cas23b, Cas22a]. loving [Der22b]. Low [Bow21, Ver24]. Low-Cost [Ver24]. Low-Dropout [Bow21]. Lower [Lew23]. LTA [Koz23d]. Lu [Der22c]. Luke [Str23]. lunar [LSM23]. Lurk [Cas23c]. luxury [GA20, Has20b].

Mach [Ros20a, Sch21e]. Machine [Ano21-125, Ano21-126, Bro23a, Bro22c, Bro22b, Bro22d, DHOW20, GH23, Hig20d, MMM24, Mar24i, Pau21, RZ21, SB22b, SB22c, Str22a, Van23, Cas20d, Cas23a, HPK22]. Machine-Learning [Str22a, Cas20d]. Machines [Gen24f, AG20, Smi20c]. Macro [HP21a, HP21b, HP21e, HP21f, HP21c, HP21d, Pes20a, Pes20b, Pes20c, Pes22, SP21a, SP21b, SP21c]. Made [Cas21c, Cha21a, Gol24f, Gol24g, Pot24, Arm20, Hsu20a, Per20i, Wea20]. Madni [Per22a]. Magaly [Der22d]. Magazine [Ano21-127]. Magdalene [Der22e]. Magic [Gen23a]. Magnetic [Shi22]. Magnetometer [Sch22d]. Magnets [BBFS22]. Mail [Mar21h]. Mainspring [Svr23]. Major [Der21f, Pau21, Cly20]. Make [BVKR23, Bro23a, Der22d, Gol22a, Gol23g, Kim22, LH24, Lew20, Pat23b, Pat24a, Poo24b, RZ21, AS22, Ano20-66, Baj22, Fai23b, Has20a, Hus24, Jes21, Koz21b, Moo24c, Pat20, Pat24b, Per20b, PSD22]. Makeover [Ulr22]. Maker [Jon24b]. Makers [Cas24d, Hsu20a]. Makes [Coo21, Gen23d, Tho21a, Cas20d, Gen24k, Hut21]. Making [Bro22a, Fai23b, Gol24f, HVW20, Hig20e, Gol23d, Hut21, Per21b]. male [Per20b]. Maluta [Der22e]. malware [Ham20c, Sha23]. Man [Koz24a, Mar23c, Abb23, Reg23]. Manager [Aro24, SG24]. Managing [Cha20a]. Manan [Gen24h]. manifests [Str24b]. manufacture [Gib21]. Manufacturers [ACDG21, Pes21b]. Many [BVKR23, And20h, Hut20]. Map [Tow21]. Mapped [Wal22]. Maps [Wal22]. Marcellus [Der21f]. Maria [Der22f]. marine [Der22h]. Mark [Cho23, Zor24b]. Market [Ash23, Gen24b, Has20b]. Mars [Cas20c, Pot20a]. masks [Pre20]. Mass [Smi23a, Has20b]. mass-market [Has20b]. Mass-Produce [Smi23a]. Masses [Gen23e, Per22b]. Massive [BBFS22]. Masterpiece [Jon22d]. Masthead [Ano22-124, Ano22-125, Ano22-126, Ano22-127, Ano22-128, Ano23-123, Ano23-124, Ano23-125, Ano23-126, Ano23-127, Ano23-128, Ano23-129, Ano23-130, Ano23-131, Ano23-132, Ano24-131, Ano24-132, Ano24-133, Ano24-134, Ano24-135, Ano24-136, Ano24-137, Ano24-138, Ano24-139, Ano24-140, Ano24-141, Ano24-142]. Match [Gol24g, Smi22k]. Materials [ROK24, WA24]. Math [Cho21]. MATLAB [Ano21-128, Ano22-50]. Matrix [SZ23]. May [MMM24, Pat23b, Pot20b, Pow24, Smi23c, Ano21-113, Cha24, Dol23b, Ros21, Sch22e, ZL20b]. Mayflower [And20a]. McDonald [Ano20-87]. Me [Ano22-122, Bro23c, Sha23, Ano21-116]. mean [AMK^{+}20, Der20]. Means [Ack22c, Cla24]. Meanwhile [Ano20-85]. Measure [Mar23j, Smi21n]. Measures [Gol24h, Pre20]. Measuring [Smi21e]. Meat [Smi22h]. Meat-Eating [Smi22h]. Mechanical [Ano24-143, Mar24h, Mie22]. Media [Str24b, Nor20a]. Medical [Bea22, Ano20-68]. medicine [CFPP22]. Megatalent [Ano24-143]. Megatruck [Jon22c]. Melanie [Der22h]. Melt [NYI24]. melts [Gal20a]. Member [Ano21-117, Ano21-118, Ano21-119, Ano21-121, Gol21, Ano20-71, Ano20-93, Ano21-120, Ano22-123]. membership [Ano20-71]. memories [Ano20-88]. Memory [BSAH21, Cas22g, Moo22b]. MEMS [Zor20]. MEMS-based [Zor20]. Mental [SHP24a]. Menus [Bro23a]. Merchant [Pat24b]. Merely [Smi21f]. Merryl [Der20]. Mesh [Cas24d]. Meshtastic [Cas24d]. Mesmerizing [Sch23c]. mess [Hut20]. Messages [Ham20b]. Messy [She24]. Meta [Gen22, Gen24j]. Metal [Cha21a, GO22, Cas20h, Kum20, Sch22c]. metal-rich [Sch22c]. Metalenses [Gob23]. Metallic [GO22]. Metals [GH23]. Metamaterials [Pou22]. methane [MJS20]. metric [Smi22i]. Micro [HP21a, HP21b, HP21e, HP21f, HP21c, HP21d, Pes20a, Pes20b, Pes20c, Pes22, SP21a, SP21b, SP21c]. Microcomputer [Ack23d, Cas23d]. Microcontroller [Ros22b]. microcredit [Fai21]. Microgrids [FOSMD23]. Microgried [FOSMD23]. MicroLED [Ham20b]. microLEDs [Moo23b]. Micromanufacturer [Cas22d]. Microphone [Sch24a]. Microsatellites [MJS20]. microscope [Tem20]. microscopy [WAC^{+}20]. Microstory [Gol24j]. Midair [Mie24]. Middle [Aro24]. midst [Der22b]. might [AMK^{+}20, And20h, Ham21, HS23, WB21]. mighty [MS22, Smi22g]. Migrating [Sch24a]. Mile [Koz21b]. Military [Cla24, Har21]. Millimeter [Gal20a]. million [Hec20, Pot20b]. millions [Smi23b]. millisecond [Pan20]. Mind [Gen24c, Kim22]. Mind-Set [Gen24c]. Minerals [GH23]. Mines [Bau24]. minikit [Cas20g]. Minimalist [Str22a]. Mining [Sch22c, Zor22d]. Miracle [Smi23b]. Missile [Cha21b]. Missiles [Ros23b]. Mission [GO22, Red20, Jon24a]. Missions [BD22, BMRW20]. Missy [Mar23k, Ros23a]. mist [And20b]. mix [Zor22a]. Mobile [Ham20c]. modded [Cas24e]. Model [Cas22f, Jon24b, Ras20b]. modeler [Fai23c]. modeling [Pre20]. models [Hut20]. Modern [Ack24a, Bro23a, Cor21b, Ano22-130, Lee20, Smi22i]. modular [Pat24b]. Molly [Koz23c]. Moment [Gol23b, HF22]. Moments [Lau24a, PRR23]. Momentum [Kum21]. Money [Bow21]. monopoly [Koz21a]. months [Scu20b]. Moon [Jon24a, BD22, Fou21, LSM23, Zor22b]. Moore [Boy24, CPBZ21, Gol24h, Moo24d, Moo24b, Per20c, RK22, Smi22i, Van23]. Morgan [Gen23b]. Mortalities [SB22a, Ano20-88]. Most [Ack23b, ALRW21, Cas24e, Lev21, Ano22-130, Smi22i]. motion [Rae20]. Motor [Der21h, GP22, Mar21k]. Motorized [Ham20d]. Motors [ROK24]. mouse [Cas24e]. Mouse’s [Ano24-144]. Mouse’s-Eye [Ano24-144]. Move [Moo23a, Str24a, Ack22b]. Movement [Bou21, Smi22l]. Movie [Smi21l]. Moving [Gob23, Gen24k]. MPGs [Sch23b]. Mr. [Per23]. Much [BF24, WB21]. Muddy [Sch22c]. Muffling [Lew23]. Multiphysics [Ano21-63, Ano20t]. Municipal [Jon24b]. murky [Fou22]. Musk [Smi21f]. must [LSM23]. mutates [Ham20c]. mute [Ros20b]. My [Bro23c, Coo22, GK24, Mar24e, Lev20]. Mysteries [Jon23i]. mysteriously [Fru20]. mystery [MJS20]. Myth [Bro22c, Bro22b, Bro22d, SB22b, SB22c].

Naffziger [Moo23a]. Naked [LA22]. nano [Cas20d]. Nanometer [TB21]. Nanometer-Scale [TB21]. Nanoparticles [Cha24]. Nanoscale [Gen24m]. Nanostructures [Gob23]. Narrative [Mar21i]. narrow [Fou21]. NASA [BBFS22, Der21f, Fou21, Fou22, GO22, JAR24, Jon24a, Pot20a, Pot24]. nasal [And20b]. nascent [Rak24a]. nasty [Sca22]. national [Sto20]. Nature [Gol23e]. navigation [Ano21-134]. Navy [Gen23c]. Nazi [Mar21f]. Near [Smi21e]. nearby [Hsu20c]. Need [Ack23b, BBS23, Cab24, Cas24f, Fit21, KH24, Cas20f, Gal21b, Gol24e, MMP22, Pan20, Ros20a]. need-for-speed [Ros20a]. Needed [MZR21]. Needs [BD22, Fit21, BSG22, Ham21]. neighbors [Ano20-66]. Nerve [Dol23a]. Nest [Has20b]. nets [Moo22b]. Network [BSAH21, Cha21b, FOSMD23, Fra21, Gen24j, Hig20f, Tow21, Fai24a, Koz21b, SMD20]. Networked [PRR23]. Networking [Bro23a, Has20b]. Networks [Cho21, MSS21, Ham21, Pan20]. Neural [BSAH21, Cha21b, Cho21, Fra21, MSS21, Dut22, Ham21, Moo22b]. Neural-Network [BSAH21]. Neuralink [Wal24b]. Neuromorphic [Coh22]. Neurons [FBMI23, Poo24a]. Neuroprosthetic [Str22b]. Neurorights [SG22]. neurotech [SG22]. Neutrino [Jon23i]. never [BS22]. News [ACDG21, ARPH22, ACLS22, And20b, AMK^{+}20, And20d, And20e, And20g, Ano20h, Ano20-46, Ano20-49, Ano20-51, Ano20-83, Ano20-82, Ano20-85, Ano20-87, Ano20-89, Ano20-50, Ano21-129, Ano21-130, Ano21-131, Ano21-132, Ano21-139, BGR^{+}21, BJZM24, CJGC24, CKG^{+}21, CSB22, CKX^{+}23, Dum20, Fai20a, Fai20b, Gal20a, Gar20, GHG^{+}22, GMB^{+}24, Gen24b, GDUF22, GAS^{+}24, Ham20a, Ham20c, Hec20, HPHU21, Hsu20a, HCP^{+}23, JPWB21, Koz20b, KPG^{+}21, KPCA23, Lau23a, MLF^{+}24, Moo20c, Moo20a, MJH^{+}21, MDC^{+}22, MHLC23, NS20, Nor20a, Ong20, PCRC24, Pat20, PPUS22, PMPG23, PLUM23, Per20a, Pot20b, Pot20a, Raf20a, Raf20b, RCAP23, Ras20a, Ras20b, Red20, Ros20b, RPC^{+}23, SAS^{+}21, Scu20a, SCPG22, SCPC23, SZCB24, Str20b, SCMP22, SGS^{+}23, SHC^{+}24, SHP^{+}24b, Ulr20a, WAC^{+}20, ZJL^{+}24, ZSG^{+}24]. Next [Moo24b, Pal22, Pau22, Van23, Zor22b, Ano21-113, Ano22-58, Fai24a, Gui20, Koz24d, Ano20-68]. Next-Gen [Pau22]. next-generation [Fai24a]. Ng [Str22a]. Niche [Cas22d]. Night [SC23, Sch24a, AF23, LSM23]. Nile [Smi20b]. No [Baj24, Cha20b, Jon23k, Kuh24, Lau23b, LA22, Sch21b, Ack21a, BMRW20, Has20a, Smi22d]. No-Fuss [Kuh24]. no-tech [Has20a]. Nobelist [Mar22l]. node [Moo20d]. nodules [Sch22c]. Noise [Ros20b]. Noise-canceling [Ros20b]. nonsense [Moo20d]. normal [Per20d]. North [Ano20t]. Nose [Str22b]. Nothing [Gol22e]. notices [Cha20b]. novel [Gol24e]. November [Bro22b]. now-famous [Bou20]. Nowhere [Smi21e]. Nuclear [Baj22, Jon22h, Kon24, Pat24b, And20d, Cly20]. Nuclear-powered [Pat24b]. number [VH20]. Numbers [Smi21d, Smi21e, Smi21f, Smi21b, Smi21c, Smi21g, Smi21h, Smi21i, Smi21j, Smi21k, Smi22a, Smi22b, Smi22c, Smi22e, Smi22j, Smi22i]. Numerical [Van23]. Nvidia [Cas20d, Smi24].

obey [Ack20]. Object [Mar23b]. obscure [Cho20]. Obsolete [SH22b]. occur [Koz20c]. Ocean [Der22g, Sch22c, Zor22d]. Oceans [Baj22]. Ode [Ano23-133, Mar22k]. Off [Cly24, Fai21, Fou22, DM20, Per23]. Off-Grid [Fai21]. Off-the-Shelf [Cly24]. offense [And20h]. offers [Ano20-71, Lew20]. officer [Gen23c]. Offices [GEN24g]. offs [Lau24b]. Offset [Smi21j]. Offshore [Fai23a, MD23]. Often [Ano23a, You22, Smi22e]. Ohm [BSAH21]. oil [Her20, Smi20f]. Old [Abb23, Ack22a, Cly24, Smi21f, WA24, Zor22c, Cas20e, Hec20, Lee20, LDO22, Ano20-87]. oldest [Mie22]. OLED [Pal22]. Olsen [Der22h]. Olympics [Lau24b, Lau24b, Pat22]. Omicron [Wal22]. Omnivorous [Svr23]. on-the-spot [Per20a]. Once [Gol23a, RHBM22]. One [Ake23, Ano21-116, Aro24, BVKR23, GEN24g, Gol23h, Gol24g, Koz24a, Sch21b, Smi22a, Zor22b, Ack21a, Ano21-138, Cha20b, Chi21, ZL20b]. Online [Gen22, RS23, Pes21a]. only [Gol22a, RK22]. opacity [Hut21]. Open [Ack24c, Gen24b, Ver24, Koz21b]. Open-Source [Gen24b, Ver24]. opener [Lew20, Zor23c]. operations [Sch22e]. operators [Koz21a]. Optical [Gen24c, Nor20a, Sch21a]. optics [Ano21-134]. Optimized [Ano20-78]. Option [Pat24b]. Options [Mul24]. optophone [Tho21b]. orbit [Mar20c]. organizations [Str24b]. organize [Dum20]. origin [Per20l]. original [Ros20a, Smi23b]. Oscar [Gen23a]. oscillators [Cas20b]. Oscilloscope [Cas20e]. Other [Bro22d, Gol23g, Kon24, RZ21, Mit20]. Our [Ano21-122, Ano21-124, Cas20i, Cas21d, Gol23e, Gol23g, GK24, MZR21, Smi22d, Ano20-72, Ano20-73, RS23, WB21]. outposts [LSM23]. output [Cas20f]. outputs [Sch23b]. Overnight [Bro22a]. Oversees [Der21f]. Overtake [Ack21a]. Own [Cas22g, Sch23c, Szk21, WB21]. oxide [Jes21].

Pac [Reg23]. Pac-Man [Reg23]. Package [Rab22]. Packetizing [AFH22]. packs [Cas20d]. paid [Per20b]. Painless [Cas20f]. Painting [Fra21]. Pallets [Ros22d]. Pan [Bro22b]. Pancreas [KK21]. Pandemic [ALRW21, Ano20-88, Hig20g, Pes21a, Ano20-92, Cha20a, Ham20a, Has20a, NS20, Per20d]. pandemics [YS20]. Panel [Bor24]. Panels [WA24]. Paper [HF22]. Parachute [Ros22a]. Paradox [Smi21n]. Parcel [Smi22f]. Paris [Lau24b]. park [Fai20b]. Part [Pot24, Smi22f, SH22b, Cas22f]. Particle [Boy24, Gen24l, Kuh24, Arm20, Gar20]. particle-accelerating [Gar20]. particles [AMK^{+}20]. Partridge [Der22b]. passes [Gar20, Koz20c]. Passive [Cas22b]. Password [Coo22]. Past [Ano23-134, Bre20, Bro23b, Cas24a, Cho23, Cor21b, Goo23b, Mar20a, Mar20b, Mar20c, Mar20f, Mar20g, Mar20i, Mar21d, Mar21e, Mar21f, Mar21h, Mar21k, Mar21i, Mar21j, Mar21c, Mar21g, Mar22i, Mar22j, Mar22l, Mar22k, Mar22a, Mar22b, Mar22c, Mar22d, Mar22e, Mar22f, Mar22g, Mar22h, Mar23b, Mar23d, Mar23e, Mar23f, Mar23g, Mar23h, Mar23i, Mar24b, Mar24k, Mar24f, Mar24d, Mar24g, Mar24e, Mar24h, Mar24c, Mar24i, Mar24j, Str21, Ano22-130]. paste [Per20o]. Patch [Sha23]. Patent [Gol24i]. Path [KH24, LW24]. Pathfinder [Koz23d]. Paths [Moo22a]. patients [Ano20-68, Arm20]. Patryk [Gen24i]. pattern [Smi22j]. Pay [Cal24, Per21c, Per20h]. PC [Bro23a]. PCB [RG20]. Peak [Smi20f]. Pedaling [Sch20c]. Peer [PSD22]. Peer-to-peer [PSD22]. Peering [Pes21a]. People [Ack23b, Cas24a, SH22b, Ano20-68, Fou21, Gol22d, Sch20b, Sch20d, Tho21b, Wea20]. Pepsico [McC20]. Percent [KH24]. perform [AG20]. Performance [Poo24a, Ver24, Moo20b, Moo22b]. Perhaps [Baj22]. Peril [Sch21e]. Perilous [Mar21i]. Perry [Gol23f, Gol24c]. Perseverance [Pot20a]. person [And20e]. Personal [Der22e, Pot20b]. perspective [Ano20-47]. Peter [Gen23c]. Pets [Fit21]. Phase [Moo24b]. phenomenon [Cho20]. Philip [Gen22, Zor23b]. Phoenix [Rae20]. Phone [MZR21, Tho21a]. Phone-Free [Tho21a]. Photon [Ano20-89]. Photonic [Ham21, NYI24]. Photonics [BVKR23, Cho20]. Photons [SC23]. physics [Arm20]. Pi [Sch23a, Cas21c, Ros22b, Sch23c]. Pi-Based [Sch23a]. Pichardo [Der22d]. Pico [Ros22b]. Picture [Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-51, Ano21-52, Ano21-114, Ano21-136, Ano22-56, Ano22-57, Ano22-58, Jon22f]. Pie [Smi22h]. pieced [Bou20]. Piezoelectrics [Gan24]. Pillai [Goo23a]. Pilot [Ros23a, Ong20, Sch22c]. Pilots [Ros22d]. Pinch [Cly22]. Pioneer [Str22a]. Pioneering [Moo23a, Moo21]. pivot [Per22a]. Pixar [Smi21l]. Placentas [Zim21]. places [And20h]. plain [MS22]. Plan [HH24, Pot24, RS23, Per20a]. plane [Per22a]. Planet [Lev21, MZR21, Pot20a, WB21, Ano21-136]. Planetarium [Mar24c]. Planets [Pow24]. planning [MMP22]. Plans [Jon23b, Ham20a, SG22]. plant [Ulr20b]. Plastic [Sch24a]. plastics [Gol22a]. Plate [Smi22b]. platform [Sto20]. Play [Cas24a, Coh22]. players [Gol22b]. plea [Hub24]. Pledges [Fis24]. Plenty [Smi22a, Pou22]. plotter [Cas20g]. Plotting [Cas20g]. Plugging [LSM23]. Plus [Ake23, Ano20-71]. Plutonium [Pat23b]. pneumatic [Fru20]. Pocket [Rak24b]. Podcast [Gol23g]. Podcasts [Ano20-100]. Point [MCG24, Cas22e]. points [Ano20-99]. poised [And20c]. polarization [Koz20c]. Polaroid [Zel20]. Pole [LSM23]. Policymakers [Cha23, Hut20]. pool [Per20f]. Pope [Gen23b]. Popular [Ros22b]. Population [Bro22c, Smi21k]. Portrait [Ano21-133, Bou20, Gol23f]. Possible [Cla24, Gol24f, Fou21, Has20a]. Post [Rao24, Per20d]. post-pandemic [Per20d]. Post-Quantum [Rao24]. Potassium [Pat20]. Potent [Fai23a]. potential [Ano20-70]. potentials [Ano20q]. Power [AFH22, Ano21-56, AF23, BBFS22, Bar24, Bow21, CPBZ21, Eis23, Fai24a, Fit21, Gol24a, Jaf22, MSS21, Poo24a, And20d, Ano20i, Cas20d, Cly20, Cly23, Fai20a, Fai23c, Gar20, Hig20b, JJS22, LSM23, Lee20, Moo20b, Rak24a, Sch20c, Smi22i, Ulr20a, Ulr20b]. power-delivery [Rak24a]. Power-Line [Poo24a]. Powered [Ano23-137, Lau24a, MW20, Pat24a, Pat24b]. Powerful [BBFS22, Jes21]. Powering [Fai24b]. Practical [Cho20, Smi21d]. Pratt [Ack22a]. Precise [Lev21, WB21]. precision [BSG22]. Prepping [Rao24]. preps [Pot20a]. Prescription [And20e]. Prescription-strength [And20e]. present [Ano22-130]. Preserve [Baj24]. Preserving [RS23]. press [Ano20-50]. pressure [ALRW21]. prevent [Kum21, Per22a]. preventive [Pre20]. price [GA20]. prices [Ano20-45]. print [Ano20-50]. printed [RG20, Tho21b]. printer [And20f, Zel20]. PrintFixer [And20f]. Printing [BF24, SC23]. Privacy [Pes20e, RS23, Ano21a, Hig20g]. private [Hoc22, Hsu20b]. Prize [Pot20b]. Pro [Per24a]. Probe [Cas21d, GO22]. problem [Gol22a, Hig20c, Sch21a]. Problems [Chi21, Per20n]. Proceedings [Ano21-59]. process [Cro20]. Processing [BSAH21, Moo22b]. processing-in-memory [Moo22b]. Processors [Moo22a, Szk21, Moo20b]. Proctor [Der21f]. Produce [Pat23b, Smi23a]. producing [Str23]. Product [Ash23]. production [RG20]. productive [AS22]. productivity [Fun20]. products [Der22f]. professionals [Per20b]. Professor [Ano20-90, Der23d, Mar22j]. Profile [Der22g, Der22h, Sac20]. Program [Ano21-116, Cas20a, Koz24b, Sch22c]. programmers [Gol22b]. Programming [Cas21d, CG23, Cas20f, Cas20i, Sca22]. Programs [Pat23a]. progress [And20d]. Project [MD23, Dum20, Sch23b, Tem20]. Projects [Coo21, Der21f, Gol23e, And20d, Cas20d, Cas20f, Cas22a]. Promise [Sch21e, LDO22, Pat20]. Promises [Gol24b].

Promising [ROK24]. Prompt [Gen24a]. Proof [Wal24a, Lew20]. Propelled [Gen24c]. Proposals [Fai23b]. Proprietary [Bor24]. Propulsion [Zor22c]. Pros [SG24]. Prosthetic [Jon23c]. protect [Gib21, Ras20b]. Protecting [Mar23j]. Proto [Mar20g]. Proto-Roomba [Mar20g]. Protocol [Cas23c, Cro20]. Prototyping [Cly24]. protracted [Smi22j]. Provide [Gan24]. Provides [Tow21]. providing [Har22]. proximity [Ano20-91]. Psyche [GO22]. ptychographic [LA22]. public [Ack21a]. published [Ano20p]. pull [HS23]. pumps [Fai21]. Purdue [Der22c]. Purpose [Chi21]. Pursuing [Ano20-92]. pushing [Ano21-123, Per21c]. Put [BVKR23, Gol22e, Som21, And20h, Cas20i]. Putting [Ros20b]. Pyramid [Smi20a]. pyramids [Smi22e]. Python [Ano20-95, Cas20i, CG23].

Qestions [Moo24a]. Quakes [Sch23a]. Quality [Gan24, Smi21n, Has20b]. Quantum [BS22, CFPP22, Cho23, Gen24c, MCG24, Pal22, Rao24, Ros22e, Scu20a, VH20, AMK^{+}20, CD21, Cas22e, Cho20, PK20]. Quantum-Centric [MCG24]. Quantum-dot [Scu20a]. Quantum-Entangled [Ros22e]. Quantum-Sensor [CFPP22]. Quartz [Wat24]. Quasiparticles [AMK^{+}20]. Qubit [Cho23]. qubits [Moo20a, VH20]. Quest [Jon23d, Kon24, Koz24a, Per20n]. Question [Cor21a]. Questions [Ack24a, Ano23a, Cas23a, Cas24a, Jon23a, Koz23a, Koz23b, Koz23c, Koz24a, Koz24b, Koz24c, Moo23a, Moo24b, Rao24, Ros23a, Sch22c, Str23, Str24a, Wal24a, Zor24a, And20d, Cha20a]. quick [DM20]. Quiet [Kon24, Bro22d]. Quixotic [Koz24a]. QWERTY [Mul24].

R&D [Kon24]. Race [Fai23a, Gen24d, You22, Moo23b, Ros20a, Wal20a, YS20].

Racers [Ack23c]. racial [Per20b, Str20b]. racing [And20e]. racism [Moo20c]. Radar [Cas22b, Cas24f, Tei23, RHBM22]. Radical [Har22, Cro20, Moo21, Ros22d, Zor23c]. Radio [Bra21, Cas22a, Cas24e, Cas24d, Laz22, MZR21, Mar24g, Ver24, Zor22b, BMRW20, Cas22b, DHOW20, Mar20d, Mar21f]. Radio-Spectrum [Laz22]. Radioactive [Cha21a, Gal20b]. radios [BMRW20, Bra21, Jes21, Koz20b]. rail [MMP22]. raise [And20d]. RAN [Koz21b]. Range [Der22d, Ros22d]. rankings [Cas20i]. Rao [Moo24a]. Rare [ROK24]. Raspberry [Sch23a, Sch23c]. Rates [Koz21a]. Ratio [Smi21d]. Rattner [Gen24k]. Ray [Ano21-124, Ano21-134, Ros22b, Ros23b, Sch22a, Cas20e, LA22, Mar20f]. Ray-Gun [Ros23b]. reach [Sch22e]. reached [Ano21-113]. Reactor [Cly24, Cly20, Cly22]. Reactors [Pat23b, Pat24b]. readers [Cas22a, LCBH20]. Ready [AS22, Cha23, HS23, MMM24, Smi23c]. readying [Fou22]. Real [BSG22, Smi21l, LCBH20, Zor20]. real-world [Zor20]. realistic [LDO22]. Reality [Fai23b, Ros22a, She24, Ano20-102, HVW20, Per21a]. Realize [Ano20-70]. Really [Cas24f, Sch24c]. Realm [Dum23]. reams [Har22]. reappear [Per21b]. Reasons [Ano21-116]. Reborn [Ano21-127]. receive [Bra21]. Receiver [Sch22b, Ver24]. Reckoning [Ano21-113]. reconfigurable [Pou22]. Reconnaissance [Zor22e]. Reconstruction [Abb23]. record [Hec20]. record-breaking [Hec20]. records [Dum20]. Recovers [WA24]. recovery [Gar20]. Recruit [Ano20-93]. recruiters [Per20f]. recruiting [Hsu20a, Per20f]. Recruitment [Ano21-135]. Rectangle [Poo24b]. recyclables [Cal22]. Recycling [Gol22a, Kum21]. Red [Mar21d, Ano21-136]. redefine [Per21a]. reduce [Dol23b, Gar20]. Reducing [Ham21]. ReefWorks [Der22h]. reengineered [Pan20]. reexamine [Hsu20a]. Reflections [Luc20b, Ano20-48]. Reflects [Ack24a]. regardless [Per20b]. Regularly [Lev21]. regulate [SG22]. Regulating [Gol23a]. Regulation [Mar23k]. Regulator [Bow21]. Reich [Shi22]. reinvent [AG20]. Reinvented [SM24]. Reinventing [Gol24i]. Reinvention [Fis24]. Relay [BD22]. relegated [Smi22g]. Reliability [Koz24d]. Reliable [WPBS23]. rely [Per20g]. Remade [Pot24, Cas20h]. remain [Pat20]. remakes [DHOW20]. Remembering [Per20o, Zor24b]. renaissance [Fai21]. Renamed [Smi21f]. rendition [WB21]. Renewable [KH24, Smi22g, Fai23b]. Repair [Smi22l]. Repairs [Pot24]. Report [Ano21-143]. Reports [Der22f]. reproduction [LDO22]. Require [Fis24, Ano20i, Ham21]. Required [Smi22b]. Requires [Ash23]. Rerecich [Der22f]. Reroute [Bou21]. Research [Ano21-56, Der22c, Ham20a, Koz23d]. Researchers [Jaf22, Per20e, Ros22e, Scu20b, WAC^{+}20]. reset [And20g]. Resignation [Per21c]. Resolution [Coo22]. Responds [Smi22l]. response [NS20]. responsibility [Cha20a]. Restore [Bou21]. Restoring [Mos22]. Rests [Fis24]. Results [Dol23a, Der22c, Per20e]. resurfaced [Cha20a]. Resurgence [Ros23b]. retackles [Koz24d]. Retinal [SH22b]. retires [Zor23b]. Retro [Cas22a, Sch23b]. retrospective [Smi20f]. Return [Koz23d, Kum20]. Returns [Ros21, TGLM21]. Reusable [Bro22a]. Revamping [Jon22d]. Revealed [Ack23a]. Revealing [Cho21]. revenue [Der22c]. reverse [Cas20h]. review [Raf20a]. Revisiting [BCM^{+}19]. Revives [Cly24]. Revolution [Usk21, Fun20]. Revolutionize [AZPC22]. revolutionized [Per20o]. revolutions [CFPP22]. rewards [Ano20-93]. RF [Zor20]. Rice [Mar24f]. Rich [CP21, Jon22e, Sch22c]. Richard [Zor24a]. Riddled [Bau24]. ride [Per20l, Pot20a]. Right [Smi22l, Hut20, Ong20, Zor22d]. Ripples [Zor22c]. RISC [Gen24b, Szk21]. RISC-V [Gen24b, Szk21]. rise [Ras20b]. Rises [Per21c]. rising [Smi22i]. risk [Cha20a]. Risks [Mar23k]. Risky [Ros23a]. Rival [Shi22, Fou22, Pal22]. Rivers [Jon22g]. Road [Bro22a, Dum22, HF22, Mur23, And20g, Goo23a, MMP22]. roads [Ack21a]. Roars [Fai24a]. Roasted [Ano22-57]. Robocalls [Koz24a]. Robocars [RHBM22]. RoboCup [Gol23c, Jon22f]. Robot [Ack21a, Ack22b, Ack23b, Ack24c, Bro23c, Chi21, Coh22, GK24, Mar21a, AG20, Fru20, GA20, Gui20, HP20, Lev20]. Robotic [Bea22, Jon24c, Usk21]. roboticist [Gen23b]. Roboticists [FBMI23, Gol23g, Sto23]. Robotics [Ack24a, Bea22, Bar20]. Robots [Ack21b, Ack22a, Ack22c, Ack22d, Ack23a, Ack23b, Ano21-144, Ano21-145, Ano22-131, Ano22-132, Gol23c, Gol23g, GK24, LH24, Pan20, Sto23, WPBS23, Ack24d, Ano20-68, Ano20-91, Ano20-103, Ano20-104, Bar20, Gol23c]. Robust [Kuh24]. rock [Gal20a]. Rocket [Bro22a, Per20f]. rockets [Fou22]. Rogue [Pow24]. role [Cha20a]. Roll [RZ21, Sch23c]. Rollout [Smi22j]. rollovers [Per22a]. Roomba [Jon22g, Mar20g]. rooms [Ano20-68]. Roots [Gen24b]. Rose [Cas24a]. Rosedale [Gen22]. Roses [Str22b]. Ross [Zor23b]. Round [Mar21d]. Routers [Mit20]. Rover [Ulr22, Pot20a]. Royal [Gen23c]. rules [AFH22, Smi20g]. Run [Svr23, Cha20b, Gen24h]. Running [Car24]. Runs [Jon22c, Reg23]. Russia [Cla24]. Russian [Usk21]. Ruzena [Ack24a]. Ryckaert [Moo24b].

S [Gol23f, Gol24c]. S. [Der21g]. Sackbut [Mar24a]. Safe [ACDG21]. safeguard [RS23, SMD20]. Safety [Der21h]. salad [Gui20]. salaries [Per21c]. salary [Per22c]. sale [SG22]. salivate [Koz21a]. salmon [Nor20b]. Sam [Cas20b, Moo23a]. Same [Hec20, Bra21]. Samsung [Moo22b, Pal22]. Sandia [Der21d]. Sandoval [Der22d]. Sandoval-Pichardo [Der22d]. Sapphire [Lev21]. Satellite [Jag22, Koz21a, Zor22e, Lau23b, Mar20h, Raf20b]. Satellite-Imaging [Jag22]. Satellites [BD22, Koz20a, Mit20, Red20]. Satiate [Gol24h]. Sausage [Zor24a]. Savant [Has20b]. Save [Bow21, BSAH21, Car24, CPBZ21, GP22, Moo20b, Ong20, Per20c]. Saving [Moo24d]. Savings [Smi21j]. Saw [GOL24d]. Say [ACDG21, AS22, Gal21b]. Says [Str22a, DM20]. Scale [Der21h, Gol23h, TB21, CFPP22, VH20]. Scaling [Gol24h]. Scan [Gen23e]. Scandal [Pau21]. scanner [GAG20]. Scanning [AS22, Pri21]. Scenarios [BSG22]. Schematics [Bro23c]. scheme [CD21]. Schneider [Gol23e]. Schrödinger [Ros22e]. Sci [Gol24j, Mie24, Gol24a]. Sci-Fi [Gol24j, Mie24, Gol24a]. Science [Der21g, Gen24c, JPWB21, MJH^{+}21, SAS^{+}21, Zim21, CKG^{+}21]. Scientific [CP21, Dum20, Cas20h]. Scientist [Gol23e, Per20o]. scientists [NS20]. Scientology [Mar24i]. Scope [Har22]. Scotland [Fai24a]. Scott [Rao24]. SCR [Bre20]. SCR-536 [Bre20]. Scramble [Baj24, Per20e, Ulr20c]. scrambles [Fai20a]. scrap [Smi22g]. Scratch [Moo24a]. screen [Som21]. Screens [Poo24b]. SDR [Sch22b, Ver24]. Sea [Cla24, MD23, Sch22c]. Seabed [Zor22d]. Search [MMM24, Sch22d]. searches [Ano20-86]. Second [Cas22c, Gen22, SH22b, RHBM22, Sch22e]. secret [LA22]. Secrets [Gol23e, Tei23, Der21d, RS23]. sector [Gol24e]. secure [Pes20b, SMD20]. Secures [Gen24b]. Security [Coo22, Gen24b, Lau24b, RS23, Lew20, Ras20b]. See [Cla24, Dol23a, Gol23d, Kuh24, Poo24b, RHBM22, PN20, Pes21a, Sch21b]. See-Through [Poo24b]. Seeing [Baj24, Tei23, GAG20]. Seek [Baj24]. seeks [Raf20b]. Seen [Sch21b]. segregation [Moo20c]. Seismometer [Sch23a]. Self [Bra21, Mar23k, WPBS23, And20g, LG20, Pan20, Smi23a]. Self-Driving [Mar23k, WPBS23, And20g, LG20, Pan20, Smi23a]. Self-interference [Bra21]. Selfie [Ano21-136]. Semiconductor [Gob23, Jes21, Kim22, Moo24e, NYI24, Pat23a]. Semiconductors [Moo24e]. Sensation [Bou21]. Sense [Ano20-94, Pat20]. Senses [KK21, Dut22, Pes20f]. sensing [CFPP22]. sensitive [Har22, SMD20]. Sensor [Ano21-138, Ano21-137, CP21, CFPP22, Der21e, Per22a, Rab22, Ano20-91]. Sensor-Rich [CP21]. sensoribility [Ano20-94]. Sensors [Baj24, CFPP22, Fit21, WPBS23, Ano20-79, Gen24h, Lee20, PK20, Per21a, WB21]. sequestering [Pat24a]. Serious [Gol23a, Fai23c]. Service [Lau23b]. Set [Der21e, Gen24c, Has20b]. Seth [Koz24b]. Seven [Ano21-116]. several [RHBM22]. Severestorms [FOSMD23]. Sews [Jor21]. Shakes [Sch23a]. Shakespearean [LCBH20]. Shane [Ano23a]. Shaped [PRR23]. Shaping [Zor22c]. shares [Goo23a]. sharing [Per20l, PSD22]. Sheer [Abb23]. Shelf [Cly24]. Shelves [Nor20b]. Shield [Fai24b, Gol24b]. Shift [Gen24c, And20c]. shifts [And20g]. ship [And20a]. Shipping [Ano21-139, Pat24b]. ships [Gal21a, Pat24b]. Shipt [Cal24]. Shoebox [HH24]. Short [Bea22, Fru20]. shortage [Per20e]. Shortcut [BBFS22]. Shot [Jon22h, Ros23b, Wal20a]. Should [Cas22c, Pes21b, Ros23b, LG20, Per20h, Str24a]. Shouldn’t [Fit21]. show [Pat20, Pat22, Str20b]. Shows [Moo24e, WB21]. Shrink [HH24, Rak24b].

Side [Bro22d]. Sight [SH22b, RHBM22]. sights [Has20b]. Signals [Bou21, Har21]. significantly [Fun20]. Silicon [BVKR23, Gen24m, MZR21, Fai24b, Gol24b, Per20m, She24]. Simple [Cas23c, AG20, Sch20b]. simpler [Cly22]. simplified [Per22b]. simplify [Cas20f]. Simulation [Ano21-124, Ano21-140, Ano21-141, Ano20t, Ano21-134]. Simulator [Smi23b]. Sinclair [Cas20h]. sing [Cas20b]. Singh [Der23a]. Singhania [Gen24d]. single [Cas20a]. sirens [Ros20b]. Sister [MZR21]. six [Scu20b]. Skeptic [Bar24]. Skeptical [Cha23]. sketch [Per23]. Skies [JAR24, Ros22a]. Skill [Der21e]. Skills [Ano20-95, Bea22, Gol22b]. Skin [Som21, ZL20b]. Skinny [Smi21d]. Sky [Bar24, Sch23c, Zor22e]. Skydiver [Jor21]. Slam [Ros22c]. slash [Gal21a, Gal21b]. Sleuths [Ano21-133]. Slow [Smi22j, Str24a, LG20]. Slowed [Smi21k]. SLS [Fou22]. sludge [Gal20b]. Small [Cha21b, Deg24, Koz21a, Str22a, Zor22b, Koz20a, LG20, Nor20a, Pat24b, Per20f]. smaller [RK22]. Smallest [Der22d]. Smart [Bor24, Fit21, Hsu20c, Ano20-96, Ano20-97, GA20, Ros20b]. Smart-Home [Bor24]. Smartly [CP21]. Smartphone [Sla21, Som21]. Smartphones [Ano20-96, Ano20-97, ZL20b]. smashes [Moo20e]. Smashing [SC23]. Smell [Str22b]. Smelling [Fit21]. Smoke [Fit21]. Smoothly [Gen24f]. SMS [Tho21a]. Snapshots [PK20]. Soar [NG23]. Socially [Ano21-142]. Software [Cab24, Hub24, Hus24, Kas23, Pes21b, VA23, Ano21-29, Ano21-30, Bar20, Cas22b, Cha20b, Hsu20a, Lee20]. Software-Defined [Kas23, Cas22b, Lee20]. Solar [Ano21-124, Fai21, Gol24a, Jon23f, Pat24a, SC23, Sch22a, WA24, Fai20b, JJS22, LSM23]. Solar-powered [Pat24a]. Solution [Gal21a, Pou22, Zor22d]. solutions [Has20a]. Solve [Jon23i, NG23, DM20, Gol22a, Per20n, Sch21a]. Some [Ake23, Der21g, Smi21k, ZL20b, Ano20-87]. someday [Pat20]. someone [Ano21-27, Ano22-55]. Somerton [Abb23]. Soon [Bro21, Fou21, HS23]. Soonish [Zor23c]. sort [Cal22]. Sound [Mar24l, Arm20, LDO22, Ros21]. Sounds [Lew23, Ros20b]. Source [Gen24b, Ver24, Smi22g]. South [Kim22, LSM23]. Space [AF23, Bar24, CP21, Coo21, Der21f, Fou22, Gol24a, Jon23j, Sch21b, Sch22a, BMRW20, Jon24a, Mit20, Pri21]. Space-based [Gol24a]. Spacecraft [CP21, Der21e, Zor22c, Red20]. Spacewalk [Pot24]. Spafford [Der23c]. spawned [Cro20]. Spawning [Pat23a]. speare [LCBH20]. Special [Ano21-143, Mar20h]. Spectacular [Smi22m]. Spectators [Bea22]. Spectral [Cha20a, Has20a, Has20b, Luc20a, Per20b, Per20l, Per20n, Per20e, Per20h, Per20o, Pre20]. Spectrometer [Coo22]. Spectroscopy [Ros22b]. Spectrum [Laz22, Bar20, Ano20s, Cas22a, Gol21, HPK22, SM24, Zor23b, Zor24b]. Speed [Ack23c, Moo23b, Moo24e, Sch21a, Cal22, Moo20e, Ros20a]. Speeds [Gen24d, Hec20]. Spent [WA24, Kum21]. Spidey [Pes20f]. spies [Luc20b]. spin [Lau24b]. spin-offs [Lau24b]. Spinning [Gen24f]. sponsored [Sto20]. Spoof [Har21]. Spooky [Jaf22]. Spot [Cas22b, GA20, MJS20, Per20a]. spray [And20b]. Spraying [Gen24e]. spread [Smi22j]. sprung [Gen24b]. Spy [Mar21a, Pau21]. SQL [Ano20-95, Cas22c]. Square [Per22b]. squares [Nor20a]. Squash [Sca22]. Squeezing [Gen24c]. St. [Koz21a]. Staff [Ano23-135]. Staggering [Cas22g, Gol23h]. standard [Ros22c]. stands [Koz21a]. star [Gol23d, Per20g]. Start [Gen24a, Moo24a, Ros22a, LG20]. Started [Gol24a]. Starts [Eis23]. Startup [Der22g, Moo24e, WA24, Ano20-100, DM20, Gen24k, Ros22d, Str23]. Startups [AZPC22, Cly20, Moo21]. State [Fra21, MS22, Sto20]. state-sponsored [Sto20]. Stay [Per20g]. Stay-at-home [Per20g]. Steal [Baj22]. Stealth [Baj22, Baj24]. Steel [NYI24]. Steep [Smi22c]. Stefany [Cas22d]. Stellarator [Cly24]. Step [Zor22b, Raf20b]. Still [Gol24a, Gol24i, Hub24, Smi21d, Smi22a, Per20b, Smi22g, Smi22k, Smi23b]. still-mighty [Smi22g]. Stimulation [Dol23a, Dol23b]. Stirs [Sch22c]. stood [Arm20]. Stop [Koz24a]. Storage [Car24, Moo21, Dum22, Nor20a, Pat20]. Store [RZ21, Scu20a]. Storing [Nor20a]. Storm [Gol24b]. Story [Ano20g, Ano20r, Ano20q, Ano20-69, Ano20-47, Ano20-48, Ano20-81, Ano20-92, Ano20s, Ano20-80, Ano20-102, Ano21-49, Ano21-50, Ano21-47, Ano21-48, Ano21-62, Aro24, Smi21l]. Storyteller [Gol24j]. straddles [CD21]. Strange [GO22, Mul24, Smi21m, Fru20]. strategies [Fou21, Pat22]. Strategists [Cla24]. Strategy [Gol24b, Koz20c]. street [Ros20b, Wea20]. strength [And20e]. stretch [ZL20b, Ack22b]. stretchable [Som21]. Strikes [Cas23c, Cha20a]. Structures [BF24, CP21]. students [Bar20, Sac20]. Studio [Smi21l]. Study [Ano21-140, Ano21-141, GO22]. stuff [Cas22b]. Stunning [Gol24g]. stutters [Per20j]. Submarine [Baj22, Baj24]. Subspace [Tow21]. substantial [Jaf22]. Subterranean [Ack22c]. Success [Bro22a, Goo23a]. Sucking [Gal21b]. Sucks [Hus24]. Suddenly [Cal24]. Sun [Ano22-57]. sunk [LSM23]. Sunless [Pow24]. Sunny [Ano21-55, Tow21]. Superbatteries [Smi22k]. Supercharged [Jes21]. Supercharges [SC23]. Supercold [Lev21]. Supercomputer [MCG24, Sch22e]. Supercomputing [Dum23, And20i]. superconducting [Cly23]. Superconductors [HH24]. superheroes [Gen23b]. Superhuman [Ack23c, Cal22].

Superior [Wat24]. Superpower [Kim22]. Supersonic [Ros21, Sch21e]. supplies [Ano20-68]. Supply [AFH22, Pes21b, Fai20a, LSM23, RG20, Sch20c]. sure [Hig20e]. surfaces [Pou22]. Surgeons [Bea22]. Surgery [Bea22]. Surgical [Bea22]. Suri [Gen24h]. Surpass [Bro21]. Surprises [Cas23c, Sca22]. Surprisingly [Cho21, Sar20]. Surveillance [Lau24b, Pes21a]. survey [Per22c]. Survivors [Str22b]. Susana [Gen24j]. Sushama [Gen23d]. Sustain [Smi21e]. SUV [Per20l, Per22a, Smi21j]. SUVs [Smi21j]. swabbed [Ano20-68]. Swallowed [CG23]. Swarm [Jon24c]. sweat [MW20]. Swimmer [Mar21a]. Swiss [Lau24a]. Switch [SG24, Jes21, Zor20]. switches [Zor20]. Synchron [Wal24b]. System [Ano21-124, Fou22, Kas23, Ros22a, Moo23b, NG23, PSD22]. System-Level [Kas23]. systemic [Moo20c]. Systems [CD21, Rao24, Ano20g, Ano20-90, Cal22, Cha20b, CHM21, Der21d, Hsu20a, Pat24a, Has20b].

T [Sch20b]. T-coil-equipped [Sch20b]. Table [Ano20-98, Ano21-113]. tackle [Per20e]. Tackling [Chi21]. Tactics [Cla24]. Tags [TB21]. Taiwan [Cla24, Fai24b, Gol24b]. Take [Ano22-56, Bar24, Bor24, Cho23, Gen24a, Gol24h, Jon23b, MZR21, Per20h, Sch21e, Sch24b, WAC^{+}20, Ano20-80, DM20, Fun20, Jag22]. takeoff [DM20]. Takes [Cal22, Sto20, Str20a, Wal20a]. Taking [Mar23j, RK22]. Talkie [Bre20]. Talking [Per24a, CD21]. Tan [Str23]. Tandy [Cas22f]. Tanghal [Der23b]. tanks [Gal20b]. Tap [Fai23a]. tape [Cly23]. Tapped [FBMI23]. Tardigrade [Ros22e]. tasks [AG20]. Taste [Ano24-145]. tattoos [Scu20a]. taught [ALRW21, SZ20]. Taxi [AZPC22]. Taxis [NG23]. Taylor [Koz23b]. teach [Wea20]. teacher [Bar20]. Team [Ack21b, Bou20]. Teams [Per24a]. Tech [Ano21-114, Ano21-143, Ano22-129, Ano23-136, Ano24-146, Bor24, Cas24d, Cha23, Cho23, Dum22, Fis24, Gen24c, Gol24c, Lau24a, Moo24d, MD23, Per21c, RZ21, Shi22, SM24, Szk21, Wat24, You22, And20h, CFPP22, Der22h, Gar20, Has20a, HS23, Koz24d, McC20, Ong20, Per20a, Per20b, Per20m, Per20d, Per20g, Per20i, Raf20b, Rak24a, Ras20b, Sac20, Smi23a, SZ20, Ulr20a, Ulr20c, YS20]. Technical [Gen23a]. Technicolor [HF22]. Technologies [GOL24d, Gol23d, LG20, Pal22]. technologists [BS22]. Technology [Der23d, JPWB21, Jon23a, LW24, MJH^{+}21, SAS^{+}21, Ano20-100, CKG^{+}21, Gol22d, Gol22e, Gol22f, Hsu20b, Lee20, Smi22i]. technophobes [Per20g]. teenager [Sha23]. Tekla [Gol23f, Gol24c]. telecom [Koz21a]. Telepresence [Ack23a]. telescope [Bou20, HPK22, Pot24, Sch21b]. television [Mie22]. Tell [Fra21, Mar23k, Hut21]. Tells [Gol22f]. temperature [Cly23]. term [And20d]. Terminators [BSG22]. Tesla [Har22, Ulr22]. Tesler [Per20o]. Test [Dum22, Gar20, Koz20c, Koz23d, Per20e, Ros21, Ros22d, YS20]. Testing [YS20, Der22h, Per20a]. Tests [Har21, Der22f, Per20a, YS20]. Texas [Gal21b, NG23]. text [Lau23b]. Textiles [CP21]. Texting [Tho21a]. thank [Per20o]. Their [Cal24, Gal21a, Gol23d, NS20, Per20b]. Them [BVKR23, Gol23g, Mar23j, Pow24, Ack23b, And20h, Sch22b]. Themselves [Ros22c, Ack21a, Koz20b]. theory [SHD20, ZL20a]. therapy [HVW20]. There [Pou22, Pes20a, Str21, PK20]. thermal [Zel20]. thermodynamics [AF23]. These [Arm20, WC23]. they’ll [Ano20-68]. they’re [Hsu20a]. thick [Rad20]. Thin [Gob23]. thing [LCBH20]. Things [BVKR23, Pes21b, Str24a]. Thinks [Cha23]. Third [FBMI23, Per20c, Shi22]. Thomas [Der21d]. thousand [Cas20b]. Thousands [Smi21e, Jes21, Kum21]. threatened [Sha23]. threats [Der23c]. Three [Fou21, Moo20b]. Threshold [BS22]. throats [Ano20-68]. Thrones [Smi24]. throughout [Moo20c]. Throw [MW20]. thrust [Rae20]. Thwart [TB21]. ticket [Per20j]. Ticks [Lev21]. Time [Bow21, Gol23a, Sch23c, Ano20-68, Per20m, Pes20e, Zor22d, Dum20]. Time-Lapse [Sch23c]. Timekeeper [Lev21]. Timeline [Cas22e]. Times [Cor21b]. Timing [Wat24, Raf20a]. Tinier [Van23]. Tiny [Gob23, NYI24, WPBS23, MS22, Per22a]. Tips [Ash23, Per24a]. tissue [SMD20]. Today [Bea22, MSS21, PRR23, Rao24, Ano20-78, Per20o]. together [Bou20, Pal22]. Tokyo [Mar23j]. toll [SB22a]. Tomorrow [Bea22, Rao24, Hut20]. Tones [Tho21b]. Tonnnes [Fis24]. tons [Kum21]. Too [Hut20, Smi21d]. Took [HF22, Per20c]. Tool [Cor21b, Cro20, Gen24k, Hsu20a, Tem20]. Tools [RZ21, VA23, Gol22b]. Top [Ano21-143, Ano22-129, Ano23-136, Ano24-146, Cas21d, CG23, GEN24g, Ulr20c, Ano20-95, Cas20i]. top-again [Cas20i]. topological [Cho20]. Touch [Gan24]. Touchscreen [Mie24]. Tough [Smi21n]. tour [WAC^{+}20]. toured [Fru20]. town [PK20]. Toyota [Ack22a]. Tracing [Ano21-124, Hsu20b, Pes21a]. Track [Gen24l]. Tracker [Per24b, Sla21]. Tracking [Rab22]. Tracks [SG24]. trade [LA22]. traditional [Gal20a]. Traffic [Tow21, Der22b, NG23, Raf20a]. Trail [Gol24c]. Train [Chi21]. Trainees [Bea22]. Training [Bea22]. transactions [Per22b]. Transform [Ano20-73, Ano21-122, Gob23, Ano20-72, Cly23]. transformation [Fai24a]. Transformed [Bro23a]. Transformer [Fai23c]. Transforming [Ano21-56]. Transgenic [Nor20b]. Transistor [Ano22-130, Cas22e, LW24, Mar24g, MS22, Moo22c, Zor22a]. Transistors [Gol22c, Van23, Per20c, RK22, Rad20, Rak24a]. Transition [Cha23, Gol23h, Moo24e, Gol22f, JJS22]. transmission [Fai24a]. transmit [Bra21, Sch20b]. Transparent [Baj22, Smi22b, HVW20]. transport [MMP22, Ros22d]. Transportation [ACDG21, AZPC22, Der22a, Gol24e, Per20l]. Trap [ROK24, Pes20c]. Trashed [WA24]. Travel [Ros21]. Travels [Gol24f]. treasure [Sch22d]. treated [Arm20]. treating [Ano20-68]. treatment [And20b, And20e]. trees [PN20]. Trend [Ulr21]. Trick [Van23]. tricky [Cas22f]. tries [BBS23]. triggering [Fai21]. Trillion [LW24, Pow24]. Trip [Smi21m]. Trippe [Bro22b]. trips [WAC^{+}20]. Triumphs [Ano20-99]. trojan [LA22]. Troynikov [Koz23a]. Trucks [Ack21a]. truly [LDO22]. try [Pri21]. Tube [Shi22, Ano20-81, Cas20e]. Tummala [Moo24a]. tunes [And20f]. tuning [Per22a]. tunnels [Cas22e]. Turbine [Jon23k]. Turbocharge [Boy24]. Turbulent [Str21]. Turf [Laz22]. Turing [Cas21c, Reg23]. Turn [Smi22h, Der22c, Gol23d, Per20g, SMD20]. Turned [Smi21l]. Turning [Ano20-100, Bar20, Tho21b, Ano20-99]. Turns [Bea22, SM24, Rae20]. tutor [BBS23]. TV [Cas22b, Mar24h, Mie22, Pal22, Poo24b]. TVs [Poo24b]. tweet [Cas20a]. Twilight [Pau21]. twist [Mie22]. Two [Ake23, And20d, CD21, GP22, HF22, Mar24l, Per20a, SB22a, Zor20]. type [Mie22]. types [Dol23b].

U.S. [Ack21b, And20d, Gol24b, Moo20c, Moo24e, Nor20b, Ong20, Pat23a, Per22c, Raf20b]. ubiquitous [MS22]. UCDAVIS [Ano20-101]. UI [Gen23d]. Ukraine [Cla24, Bau24]. Ultimate [Cas22e, Ver24, Fou22]. Ultium [Ulr20b]. Ultra [Deg24, BSG22]. ultra-deepfakes [BSG22]. ultrafast [Koz20c]. Ultrahigh [Ver24]. Ultralight [Cri20, Zor23c]. Ultrasound [Deg24]. ultraviolet [And20h]. Unappreciated [Cab24]. Uncertain [Str21]. Uncovering [GH23]. underappreciated [RG20]. Underground [Ack22c, Ack22d]. Undermine [Baj22]. undermined [Moo20c]. understand [HVW20, Pre20, Wea20, SG24]. unfairly [Smi22g]. unfortunately [Smi22i]. unhackable [Moo24c]. Unicorn [She24]. Unilever [Gen24i]. Universal [Gen23e, VH20]. Universe [Gen22, HPK22, Jon23i]. universities [Cly20]. University [Ano20-101, Pat23a, Der22c]. unleash [Gol22b]. Unlikely [Mar24f]. Unsustainable [TGLM21]. until [Cha20b, Fru20]. Unused [FBMI23]. Unveil [Gol23e]. unveils [Per20l]. up-and-down [Rae20]. Upcycling [Cas22f]. Upgrades [Pot24]. Upon [Sch22e]. Ups [Moo21]. Upside [Bro22c]. Upstarts [Smi24]. uranium [Mar20e]. urban [Per20l]. USA [Per22c]. Use [SHP24a, Sch23c, Smi21j, HH24, Pou22, Ras20b, Szk21, YS20]. use-anywhere [YS20]. used [Hsu20a]. Useful [MCG24]. user [Der20, Gol22e, Per20o]. users [Hsu20a]. uses [Cho20, Zor20]. Usher [Wal21, Ros21]. Using [LH24, Pre20, ROK24, RHBM22, Cas20e, Dum22, Koz23b, Per20f, Per20e]. usual [Pes20d]. Utrecht [Dum22]. UV [And20h, Cas24e]. UV-K5 [Cas24e].

V [Gen24b, Szk21]. V2G [Gol22f].

Vaccinate [Gib21]. vaccine

[Scu20a, Wal20a]. Vaccines [Wal21, Wal22]. Vacuum [Shi22, Arm20]. Valley [Gol24c, Per20m, She24]. Valuable [WA24, Tem20]. value [Fou22, Per20h]. Vanns [Gen23a]. Variants [Wal22]. variety [Pat22]. Varkey [Sac20]. Vast [Der22d, Fai20b]. Vastly [Wat24]. VC [Der22a, Der23b]. vector [Cas20e]. vector-graphics [Cas20e]. Vegan [Smi22h]. Vehicle [Dum22, Eis23, Gol22f, MMP22, Ong20]. Vehicle-to-Grid [Dum22, Eis23, Gol22f]. Vehicles [Der23a, Ros23a, Smi21j, And20c, Ano20i, LG20, Ong20, Red20, Ulr20b, Wea20]. Venus [Ano24-145, MZR21]. verse [LCBH20]. Versions [Bow21]. vertical [DM20]. vertical-takeoff [DM20]. very [GA20]. vessel [Wal24b]. VHF [Tho21a]. via [Lau23b, Wal24b]. vicious [And20h]. Video [Gen23d, Lau24b, Smi23b]. videoconferencing [Sch20d]. videos [Hec20]. vie [Ack24d]. View [Ano24-144, Jag22]. Vince [Koz24c]. Vint [Per23]. vintage [Sch23b]. Violence [Moo20c]. violin [Ano20-50]. virtual [Ano20-102]. virtually [Pes20a]. viruses [And20h]. Vision [Ano22-108, Aro24, SC23, Cal22, Ros22d]. Visually [Gol24g]. Vital [GH23]. vodka [Ano20-100]. Voiceprinting [Smi23c]. Voices [Smi23c]. Voltage [Bow21, Fai24a]. volts [Jes21]. Volume [Lew23]. votes [Hig20e]. voyage [And20a]. VR [Per20j, WAC^{+}20]. vs [Hig20g]. Vulnerability [Hub24, RG20]. vying [Cly20].

Wage [Per20i]. Waiting [Har21, Smi22k]. Waits [Koz21a]. Wallet [Per22b, ZL20b]. Wanamaker [Bar20]. Wandering [Pow24]. Want [FBMI23, SG24, Per20h]. Wanted [Jon22b, Raf20b, Ano20-95, Cas20g]. wants [AG20, Ack21b, Per22b]. War [Smi22f, Arm20, Cla24]. Warehouse [Ack22b]. warn [Hsu20c]. Wars [Laz22, Mis23]. Was [Hoc22, Bro22b, Cas20h, Fru20, Smi23b]. Washing [Pre20]. waste [Kum21]. Watch [And20f, Ano20-91, Ham20d, Ham20b, Hsu20c, Wal20b]. Watchdog [Fit21]. Watts [Pat23b]. Wave [AZPC22, Koz20c]. waves [Gal20a]. Way [Cab24, Kuh24, MMM24, MCG24, Sca22, Koz21a, Sch20b, Str21]. Ways [BVKR23, Cho21, Fou21, Moo20a, RG20]. Weaning [JJS22]. Weapons [Pat23b]. Weapons-Grade [Pat23b]. wear [Sch20b, ZL20b]. Wearable [KK21, Mar24j, MW20]. Wearables [Bou21, Per21a]. wearing [Pre20]. Weather [Sch22a, Sch22b, Tow21]. Weave [Cas22g]. Weaving [Mar20i]. Web [Pau22]. Web-Enabled [Pau22]. Web3 [Koz23c]. Webb [Sch21b, HPK22]. Website [Gol21]. weeks [Ano20p]. Weight [GP22]. Weird [Ano23a, Cas22f, MS22]. We’ll [Gol24e, FOSMD23, Ros23b, Pes21a]. Well-Connnected [FOSMD23]. wells [Gal20a]. We’re [Mar23j, Pot24, Bro22a, Mul24, Raf20b]. west [Ano20f, Gal21b]. We’ve [GOL24d]. Where [BMRW20, GK24, Jon23k, Sch21b, Per20h]. wherever [Per20h]. Whether [Fra21]. while [Gol24b, Ras20b, Sch20d]. Whirlpool [Gen24f]. white [Per20b, Koz23c]. Who [Ack23b, BSG22, Cal24, Gol23g, Gol24j, HP20, Jor21, Gen24j, Per20o, Sch20b]. Wi [Koz24d, Hig20j, Zel20]. Wi-Fi [Koz24d, Hig20j, Zel20]. Wide [Ack24c, Mis23, Smi21e]. Wide-Bandgap [Mis23]. Wide-Body [Smi21e]. Wiggling [Boy24]. wild [Ano20f]. Wildfire [Fit21]. Will [And20b, And20i, Baj22, Cho23, CPBZ21, Dol23a, Fou22, Gol24h, Mis23, Moo24e, Smi23a, Ack20, Ack21a, AS22, BS22, Bow21, BSAH21, Cly22, Dum20, Dum23, Fou21, Fun20, GOL24d, Gen24a, GO22, GA20, Has20a, Hut20, Koz23d, KH24, Lau23b, MW20, Mit20, Moo22b, Moo22c, MD23, Pal22, Pan20, Pat22, Pot20a, Ros22c, Ros22d, Sch21b, Sch22c, Smi22h, Sto20, VH20, WPBS23]. Wimshurst [Mar23c]. Win [Gen24b, Mis23, SG22, Pot20b, ZL20a].

Wind [Ano23-137, Jon23k, MD23, JJS22]. Wind-Powered [Ano23-137]. Wind-to-Hydrogen [MD23]. windfall [Koz24c]. Winding [Mur23]. Windows [Bro23a, Ros20b]. Winds [Fai23a]. wine [Smi20e]. Wingman [Lev20]. Winners [Sto23]. winning [SZ20]. Wins [Gen23a]. Winter [Pat22]. WinZips [Per21b]. Wireless [BD22, Ulr20a, Koz21b, Koz24d, SMD20]. Wiretapper [Hoc22]. wish [Ano20-95]. Without [Gol22e, Lew23, Ros22d, WC23, And20h]. Withstand [FOSMD23, Cha21a]. Wizard [Zor23b]. Wizards [Ano21-50]. Wombs [Zim21]. Women [GEN24g, Gol24c, Per20i, Per20b]. wonder [Smi23b]. Wonderful [Mar21j, Gol23c]. Won’t [Pau22, Bro21]. word [Tho21b]. Work [Moo24e, Ros23b, Cro20, Der21d, Der22f, NS20, Per20h, Pre20]. Workbench [Coo21]. Worked [Zor22a, Smi23b]. Workers [Cal24, AS22, Ano20-68]. Workforce [Pat23a]. Working [Der21g, Pes21b, NG23]. Workplace [AS22, Per24a, Per20d]. Works [MSS21]. World [Ano21-144, Ano22-131, Ano22-132, Car24, CG23, Gen23e, Gib21, GO22, GP22, Lev21, McC20, PRR23, Smi21k, Smi22h, Ano20-103, Ano20-104, Ano21-145, Ano22-130, Cha20b, Dum22, Gol23c, Jon24a, Zor20, Sto23]. Worldcoin [Gen23e]. Worlds [Pow24]. Worst [Ack22b, BSG22]. Worst-Case [BSG22]. would [DM20, PSD22]. Wrapped [CP21]. Write [Cab24, Cas20a]. Writes [Der21g]. Wrong [Ano23a, Hig20i]. WW [Hoc22]. WYSIWYG [Bro23a].

X [Cas20g, LA22, Mar20f]. X-ray [LA22, Mar20f]. X2 [Mar24k]. XB [Ros21]. XB-1 [Ros21]. Xerox [Bro23a]. XPrize [Ack23a, Ano20-100].

Year [Abb23, Ack24d, Cly24, Smi21f, Str24b, Zor22c, Ack21a, Ano20t, Per20j, Per20a, Smi20h]. Years [GOL24d, SB22a, Der22b, Fun20, Jon24a, MS22, Scu20b, Smi20d, ZL20b, Zor22d, Zor23b]. yellow [GA20, Raf20a]. yellow-light [Raf20a]. yields [Sca22]. Yilu [Wal24a]. York [Hoc22]. Yorktown [Gol24g]. You’ll [Cas23b]. Younger [Ack22a]. Yours [Cas22g]. yourself [CD21]. Yugoslavia [Ack23d, Cas23d]. Yung [Der22c]. Yung-Hsiang [Der22c].

Z [Cly22]. Z-pinch [Cly22]. Z80 [Ack23d, Cas23d]. Z80-based [Ack23d, Cas23d]. Zach [Gen24k]. Zap [Cly22]. Zero [Cas23b, Fit21]. Zero-Power [Fit21]. Zittrain [Der23d]. Ziv [Per21b]. Zoom [Hec20, Sch20d]. Zurich [Ano22-51].

References

Abbott:2023:FSM

[Abb23] Derek Abbott. Finding Somerton Man: How DNA, AI facial reconstruction, and sheer grit cracked a 75-year-old cold case. IEEE Spectrum, 60(4):22–30, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2021:THS

[ACDG21] Evan Ackerman, Stephen Cass, Michael Dumiak, and Maria Gallucci. Transportation: How safe are eVTOLs?: Extremely safe say manufacturers: News. IEEE Spectrum, 58(11):6–13, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2020:EWO

[Ack20] E. Ackerman. This exoskeleton will obey your brain. IEEE Spectrum, 57(1):50–51, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2021:RTO

[Ack21a] E. Ackerman. Robot trucks overtake robot cars: This year, trucks will drive themselves on public roads with no one on board. IEEE Spectrum, 58 (1):42–43, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2021:DLG

[Ack21b] Evan Ackerman. Deep learning goes to boot camp: The U.S. Army wants to team humans and robots on the battlefield. IEEE Spectrum, 58(10):56–62, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:HRC

[Ack22a] Evan Ackerman. How robots can help US act and feel younger: Toyota’s Gill Pratt on enhancing independence in old age. IEEE Spectrum, 59(8):26–31, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:RWJ

[Ack22b] Evan Ackerman. A robot for the worst job in the warehouse: Boston Dynamics’ Stretch can move 800 heavy boxes per hour. IEEE Spectrum, 59(1): 50–51, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:RCU

[Ack22c] Evan Ackerman. Robots conquer the underground: What DARPA’s subterranean challenge means for the future of autonomous robots. IEEE Spectrum, 59(5):30–37, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:UR

[Ack22d] Evan Ackerman. Underground with robots. IEEE Spectrum, 59(5):2, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2023:HLW

[Ack23a] Evan Ackerman. Human in the loop: What the avatar XPrize revealed about the future of telepresence robots. IEEE Spectrum, 60(5):38–44, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2023:RHH

[Ack23b] Evan Ackerman. A robot for humanity: How robots can empower people who need them the most. IEEE Spectrum, 60(11): 22–29, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2023:SSH

[Ack23c]  Evan Ackerman.  Superhu-

man speed: How autonomous drones beat the best human racers. IEEE Spectrum, 60(9): 30–37, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2023:YHB

[Ack23d] Evan Ackerman. Yugoslavia’s home-brewed microcomputer: The Z80-based Galaksija dared to be different. IEEE Spectrum, 60(8):21–22, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2024:QRB

[Ack24a] Evan Ackerman. 5 questions Ruzena Bajcsy: a founder of modern robotics reflects on her legacy. IEEE Spectrum, 61 (12):23, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2024:BHA

[Ack24b] Evan Ackerman. A brief history of Aibo. IEEE Spectrum, 61 (12):16–17, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2024:OWR

[Ack24c] Evan Ackerman. Open wide for this robot arm. IEEE Spectrum, 61(10):14–15, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2024:YHL

[Ack24d] Evan Ackerman. Year of the humanoid: Legged robots from eight companies vie for jobs. IEEE Spectrum, 61(01): 44–48, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:Nb

[ACLS22] Evan Ackerman, Charles Q. Choi, Lucas Laursen, and Rebecca Sohn. News. IEEE Spectrum, 59(9):6–13, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Assawaworrarit:2023:CSC

[AF23] Sid Assawaworrarit and Shanhui Fan. The coldness of space can power earth: We can hack thermodynamics to cool buildings and light up the night. IEEE Spectrum, 60(12):40–45, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Almassalkhi:2022:PPG

[AFH22] Mads Almassalkhi, Jeff Frolik, and Paul Hines. Packetizing the power grid: The rules of the Internet can also balance electricity supply and demand. IEEE Spectrum, 59(2):42–47, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2020:RKI

[AG20] E. Ackerman and E. Guizzo. A robot that keeps it simple:

Hello robot wants to reinvent how autonomous machines perform tasks at home. IEEE Spectrum, 57(10):11, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Akesson:2023:CTC

[Ake23] Linus Akesson. The Commodordion: Two C64s plus some floppy disks equals one instrument. IEEE Spectrum, 60(1): 20–22, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Acuto:2021:WWL

[ALRW21] Michele Acuto, Shaun Larcom, Ferdin Rauch, and Tim Willems. What we learned from the pandemic: Most of all, it taught us how to adapt under pressure. IEEE Spectrum, 58(8): 22–27, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:FPC

[AMK^{+}20] M. Anderson, S. K. Moore, M. Koziol, R. N. Charette, and M. Scudellari. Fractional particles could mean new electronics: Quasiparticles might lead to new kinds of quantum computers — [news]. IEEE Spectrum, 57 (11):6–13, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:BVA

[And20a] M. Anderson. Bon voyage for the autonomous ship Mayflower. IEEE Spectrum, 57 (1):36–39, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:EAM

[And20b] M. Anderson. Engineered antibody mist blocks coronavirus: Will the first breakthrough COVID-19 treatment be an inhaler or nasal spray? — [news]. IEEE Spectrum, 57(10):6–7, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:EBS

[And20c] M. Anderson. EV batteries shift into high gear: Advances in anodes, cathodes, and electrolytes are poised to appear in future electric vehicles. IEEE Spectrum, 57(4):8–9, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:LPU

[And20d] M. Anderson. Limited progress for U.S. nuclear: Two fission projects raise questions about the long-term future of nuclear power — [news]. IEEE Spectrum, 57(6):6–7, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:PSG

[And20e] M. Anderson. Prescriptionstrength gaming: ADHD treatment now comes in the form of a first-person racing game — [news]. IEEE Spectrum, 57 (8):9–10, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:PTP

[And20f] M. Anderson. PrintFixer tunes any 3D printer — [journal watch]. IEEE Spectrum, 57(4):7, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:RAS

[And20g] M. Anderson. The road ahead for self-driving cars: The AV industry has had to reset expectations, as it shifts its focus to level 4 autonomy — [news]. IEEE Spectrum, 57(5):8–9, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:UOG

[And20h] M. Anderson. The ultraviolet offense: Germicidal UV lamps destroy vicious viruses. New tech might put them many more places without harming humans. IEEE Spectrum, 57 (10):50–55, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:WCA

[And20i] M. Anderson. Will China attain exascale supercomputing in 2020? IEEE Spectrum, 57 (1):52–53, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Aa

[Ano20a] Anonymous. Advertisements. IEEE Spectrum, 57(8):51, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. Advertisements. IEEE Spectrum, 57(9):52, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ac

[Ano20c] Anonymous. Advertisements. IEEE Spectrum, 57(10):68, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ad

[Ano20d] Anonymous. Advertisements. IEEE Spectrum, 57(10):70– 71, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ae

[Ano20e] Anonymous. Advertisements. IEEE Spectrum, 57(11):50–51, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:AAF

[Ano20f] Anonymous. Applying AI in forestry’s wild west. IEEE Spectrum, 57(8):2, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:ALS

[Ano20g] Anonymous. Attack of the legacy systems — [back story]. IEEE Spectrum, 57(9):2, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:AAA

[Ano20h] Anonymous. Automaton administers antiviral aerosol — [news]. IEEE Spectrum, 57 (5):14–15, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:AVR

[Ano20i] Anonymous. Autonomous vehicles require batteries with lasting power. IEEE Spectrum, 57 (12):5, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BCa

[Ano20j] Anonymous. Back cover. IEEE Spectrum, 57(8):c4, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BCb

[Ano20k] Anonymous. Back cover. IEEE Spectrum, 57(9):c4, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BCc

[Ano20l] Anonymous. [Back cover]. IEEE Spectrum, 57(10):c4, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BCd

[Ano20m] Anonymous. [Back cover]. IEEE Spectrum, 57(11):c4, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BICa

[Ano20n] Anonymous. [Back inside cover]. IEEE Spectrum, 57(10):c3, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BICb

[Ano20o] Anonymous. [Back inside cover]. IEEE Spectrum, 57 (11):c3, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BPA

[Ano20p] Anonymous. Become a published author in 4 to 6 weeks [advertisement]. IEEE Spectrum, 57 (12):15, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BPB

[Ano20q] Anonymous. Biology’s potentials — [back story]. IEEE Spectrum, 57(7):02, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:BEB

[Ano20r] Anonymous. The body electric — [back story]. IEEE Spectrum, 57(12):2, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SBB

[Ano20s] Anonymous. Spectrum’s bellwether — [back story]. IEEE Spectrum, 57(5):2, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:CCN

[Ano20t] Anonymous. COMSOL conference: 2020 North America: The multiphysics simulation event of the year. IEEE Spectrum, 57 (8):17, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:CFC

[Ano20u] Anonymous. Concrete facts — [CrossTalk]. IEEE Spectrum, 57 (3):20–21, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ca

[Ano20v] Anonymous. Contents. IEEE Spectrum, 57(1):1, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cc

[Ano20w] Anonymous. Contents. IEEE Spectrum, 57(2):1, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ce

[Ano20x] Anonymous. Contents. IEEE Spectrum, 57(3):1, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cg

[Ano20y] Anonymous. Contents. IEEE Spectrum, 57(4):1, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ci

[Ano20z] Anonymous. Contents. IEEE Spectrum, 57(5):1, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ck

[Ano20-27] Anonymous. Contents. IEEE Spectrum, 57(6):1, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cn

[Ano20-28] Anonymous. Contents. IEEE Spectrum, 57(8):1, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cp

[Ano20-29] Anonymous. Contents. IEEE Spectrum, 57(9):1, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cr

[Ano20-30] Anonymous. Contents. IEEE Spectrum, 57(10):1, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ct

[Ano20-31] Anonymous. Contents. IEEE Spectrum, 57(11):1, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cb

[Ano20-32] Anonymous. Contributors. IEEE Spectrum, 57(1):3, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cd

[Ano20-33] Anonymous. Contributors. IEEE Spectrum, 57(2):3, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cf

[Ano20-34] Anonymous. Contributors. IEEE Spectrum, 57(3):3, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Ch

[Ano20-35] Anonymous. Contributors. IEEE Spectrum, 57(4):3, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cj

[Ano20-36] Anonymous. Contributors. IEEE Spectrum, 57(5):3, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cl

[Ano20-37] Anonymous. Contributors. IEEE Spectrum, 57(6):3, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cm

[Ano20-38] Anonymous. Contributors. IEEE Spectrum, 57(7):03, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Co

[Ano20-39] Anonymous. Contributors. IEEE Spectrum, 57(8):3, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cq

[Ano20-40] Anonymous. Contributors. IEEE Spectrum, 57(9):3, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cs

[Ano20-41] Anonymous. Contributors. IEEE Spectrum, 57(10):3, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cu

[Ano20-42] Anonymous. Contributors. IEEE Spectrum, 57(11):3, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:Cv

[Ano20-43] Anonymous. Contributors. IEEE Spectrum, 57(12):4, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:CIG

[Ano20-44] Anonymous. CrossTalk: Indeterminate growth. IEEE Spectrum, 57(9):20–21, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:EPC

[Ano20-45] Anonymous. Electricity prices:

a changing bargain: Crosstalk. IEEE Spectrum, 57(2):17–18, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:EBN

[Ano20-46] Anonymous. Emergency broadcast — [news]. IEEE Spectrum, 57(6):12–13, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:EPB

[Ano20-47] Anonymous. The engineer’s perspective — [back story]. IEEE Spectrum, 57(3):2, March 2020.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FRB

[Ano20-48] Anonymous. Final reflections — [back story]. IEEE Spectrum, 57(1):2, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FFF

[Ano20-49] Anonymous. Flying fast, far, and fuelless — [news]. IEEE Spectrum, 57(3):14–15, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:VPC

[Ano20-50] Anonymous. For violin, press command-print — [news]. IEEE Spectrum, 57(4):14–15, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FFL

[Ano20-51] Anonymous. Formula 1 fantasy league — [news]. IEEE Spectrum, 57(8):12–13, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCa

[Ano20-52] Anonymous. [Front cover]. IEEE Spectrum, 57(1):c1, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCb

[Ano20-53] Anonymous. Front cover. IEEE Spectrum, 57(2):c1, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCc

[Ano20-54] Anonymous. Front cover. IEEE Spectrum, 57(3):c1, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCd

[Ano20-55] Anonymous. Front cover. IEEE Spectrum, 57(4):c1, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCe

[Ano20-56] Anonymous. Front cover. IEEE Spectrum, 57(5):c1, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCf

[Ano20-57] Anonymous. Front cover. IEEE Spectrum, 57(6):c1, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCg

[Ano20-58] Anonymous. Front cover. IEEE Spectrum, 57(8):c1, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCh

[Ano20-59] Anonymous. Front cover. IEEE Spectrum, 57(9):c1, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCi

[Ano20-60] Anonymous. [Front cover]. IEEE Spectrum, 57(10):c1, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCj

[Ano20-61] Anonymous. [Front cover]. IEEE Spectrum, 57(11):c1, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FCk

[Ano20-62] Anonymous. [Front cover]. IEEE Spectrum, 57(12):c1, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FICa

[Ano20-63] Anonymous. [Front inside cover]. IEEE Spectrum, 57(10):c2, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FICb

[Ano20-64] Anonymous. [Front inside cover]. IEEE Spectrum, 57 (11):c2, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:FICc

[Ano20-65] Anonymous. [Front inside cover].

IEEE Spectrum, 57(12):c2, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:GGM

[Ano20-66] Anonymous. Good grids make good neighbors. IEEE Spectrum, 57(12):38–43, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:HL

[Ano20-67] Anonymous. The healing light. IEEE Spectrum, 57(9): 50, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:HRB

[Ano20-68] Anonymous. How robots became essential workers: They disinfected hospital rooms. They delivered medical supplies. They swabbed people’s throats. Next time around, they’ll be treating patients. IEEE Spectrum, 57 (10):36–43, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:DCE

[Ano20-69] Anonymous. I drive the car electric — [back story]. IEEE Spectrum, 57(4):2, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IFR

[Ano20-70] Anonymous. IEEE Foundation: Realize: The full potential of IEEE. IEEE Spectrum, 57(12):7, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IMD

[Ano20-71] Anonymous. IEEE member discounts: Exclusive offers are another big plus to IEEE membership. IEEE Spectrum, 57(8):31, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:ITOa

[Ano20-72] Anonymous. Innovate: transform our future. IEEE Spectrum, 57(8):23, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:ITOb

[Ano20-73] Anonymous. Innovate: Transform our future. IEEE Spectrum, 57(9):13, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IBCa

[Ano20-74] Anonymous. Inside back cover. IEEE Spectrum, 57(8):c3, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IBCb

[Ano20-75] Anonymous. Inside back cover. IEEE Spectrum, 57 (9):c3, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IFCa

[Ano20-76] Anonymous. Inside front cover.

IEEE Spectrum, 57(8):c2, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IFCb

[Ano20-77] Anonymous. Inside front cover. IEEE Spectrum, 57 (9):c2, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:IOT

[Ano20-78] Anonymous. Invented in the 1800s. Optimized for today. IEEE Spectrum, 57(11): 5, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:ICF

[Ano20-79] Anonymous. IoT calls for fast communication between sensors. IEEE Spectrum, 57(8):5, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:TBS

[Ano20-80] Anonymous. It’s a take — [back story]. IEEE Spectrum, 57(6):2, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:LTG

[Ano20-81] Anonymous. Life of a tube guy — [back story]. IEEE Spectrum, 57(11):2, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:LLNb

[Ano20-82] Anonymous. Light and lively — [news]. IEEE Spectrum, 57 (12):16–17, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:LLNa

[Ano20-83] Anonymous. Location, location — [news]. IEEE Spectrum, 57 (2):12–13, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:LD

[Ano20-84] Anonymous. Look deeper. IEEE Spectrum, 57(12):3, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:MBE

[Ano20-85] Anonymous. Meanwhile, back on Earth — news. IEEE Spectrum, 57(7):12–13, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:NFS

[Ano20-86] Anonymous. New faculty searches in electrical engineering 2020–21. IEEE Spectrum, 57 (12):51–55, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:OMS

[Ano20-87] Anonymous. Old McDonald had some chips — [news]. IEEE Spectrum, 57(9):14–15, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:PMM

[Ano20-88] Anonymous. Pandemic memories and mortalities — [CrossTalk]. IEEE Spectrum, 57(10):16– 17, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:PFN

[Ano20-89] Anonymous. Photon flight — [news]. IEEE Spectrum, 57(11): 14–15, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:PIA

[Ano20-90] Anonymous. Professor intelligent and autonomous embedded systems. IEEE Spectrum, 57 (9):51, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:PSR

[Ano20-91] Anonymous. A proximity sensor for robots — [journal watch]. IEEE Spectrum, 57(5):10, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:PPB

[Ano20-92] Anonymous. Pursuing the pandemic — [back story]. IEEE Spectrum, 57(10):2, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:RME

[Ano20-93] Anonymous. Recruit a member. Earn rewards! IEEE Spectrum, 57(9):31, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SSC

[Ano20-94] Anonymous. Sense and sensoribility — [crosstalk]. IEEE Spectrum, 57(2):20, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SWS

[Ano20-95] Anonymous. Skills wanted: SQL, Java, Python, and AWS top employers’ wish lists — [careers]. IEEE Spectrum, 57(1):59, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SSHa

[Ano20-96] Anonymous. Smartphones, smart homes, smart healthcare? IEEE Spectrum, 57 (9):5, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SSHb

[Ano20-97] Anonymous. Smartphones, smart homes, smart healthcare? IEEE Spectrum, 57(10):5, October 2020. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:TC

[Ano20-98] Anonymous. Table of contents.

IEEE Spectrum, 57(12):1, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:TTP

[Ano20-99] Anonymous. Triumphs turning points. IEEE Spectrum, 57(1):23, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:TCD

[Ano20-100] Anonymous. Turning carbon dioxide into vodka: a Brooklyn startup is an XPrize finalist for its boozy carbon-capture technology — [podcasts]. IEEE Spectrum, 57(12):21, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:UUC

[Ano20-101] Anonymous. UCDAVIS: University of California, Davis. IEEE Spectrum, 57(10):69, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:WAR

[Ano20-102] Anonymous. When all reality is virtual — [back story]. IEEE Spectrum, 57(2):2, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:WBRa

[Ano20-103] Anonymous. The world’s best robots guide is here! [advertisement]. IEEE Spectrum, 57 (11):37, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:WBRb

[Ano20-104] Anonymous. The world’s best robots guide is here! [advertisement]. IEEE Spectrum, 57 (12):56, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:AGH

[Ano21a] Anonymous. Adrian Gropper: How to become a privacy entrepreneur [careers]. IEEE Spectrum, 58(4):17, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Aa

[Ano21b] Anonymous. Advertisement. IEEE Spectrum, 58(1):1, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ab

[Ano21c] Anonymous. Advertisement. IEEE Spectrum, 58(2):46–47, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ac

[Ano21d] Anonymous. Advertisement. IEEE Spectrum, 58(3):29, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ae

[Ano21e] Anonymous. Advertisement.

IEEE Spectrum, 58(5):31, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Af

[Ano21f] Anonymous. Advertisement.

IEEE Spectrum, 58(5):47, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ag

[Ano21g] Anonymous. Advertisement.

IEEE Spectrum, 58(5):55, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ah

[Ano21h] Anonymous. Advertisement.

IEEE Spectrum, 58(6):3, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ai

[Ano21i]  Anonymous.  Advertisement.

IEEE Spectrum, 58(6):50, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Aj

[Ano21j]  Anonymous.  Advertisement.

IEEE Spectrum, 58(6):55, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ak

[Ano21k] Anonymous. Advertisement. IEEE Spectrum, 58(6):67, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Al

[Ano21l]  Anonymous.  Advertisement.

IEEE Spectrum, 58(8):47, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Am

[Ano21m] Anonymous. Advertisement. IEEE Spectrum, 58(9):3, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:An

[Ano21n] Anonymous. Advertisement. IEEE Spectrum, 58(9):5, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ao

[Ano21o] Anonymous. Advertisement. IEEE Spectrum, 58(9):57, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ap

[Ano21p] Anonymous. Advertisement. IEEE Spectrum, 58(9):69, September 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Aq

[Ano21q] Anonymous. Advertisement.

IEEE Spectrum, 58(9):70–71, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ar

[Ano21r] Anonymous. [Advertisement].

IEEE Spectrum, 58(10):3, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:As

[Ano21s] Anonymous. [Advertisement].

IEEE Spectrum, 58(10):5, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:At

[Ano21t] Anonymous. [Advertisement].

IEEE Spectrum, 58(10):21, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Au

[Ano21u] Anonymous. [Advertisement].

IEEE Spectrum, 58(10):23, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Av

[Ano21v] Anonymous. [Advertisement].

IEEE Spectrum, 58(10):62– 63, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Aw

[Ano21w] Anonymous. Advertisement. IEEE Spectrum, 58(12):3, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ax

[Ano21x] Anonymous. Advertisement. IEEE Spectrum, 58(12):5, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ay

[Ano21y] Anonymous. Advertisement. IEEE Spectrum, 58(12):67, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Az

[Ano21z] Anonymous. Advertisement. IEEE Spectrum, 58(12):68–75, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ABS

[Ano21-27] Anonymous. Advertisement: Because someone looks up to you. IEEE Spectrum, 58(12):57, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ACa

[Ano21-28] Anonymous. Advertisement: COMSOL. IEEE Spectrum, 58(8):3, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ACSa

[Ano21-29] Anonymous. Advertisement:

COMSOL software. IEEE Spectrum, 58(7):5, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ACSb

[Ano21-30] Anonymous. Advertisement:

COMSOL software. IEEE Spectrum, 58(11):5, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:AIF

[Ano21-31] Anonymous. Advertisement:

IEEE Foundation. IEEE Spectrum, 58(4):55, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ad

[Ano21-32] Anonymous. Advertisements.

IEEE Spectrum, 58(4):54, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCa

[Ano21-33] Anonymous. [Back cover]. IEEE Spectrum, 58(4):CV4, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCb

[Ano21-34] Anonymous. [Back cover]. IEEE Spectrum, 58(6):CV4, June 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCc

[Ano21-35] Anonymous. [Back cover]. IEEE Spectrum, 58(8):CV4, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCd

[Ano21-36] Anonymous. [Back cover]. IEEE Spectrum, 58(9):CV4, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCe

[Ano21-37] Anonymous. [Back cover]. IEEE Spectrum, 58(10):CV4, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCf

[Ano21-38] Anonymous. [Back cover]. IEEE Spectrum, 58(11):CV4, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BCg

[Ano21-39] Anonymous. [Back cover]. IEEE Spectrum, 58(12):CV4, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICa

[Ano21-40] Anonymous. [Back inside cover]. IEEE Spectrum, 58(4):CV3, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICb

[Ano21-41] Anonymous. [Back inside cover].

IEEE Spectrum, 58(6):CV3, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICc

[Ano21-42] Anonymous. [Back inside cover].

IEEE Spectrum, 58(8):CV3, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICd

[Ano21-43] Anonymous. [Back inside cover]. IEEE Spectrum, 58(9): CV3, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICe

[Ano21-44] Anonymous. [Back inside cover].

IEEE Spectrum, 58(10):CV3, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICf

[Ano21-45] Anonymous. [Back inside cover]. IEEE Spectrum, 58(11): CV3, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BICg

[Ano21-46] Anonymous. [Back inside cover].

IEEE Spectrum, 58(12):CV3, December 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BSa

[Ano21-47] Anonymous. Back story. IEEE Spectrum, 58(1):3, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BSb

[Ano21-48] Anonymous. Back story. IEEE Spectrum, 58(2):2, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BS

[Ano21-49] Anonymous. Back story. IEEE Spectrum, 58(6):2, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BSW

[Ano21-50] Anonymous. Back story: Wizards of light. IEEE Spectrum, 58(4):2, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPa

[Ano21-51] Anonymous. The big picture. IEEE Spectrum, 58(2): 12–13, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPb

[Ano21-52] Anonymous. The big picture. IEEE Spectrum, 58(4):12–13, April 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPBa

[Ano21-53] Anonymous. The big picture: a big bet on crypto. IEEE Spectrum, 58(8):12–13, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPD

[Ano21-54] Anonymous. The big picture: In the dark. IEEE Spectrum, 58 (11):14–15, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPS

[Ano21-55] Anonymous. The big picture: Sunny days ahead. IEEE Spectrum, 58(9):16–17, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:BPT

[Ano21-56] Anonymous. The big picture: Transforming power research. IEEE Spectrum, 58(12): 12–13, December 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IAAa

[Ano21-57] Anonymous. IEEE Access [advertisement]. IEEE Spectrum, 58(3):43, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IAAb

[Ano21-58] Anonymous. IEEE Access [advertisement]. IEEE Spectrum, 58(4):43, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:PIA

[Ano21-59] Anonymous. Proceedings of the IEEE [advertisement]. IEEE Spectrum, 58(3):52, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cc

[Ano21-60] Anonymous. Careers. IEEE Spectrum, 58(1):15, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:CAa

[Ano21-61] Anonymous. [Classified ads]. IEEE Spectrum, 58(11):50–55, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:CDG

[Ano21-62] Anonymous. Combating doom and gloom [back story]. IEEE Spectrum, 58(7):2, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:CMA

[Ano21-63] Anonymous. COMSOL multiphysics [advertisement]. IEEE Spectrum, 58(4):5, April 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ca

[Ano21-64] Anonymous. Contents. IEEE Spectrum, 58(1):1–2, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cd

[Ano21-65] Anonymous. Contents. IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cf

[Ano21-66] Anonymous. Contents. IEEE Spectrum, 58(3):1, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ch

[Ano21-67] Anonymous. Contents. IEEE Spectrum, 58(4):1, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ck

[Ano21-68] Anonymous. Contents. IEEE Spectrum, 58(5):1, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Co

[Ano21-69] Anonymous. Contents. IEEE Spectrum, 58(6):1, June 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cq

[Ano21-70] Anonymous. Contents. IEEE Spectrum, 58(7):1, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cs

[Ano21-71] Anonymous. Contents. IEEE Spectrum, 58(8):1, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cu

[Ano21-72] Anonymous. Contents. IEEE Spectrum, 58(9):1, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cw

[Ano21-73] Anonymous. Contents. IEEE Spectrum, 58(10):1, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cy

[Ano21-74] Anonymous. Contents. IEEE Spectrum, 58(11):1, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cba

[Ano21-75] Anonymous. Contents. IEEE Spectrum, 58(12):1, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cb

[Ano21-76] Anonymous. Contributors. IEEE Spectrum, 58(1):4, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ce

[Ano21-77] Anonymous. Contributors. IEEE Spectrum, 58(2):11, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cg

[Ano21-78] Anonymous. Contributors. IEEE Spectrum, 58(3):11, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ci

[Ano21-79] Anonymous. Contributors. IEEE Spectrum, 58(4):4, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cl

[Ano21-80] Anonymous. Contributors. IEEE Spectrum, 58(5):3, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cp

[Ano21-81] Anonymous. Contributors. IEEE Spectrum, 58(6):4, June 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cr

[Ano21-82] Anonymous. Contributors. IEEE Spectrum, 58(7):4, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Ct

[Ano21-83] Anonymous. Contributors. IEEE Spectrum, 58(8):4, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cv

[Ano21-84] Anonymous. Contributors. IEEE Spectrum, 58(9):4, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cx

[Ano21-85] Anonymous. Contributors. IEEE Spectrum, 58(10):4, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cz

[Ano21-86] Anonymous. Contributors. IEEE Spectrum, 58(11):4, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cbb

[Ano21-87] Anonymous. Contributors. IEEE Spectrum, 58(12):4, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cj

[Ano21-88] Anonymous. Cover 2. IEEE Spectrum, 58(5):CV2, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cm

[Ano21-89] Anonymous. Cover 3. IEEE Spectrum, 58(5):CV3, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Cn

[Ano21-90] Anonymous. Cover 4. IEEE Spectrum, 58(5):CV4, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:DBD

[Ano21-91] Anonymous. Design better devices faster. IEEE Spectrum, 58(1):1, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCa

[Ano21-92] Anonymous. Front cover. IEEE Spectrum, 58(1):1, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCb

[Ano21-93] Anonymous. Front cover. IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCc

[Ano21-94] Anonymous. [Front cover]. IEEE Spectrum, 58(3):1, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCd

[Ano21-95] Anonymous. [Front cover]. IEEE Spectrum, 58(4):1, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCe

[Ano21-96] Anonymous. Front cover. IEEE Spectrum, 58(5):1, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCf

[Ano21-97] Anonymous. [Front cover]. IEEE Spectrum, 58(6):1, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCg

[Ano21-98] Anonymous. [Front cover]. IEEE Spectrum, 58(7):1, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCh

[Ano21-99] Anonymous. [Front cover]. IEEE Spectrum, 58(8):1, August 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCi

[Ano21-100] Anonymous. [Front cover]. IEEE Spectrum, 58(9):1, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCj

[Ano21-101] Anonymous. [Front cover].

IEEE Spectrum, 58(10):1, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCk

[Ano21-102] Anonymous. [Front cover]. IEEE Spectrum, 58(11):1, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FCl

[Ano21-103] Anonymous. Front cover. IEEE Spectrum, 58(12):1, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICa

[Ano21-104] Anonymous. [Front inside cover].

IEEE Spectrum, 58(4):CV2, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICc

[Ano21-105] Anonymous. [Front inside cover].

IEEE Spectrum, 58(7):CV2, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICd

[Ano21-106] Anonymous. [Front inside cover]. IEEE Spectrum, 58(8):CV2, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICe

[Ano21-107] Anonymous. [Front inside cover]. IEEE Spectrum, 58(9): CV2, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICf

[Ano21-108] Anonymous. [Front inside cover]. IEEE Spectrum, 58(10):CV2, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICg

[Ano21-109] Anonymous. [Front inside cover]. IEEE Spectrum, 58(11): CV2, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICh

[Ano21-110] Anonymous. [Front inside cover]. IEEE Spectrum, 58(12):CV2, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:FICb

[Ano21-111] Anonymous. [Front inside cover - advertisement]. IEEE Spectrum, 58(6):CV2, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:G

[Ano21-112] Anonymous. Gadgeteering.

IEEE Spectrum, 58(10):2, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:GAR

[Ano21-113] Anonymous. The great AI reckoning: Deep learning may have reached its limits. What comes next? [table of contents]. IEEE Spectrum, 58(10):24–25, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:HTE

[Ano21-114] Anonymous. High-tech eyewear [the big picture]. IEEE Spectrum, 58(7):14–15, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IGC

[Ano21-115] Anonymous. IEEE Global Communications Conference, 7–11 December 2021. IEEE Spectrum, 58(11):19, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMO

[Ano21-116] Anonymous. IEEE ME: One program. seven reasons. IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMDa

[Ano21-117] Anonymous. IEEE member discounts. IEEE Spectrum, 58 (1):1, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMDb

[Ano21-118] Anonymous. IEEE member discounts. IEEE Spectrum, 58(5):39, May 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMDd

[Ano21-119] Anonymous. IEEE member discounts. IEEE Spectrum, 58 (11):3, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMDc

[Ano21-120] Anonymous. IEEE member discounts [advertisement]. IEEE Spectrum, 58(7):3, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:IMG

[Ano21-121] Anonymous. IEEE member group disability income insurance. IEEE Spectrum, 58 (7):CV3, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ITO

[Ano21-122] Anonymous. Innovate transform our future. IEEE Spectrum, 58(5):19, May 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:KPL

[Ano21-123] Anonymous. Keep pushing the limits! IEEE Spectrum, 58 (1):1, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:LBO

[Ano21-124] Anonymous. Looking beyond our solar system with ray tracing simulation. IEEE Spectrum, 58(3):3, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:MLa

[Ano21-125] Anonymous. Machine learning.

IEEE Spectrum, 58(1):1, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:MLb

[Ano21-126] Anonymous. Machine learning.

IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:MR

[Ano21-127] Anonymous. A magazine reborn. IEEE Spectrum, 58(5):2, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:MA

[Ano21-128] Anonymous. MATLAB [advertisement]. IEEE Spectrum, 58(7):CV4, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Na

[Ano21-129] Anonymous. News. IEEE Spectrum, 58(1):6–11, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Nb

[Ano21-130] Anonymous. News. IEEE Spectrum, 58(2):4–10, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Nc

[Ano21-131] Anonymous. News. IEEE Spectrum, 58(3):4–10, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:Nd

[Ano21-132] Anonymous. News. IEEE Spectrum, 58(4):6–11, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:PPS

[Ano21-133] Anonymous. Portrait of the portrait sleuths. IEEE Spectrum, 58(9):2, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:ROS

[Ano21-134] Anonymous. Ray optics simulation for inertial navigation.

IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:R

[Ano21-135] Anonymous. Recruitment. IEEE Spectrum, 58(1):55, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:RPS

[Ano21-136] Anonymous. Red Planet selfie [the big picture]. IEEE Spectrum, 58(6):14–15, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SCCb

[Ano21-137] Anonymous. Sensor control and characterization [advertisement]. IEEE Spectrum, 58(3):1, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SCCa

[Ano21-138] Anonymous. Sensor control and characterization: All in one. IEEE Spectrum, 58(2):1, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SN

[Ano21-139] Anonymous. The shipping news.

IEEE Spectrum, 58(3):2, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SCSa

[Ano21-140] Anonymous. Simulation case study. IEEE Spectrum, 58(5):5, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SCSb

[Ano21-141] Anonymous. Simulation case study [advertisement]. IEEE Spectrum, 58(6):5, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:SDC

[Ano21-142] Anonymous. Socially distanced celebration. IEEE Spectrum, 58 (3):12–13, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:TTS

[Ano21-143] Anonymous. Top tech 2021: a special report. IEEE Spectrum, 58(1):21, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:WBRa

[Ano21-144] Anonymous. The world’s best robots guide is here! [advertisement]. IEEE Spectrum, 58(4):3, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:WBRb

[Ano21-145] Anonymous. The world’s best robots guide is here! [advertisement]. IEEE Spectrum, 58 (11):21, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. [Advertisement].

IEEE Spectrum, 59(1):3, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. Advertisement.

IEEE Spectrum, 59(1):5, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ac

[Ano22c] Anonymous. Advertisement.

IEEE Spectrum, 59(1):59, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ad

[Ano22d] Anonymous. Advertisement.

IEEE Spectrum, 59(1):60–63, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ae

[Ano22e] Anonymous. Advertisement.

IEEE Spectrum, 59(2):5, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Af

[Ano22f] Anonymous. Advertisement.

IEEE Spectrum, 59(3):5, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ag

[Ano22g] Anonymous. Advertisement. IEEE Spectrum, 59(4):3, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ah

[Ano22h] Anonymous. Advertisement. IEEE Spectrum, 59(4):51, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ai

[Ano22i]  Anonymous.  Advertisement.

IEEE Spectrum, 59(5):3, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Aj

[Ano22j]  Anonymous.  Advertisement.

IEEE Spectrum, 59(5):5, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ak

[Ano22k] Anonymous. Advertisement. IEEE Spectrum, 59(5):6, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Al

[Ano22l]  Anonymous.  Advertisement.

IEEE Spectrum, 59(5):23, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Am

[Ano22m] Anonymous. Advertisement.

IEEE Spectrum, 59(5):50, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:An

[Ano22n] Anonymous. Advertisement.

IEEE Spectrum, 59(5):51, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ao

[Ano22o] Anonymous. Advertisement.

IEEE Spectrum, 59(6):3, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ap

[Ano22p] Anonymous. Advertisement.

IEEE Spectrum, 59(6):53, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Aq

[Ano22q] Anonymous. Advertisement.

IEEE Spectrum, 59(6):63, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ar

[Ano22r] Anonymous. Advertisement.

IEEE Spectrum, 59(7):3, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:As

[Ano22s] Anonymous. Advertisement. IEEE Spectrum, 59(7):19, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:At

[Ano22t] Anonymous. Advertisement. IEEE Spectrum, 59(7):47, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Au

[Ano22u] Anonymous. Advertisement. IEEE Spectrum, 59(8):3, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Av

[Ano22v] Anonymous. Advertisement. IEEE Spectrum, 59(8):47, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Aw

[Ano22w] Anonymous. Advertisement. IEEE Spectrum, 59(9):3, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ax

[Ano22x] Anonymous. Advertisement. IEEE Spectrum, 59(9):5, September 2022. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ay

[Ano22y] Anonymous. Advertisement.

IEEE Spectrum, 59(9):65, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Az

[Ano22z] Anonymous. Advertisement.

IEEE Spectrum, 59(9):66, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Aba

[Ano22-27] Anonymous. Advertisement.

IEEE Spectrum, 59(9):67, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abb

[Ano22-28] Anonymous. Advertisement.

IEEE Spectrum, 59(10):3, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abd

[Ano22-29] Anonymous. Advertisement.

IEEE Spectrum, 59(10):5, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abe

[Ano22-30] Anonymous. Advertisement.

IEEE Spectrum, 59(10):31, October 2022. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abf

[Ano22-31] Anonymous. Advertisement. IEEE Spectrum, 59(10):39, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abg

[Ano22-32] Anonymous. Advertisement. IEEE Spectrum, 59(10):53, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abh

[Ano22-33] Anonymous. Advertisement. IEEE Spectrum, 59(10):54, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abi

[Ano22-34] Anonymous. Advertisement. IEEE Spectrum, 59(10):55, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abj

[Ano22-35] Anonymous. Advertisement. IEEE Spectrum, 59(11):3, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abk

[Ano22-36] Anonymous. Advertisement. IEEE Spectrum, 59(11):13, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abl

[Ano22-37] Anonymous. Advertisement.

IEEE Spectrum, 59(11):46, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abm

[Ano22-38] Anonymous. Advertisement.

IEEE Spectrum, 59(11):47, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abn

[Ano22-39] Anonymous. Advertisement.

IEEE Spectrum, 59(12):3, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abo

[Ano22-40] Anonymous. Advertisement.

IEEE Spectrum, 59(12):5, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abp

[Ano22-41] Anonymous. Advertisement.

IEEE Spectrum, 59(12):71, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abq

[Ano22-42] Anonymous. Advertisement.

IEEE Spectrum, 59(12):72, December 2022. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abr

[Ano22-43] Anonymous. Advertisement. IEEE Spectrum, 59(12):73, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abs

[Ano22-44] Anonymous. Advertisement. IEEE Spectrum, 59(12):74, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Abt

[Ano22-45] Anonymous. Advertisement. IEEE Spectrum, 59(12):75, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:ABCa

[Ano22-46] Anonymous. Advertisement [back cover]. IEEE Spectrum, 59(2):CV4, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:ABCb

[Ano22-47] Anonymous. Advertisement [back cover]. IEEE Spectrum, 59(3):CV4, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:AFIb

[Ano22-48] Anonymous. Advertisement [front insde cover]. IEEE Spectrum, 59(3):CV2, March 2022.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:AFIa

[Ano22-49] Anonymous. Advertisement [front inside cover]. IEEE Spectrum, 59(2):CV2, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:AMB

[Ano22-50] Anonymous. Advertisement, Matlab [back cover]. IEEE Spectrum, 59(4):CV4, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:AZI

[Ano22-51] Anonymous. Advertisement, Zurich Instruments [front inside cover]. IEEE Spectrum, 59(4):CV2, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BC

[Ano22-52] Anonymous. [Back cover]. IEEE Spectrum, 59(1):CV4, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BICa

[Ano22-53] Anonymous. [Back inside cover].

IEEE Spectrum, 59(1):CV3, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BICb

[Ano22-54] Anonymous. [Back inside cover]. IEEE Spectrum, 59(4):CV3, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BSL

[Ano22-55] Anonymous. Because someone looks up to you [advertisement]. IEEE Spectrum, 59(3):57, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BPE

[Ano22-56] Anonymous. The big picture: Electric buses take charge. IEEE Spectrum, 59(3):14–15, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BPS

[Ano22-57] Anonymous. The big picture: Sun roasted. IEEE Spectrum, 59 (2):14–15, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:BPW

[Ano22-58] Anonymous. The big picture: What’s next? IEEE Spectrum, 59(4):12–13, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:IAA

[Ano22-59] Anonymous. IEEE Access [advertisement]. IEEE Spectrum, 59(3):67, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ca

[Ano22-60] Anonymous. Contents. IEEE Spectrum, 59(1):1–2, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cc

[Ano22-61] Anonymous. Contents. IEEE Spectrum, 59(2):1, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ce

[Ano22-62] Anonymous. Contents. IEEE Spectrum, 59(3):1, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cg

[Ano22-63] Anonymous. Contents. IEEE Spectrum, 59(4):1, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cj

[Ano22-64] Anonymous. Contents. IEEE Spectrum, 59(5):1, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Co

[Ano22-65] Anonymous. Contents. IEEE Spectrum, 59(6):1, June 2022.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ct

[Ano22-66] Anonymous. Contents. IEEE Spectrum, 59(7):1, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cy

[Ano22-67] Anonymous. Contents. IEEE Spectrum, 59(8):1, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbd

[Ano22-68] Anonymous. Contents. IEEE Spectrum, 59(9):1, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbi

[Ano22-69] Anonymous. Contents. IEEE Spectrum, 59(10):1, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbn

[Ano22-70] Anonymous. Contents. IEEE Spectrum, 59(11):1, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbs

[Ano22-71] Anonymous. Contents. IEEE Spectrum, 59(12):1, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cb

[Ano22-72] Anonymous. Contributors. IEEE Spectrum, 59(1):6, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cd

[Ano22-73] Anonymous. Contributors. IEEE Spectrum, 59(2):2, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cf

[Ano22-74] Anonymous. Contributors. IEEE Spectrum, 59(3):3, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ch

[Ano22-75] Anonymous. Contributors. IEEE Spectrum, 59(4):2, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ck

[Ano22-76] Anonymous. Contributors. IEEE Spectrum, 59(5):4, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cp

[Ano22-77] Anonymous. Contributors. IEEE Spectrum, 59(6):4, June 2022.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cu

[Ano22-78] Anonymous. Contributors. IEEE Spectrum, 59(7):4, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cz

[Ano22-79] Anonymous. Contributors. IEEE Spectrum, 59(8):4, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbe

[Ano22-80] Anonymous. Contributors. IEEE Spectrum, 59(9):4, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbj

[Ano22-81] Anonymous. Contributors. IEEE Spectrum, 59(10):4, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbo

[Ano22-82] Anonymous. Contributors. IEEE Spectrum, 59(11):4, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbt

[Ano22-83] Anonymous. Contributors. IEEE Spectrum, 59(12):4, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ci

[Ano22-84] Anonymous. Cover 2. IEEE Spectrum, 59(5):CV2, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cn

[Ano22-85] Anonymous. Cover 2. IEEE Spectrum, 59(6):CV2, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cs

[Ano22-86] Anonymous. Cover 2. IEEE Spectrum, 59(7):CV2, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cx

[Ano22-87] Anonymous. Cover 2. IEEE Spectrum, 59(8):CV2, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbc

[Ano22-88] Anonymous. Cover 2. IEEE Spectrum, 59(9):CV2, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbh

[Ano22-89] Anonymous. Cover 2. IEEE Spectrum, 59(10):CV2, October 2022. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbm

[Ano22-90] Anonymous. Cover 2. IEEE Spectrum, 59(11):CV2, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbr

[Ano22-91] Anonymous. Cover 2. IEEE Spectrum, 59(12):CV2, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cl

[Ano22-92] Anonymous. Cover 3. IEEE Spectrum, 59(5):CV3, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cq

[Ano22-93] Anonymous. Cover 3. IEEE Spectrum, 59(6):CV3, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cv

[Ano22-94] Anonymous. Cover 3. IEEE Spectrum, 59(7):CV3, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cba

[Ano22-95] Anonymous. Cover 3. IEEE Spectrum, 59(8):CV3, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbf

[Ano22-96] Anonymous. Cover 3. IEEE Spectrum, 59(9):CV3, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbk

[Ano22-97] Anonymous. Cover 3. IEEE Spectrum, 59(10):CV3, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbp

[Ano22-98] Anonymous. Cover 3. IEEE Spectrum, 59(11):CV3, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbu

[Ano22-99] Anonymous. Cover 3. IEEE Spectrum, 59(12):CV3, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cm

[Ano22-100] Anonymous. Cover 4. IEEE Spectrum, 59(5):CV4, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cr

[Ano22-101] Anonymous. Cover 4. IEEE Spectrum, 59(6):CV4, June 2022.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cw

[Ano22-102] Anonymous. Cover 4. IEEE Spectrum, 59(7):CV4, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbb

[Ano22-103] Anonymous. Cover 4. IEEE Spectrum, 59(8):CV4, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbg

[Ano22-104] Anonymous. Cover 4. IEEE Spectrum, 59(9):CV4, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbl

[Ano22-105] Anonymous. Cover 4. IEEE Spectrum, 59(10):CV4, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbq

[Ano22-106] Anonymous. Cover 4. IEEE Spectrum, 59(11):CV4, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Cbv

[Ano22-107] Anonymous. Cover 4. IEEE Spectrum, 59(12):CV4, December 2022. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FV

[Ano22-108] Anonymous. A fab and a vision. IEEE Spectrum, 59(6):2, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCa

[Ano22-109] Anonymous. [Front cover]. IEEE Spectrum, 59(1):1, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCb

[Ano22-110] Anonymous. [Front cover].

IEEE Spectrum, 59(2):1, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCc

[Ano22-111] Anonymous. [Front cover]. IEEE Spectrum, 59(3):1, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCd

[Ano22-112] Anonymous. Front cover. IEEE Spectrum, 59(4):1, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCe

[Ano22-113] Anonymous. Front cover. IEEE Spectrum, 59(5):1, May 2022.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCf

[Ano22-114] Anonymous. Front cover. IEEE Spectrum, 59(6):1, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCg

[Ano22-115] Anonymous. Front cover. IEEE Spectrum, 59(7):1, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCh

[Ano22-116] Anonymous. Front cover. IEEE Spectrum, 59(8):1, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCi

[Ano22-117] Anonymous. Front cover. IEEE Spectrum, 59(9):1, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCj

[Ano22-118] Anonymous. Front cover. IEEE Spectrum, 59(10):1, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCk

[Ano22-119] Anonymous. Front cover. IEEE Spectrum, 59(11):1, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FCl

[Ano22-120] Anonymous. Front cover. IEEE Spectrum, 59(12):1, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:FIC

[Ano22-121] Anonymous. [Front inside cover].

IEEE Spectrum, 59(1):CV2, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:IMB

[Ano22-122] Anonymous. IEEE & me [back inside cover]. IEEE Spectrum, 59(2):CV3, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:IMD

[Ano22-123] Anonymous. IEEE member discounts. IEEE Spectrum, 59 (1):15, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Ma

[Ano22-124] Anonymous. Masthead. IEEE Spectrum, 59(8):4, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Mb

[Ano22-125] Anonymous. Masthead. IEEE Spectrum, 59(9):4, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Mc

[Ano22-126] Anonymous. Masthead. IEEE Spectrum, 59(10):4, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Md

[Ano22-127] Anonymous. Masthead. IEEE Spectrum, 59(11):4, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:Me

[Ano22-128] Anonymous. Masthead. IEEE Spectrum, 59(12):4, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:TT

[Ano22-129] Anonymous. Top tech 2022. IEEE Spectrum, 59(1):25, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:TPP

[Ano22-130] Anonymous. The transistor at 75: The past, present, and future of the modern world’s most important invention. IEEE Spectrum, 59(12):22–23, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:WBRa

[Ano22-131] Anonymous. The world’s best robots guide is here! [advertisement]. IEEE Spectrum, 59 (2):3, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2022:WBRb

[Ano22-132] Anonymous. The world’s best robots guide is here! [back inside cover]. IEEE Spectrum, 59(3):CV3, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:QJS

[Ano23a] Anonymous. 5 questions for Janelle Shane: Why AI is so often weird, argumentative, and wrong. IEEE Spectrum, 60 (9):23, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aa

[Ano23b] Anonymous. Advertisement.

IEEE Spectrum, 60(1):3, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ab

[Ano23c] Anonymous. Advertisement.

IEEE Spectrum, 60(1):5, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ac

[Ano23d] Anonymous. Advertisement.

IEEE Spectrum, 60(1):7, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ad

[Ano23e] Anonymous. Advertisement. IEEE Spectrum, 60(1):58, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ae

[Ano23f] Anonymous. Advertisement. IEEE Spectrum, 60(1):59, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Af

[Ano23g] Anonymous. Advertisement. IEEE Spectrum, 60(1):60, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ag

[Ano23h] Anonymous. Advertisement. IEEE Spectrum, 60(1):61, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ah

[Ano23i]  Anonymous.  Advertisement.

IEEE Spectrum, 60(1):62, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ai

[Ano23j]  Anonymous.  Advertisement.

IEEE Spectrum, 60(1):63, January 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aj

[Ano23k] Anonymous. Advertisement.

IEEE Spectrum, 60(3):3, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ak

[Ano23l]  Anonymous.  Advertisement.

IEEE Spectrum, 60(3):13, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Al

[Ano23m] Anonymous. Advertisement.

IEEE Spectrum, 60(3):65, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Am

[Ano23n] Anonymous. Advertisement.

IEEE Spectrum, 60(3):66, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:An

[Ano23o] Anonymous. Advertisement.

IEEE Spectrum, 60(3):67, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ao

[Ano23p] Anonymous. Advertisement.

IEEE Spectrum, 60(4):3, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ap

[Ano23q] Anonymous. Advertisement. IEEE Spectrum, 60(4):31, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aq

[Ano23r] Anonymous. Advertisement. IEEE Spectrum, 60(4):54, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ar

[Ano23s] Anonymous. Advertisement. IEEE Spectrum, 60(4):55, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:As

[Ano23t] Anonymous. Advertisement. IEEE Spectrum, 60(5):3, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:At

[Ano23u] Anonymous. Advertisement. IEEE Spectrum, 60(5):5, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Au

[Ano23v] Anonymous. Advertisement. IEEE Spectrum, 60(5):6, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Av

[Ano23w] Anonymous. Advertisement.

IEEE Spectrum, 60(5):45, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aw

[Ano23x] Anonymous. Advertisement.

IEEE Spectrum, 60(5):53, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ax

[Ano23y] Anonymous. Advertisement.

IEEE Spectrum, 60(5):54, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ay

[Ano23z] Anonymous. Advertisement.

IEEE Spectrum, 60(5):55, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Az

[Ano23-27] Anonymous. Advertisement.

IEEE Spectrum, 60(6):3, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aba

[Ano23-28] Anonymous. Advertisement.

IEEE Spectrum, 60(6):13, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abb

[Ano23-29] Anonymous. Advertisement. IEEE Spectrum, 60(6):67, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abc

[Ano23-30] Anonymous. Advertisement. IEEE Spectrum, 60(7):3, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abd

[Ano23-31] Anonymous. Advertisement. IEEE Spectrum, 60(7):5, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abe

[Ano23-32] Anonymous. Advertisement. IEEE Spectrum, 60(7):50, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abf

[Ano23-33] Anonymous. Advertisement. IEEE Spectrum, 60(7):51, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abg

[Ano23-34] Anonymous. Advertisement. IEEE Spectrum, 60(8):CV2, August 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abh

[Ano23-35] Anonymous. Advertisement.

IEEE Spectrum, 60(8):3, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abi

[Ano23-36] Anonymous. Advertisement.

IEEE Spectrum, 60(8):23, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abj

[Ano23-37] Anonymous. Advertisement.

IEEE Spectrum, 60(8):37, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abk

[Ano23-38] Anonymous. Advertisement.

IEEE Spectrum, 60(8):50, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abl

[Ano23-39] Anonymous. Advertisement.

IEEE Spectrum, 60(8):51, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abm

[Ano23-40] Anonymous. Advertisement.

IEEE Spectrum, 60(8):CV3, August 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abn

[Ano23-41] Anonymous. Advertisement. IEEE Spectrum, 60(8):CV4, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abo

[Ano23-42] Anonymous. Advertisement. IEEE Spectrum, 60(9):3, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abp

[Ano23-43] Anonymous. Advertisement. IEEE Spectrum, 60(9):5, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abq

[Ano23-44] Anonymous. Advertisement. IEEE Spectrum, 60(9):6, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abr

[Ano23-45] Anonymous. Advertisement. IEEE Spectrum, 60(9):15, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abs

[Ano23-46] Anonymous. Advertisement. IEEE Spectrum, 60(9):67, September 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abt

[Ano23-47] Anonymous. Advertisement.

IEEE Spectrum, 60(10):3, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abu

[Ano23-48] Anonymous. Advertisement.

IEEE Spectrum, 60(10):50, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abv

[Ano23-49] Anonymous. Advertisement.

IEEE Spectrum, 60(10):51, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abw

[Ano23-50] Anonymous. Advertisement.

IEEE Spectrum, 60(11):3, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abx

[Ano23-51] Anonymous. Advertisement.

IEEE Spectrum, 60(11):5, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aby

[Ano23-52] Anonymous. Advertisement.

IEEE Spectrum, 60(11):51, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Abz

[Ano23-53] Anonymous. Advertisement. IEEE Spectrum, 60(12):3, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Aca

[Ano23-54] Anonymous. Advertisement. IEEE Spectrum, 60(12):13, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Acb

[Ano23-55] Anonymous. Advertisement. IEEE Spectrum, 60(12):46, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Acc

[Ano23-56] Anonymous. Advertisement. IEEE Spectrum, 60(12):65, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Acd

[Ano23-57] Anonymous. Advertisement. IEEE Spectrum, 60(12):66, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ace

[Ano23-58] Anonymous. Advertisement. IEEE Spectrum, 60(12):67, December 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:CAa

[Ano23-59] Anonymous. COMSOL — advertisement. IEEE Spectrum, 60 (1):1s–36s, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cb

[Ano23-60] Anonymous. Contents. IEEE Spectrum, 60(1):1–2, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cg

[Ano23-61] Anonymous. Contents. IEEE Spectrum, 60(3):1–2, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cl

[Ano23-62] Anonymous. Contents. IEEE Spectrum, 60(4):1, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cq

[Ano23-63] Anonymous. Contents. IEEE Spectrum, 60(5):1, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cv

[Ano23-64] Anonymous. Contents. IEEE Spectrum, 60(6):1, June 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cba

[Ano23-65] Anonymous. Contents. IEEE Spectrum, 60(7):1, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbe

[Ano23-66] Anonymous. Contents. IEEE Spectrum, 60(8):1, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbh

[Ano23-67] Anonymous. Contents. IEEE Spectrum, 60(9):1, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbm

[Ano23-68] Anonymous. Contents. IEEE Spectrum, 60(10):1, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbr

[Ano23-69] Anonymous. Contents. IEEE Spectrum, 60(11):1, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbw

[Ano23-70] Anonymous. Contents. IEEE Spectrum, 60(12):1, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cc

[Ano23-71] Anonymous. Contributors. IEEE Spectrum, 60(1):6, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ch

[Ano23-72] Anonymous. Contributors. IEEE Spectrum, 60(3):4, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cm

[Ano23-73] Anonymous. Contributors. IEEE Spectrum, 60(4):4, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cr

[Ano23-74] Anonymous. Contributors. IEEE Spectrum, 60(5):4, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cw

[Ano23-75] Anonymous. Contributors. IEEE Spectrum, 60(6):4, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbb

[Ano23-76] Anonymous. Contributors. IEEE Spectrum, 60(7):4, July 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbf

[Ano23-77] Anonymous. Contributors. IEEE Spectrum, 60(8):4, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbi

[Ano23-78] Anonymous. Contributors. IEEE Spectrum, 60(9):4, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbn

[Ano23-79] Anonymous. Contributors. IEEE Spectrum, 60(10):4, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbs

[Ano23-80] Anonymous. Contributors. IEEE Spectrum, 60(11):4, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbx

[Ano23-81] Anonymous. Contributors. IEEE Spectrum, 60(12):4, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ca

[Ano23-82] Anonymous. Cover 2. IEEE Spectrum, 60(1):CV2, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cf

[Ano23-83] Anonymous. Cover 2. IEEE Spectrum, 60(3):CV2, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ck

[Ano23-84] Anonymous. Cover 2. IEEE Spectrum, 60(4):CV2, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cp

[Ano23-85] Anonymous. Cover 2. IEEE Spectrum, 60(5):CV2, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cu

[Ano23-86] Anonymous. Cover 2. IEEE Spectrum, 60(6):CV2, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cz

[Ano23-87] Anonymous. Cover 2. IEEE Spectrum, 60(7):1, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbl

[Ano23-88] Anonymous. Cover 2. IEEE Spectrum, 60(10):CV2, October 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbq

[Ano23-89] Anonymous. Cover 2. IEEE Spectrum, 60(11):CV2, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbv

[Ano23-90] Anonymous. Cover 2. IEEE Spectrum, 60(12):CV2, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cd

[Ano23-91] Anonymous. Cover 3. IEEE Spectrum, 60(1):CV3, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ci

[Ano23-92] Anonymous. Cover 3. IEEE Spectrum, 60(3):CV3, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cn

[Ano23-93] Anonymous. Cover 3. IEEE Spectrum, 60(4):CV3, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cs

[Ano23-94] Anonymous. Cover 3. IEEE Spectrum, 60(5):CV3, May 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cx

[Ano23-95] Anonymous. Cover 3. IEEE Spectrum, 60(6):CV3, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbc

[Ano23-96] Anonymous. Cover 3. IEEE Spectrum, 60(7):1, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbo

[Ano23-97] Anonymous. Cover 3. IEEE Spectrum, 60(10):CV3, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbt

[Ano23-98] Anonymous. Cover 3. IEEE Spectrum, 60(11):CV3, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cby

[Ano23-99] Anonymous. Cover 3. IEEE Spectrum, 60(12):CV3, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ce

[Ano23-100] Anonymous. Cover 4. IEEE Spectrum, 60(1):CV4, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cj

[Ano23-101] Anonymous. Cover 4. IEEE Spectrum, 60(3):CV4, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Co

[Ano23-102] Anonymous. Cover 4. IEEE Spectrum, 60(4):CV4, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ct

[Ano23-103] Anonymous. Cover 4. IEEE Spectrum, 60(5):CV4, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cy

[Ano23-104] Anonymous. Cover 4. IEEE Spectrum, 60(6):CV4, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbd

[Ano23-105] Anonymous. Cover 4. IEEE Spectrum, 60(7):1, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbp

[Ano23-106] Anonymous. Cover 4. IEEE Spectrum, 60(10):CV4, October 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbu

[Ano23-107] Anonymous. Cover 4. IEEE Spectrum, 60(11):CV4, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbz

[Ano23-108] Anonymous. Cover 4. IEEE Spectrum, 60(12):CV4, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbg

[Ano23-109] Anonymous. Cover2. IEEE Spectrum, 60(9):1, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbj

[Ano23-110] Anonymous. Cover3. IEEE Spectrum, 60(9):1, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Cbk

[Ano23-111] Anonymous. Cover4. IEEE Spectrum, 60(9):1, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCa

[Ano23-112] Anonymous. Front cover. IEEE Spectrum, 60(1):1, January 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCb

[Ano23-113] Anonymous. Front cover. IEEE Spectrum, 60(3):1, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCc

[Ano23-114] Anonymous. Front cover. IEEE Spectrum, 60(4):1, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCd

[Ano23-115] Anonymous. Front cover. IEEE Spectrum, 60(5):1, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCe

[Ano23-116] Anonymous. Front cover. IEEE Spectrum, 60(6):1, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCf

[Ano23-117] Anonymous. Front cover. IEEE Spectrum, 60(7):1, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCg

[Ano23-118] Anonymous. Front cover. IEEE Spectrum, 60(8):1, August 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCh

[Ano23-119] Anonymous. Front cover. IEEE Spectrum, 60(9):1, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCi

[Ano23-120] Anonymous. Front cover. IEEE Spectrum, 60(10):1, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCj

[Ano23-121] Anonymous. Front cover. IEEE Spectrum, 60(11):1, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:FCk

[Ano23-122] Anonymous. Front cover. IEEE Spectrum, 60(12):1, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Ma

[Ano23-123] Anonymous. Masthead. IEEE Spectrum, 60(1):6, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mb

[Ano23-124] Anonymous. Masthead. IEEE Spectrum, 60(3):4, March 2023.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mc

[Ano23-125] Anonymous. Masthead. IEEE Spectrum, 60(4):4, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Md

[Ano23-126] Anonymous. Masthead. IEEE Spectrum, 60(5):4, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Me

[Ano23-127] Anonymous. Masthead. IEEE Spectrum, 60(6):4, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mf

[Ano23-128] Anonymous. Masthead. IEEE Spectrum, 60(7):4, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mg

[Ano23-129] Anonymous. Masthead. IEEE Spectrum, 60(8):4, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mh

[Ano23-130] Anonymous. Masthead. IEEE Spectrum, 60(9):4, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mi

[Ano23-131] Anonymous. Masthead. IEEE Spectrum, 60(10):4, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:Mj

[Ano23-132] Anonymous. Masthead. IEEE Spectrum, 60(11):4, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:OA

[Ano23-133] Anonymous. Ode to an arm.

IEEE Spectrum, 60(8):52, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:PF

[Ano23-134] Anonymous. Past forward. IEEE Spectrum, 60(9):68, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:S

[Ano23-135] Anonymous. Staff. IEEE Spectrum, 60(12):4, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:TT

[Ano23-136] Anonymous. Top tech 2023.

IEEE Spectrum, 60(1):23, January 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2023:WPH

[Ano23-137] Anonymous. Wind-powered home heating. IEEE Spectrum, 60(8):14–15, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. Advertisement. IEEE Spectrum, 61(01):3, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. Advertisement. IEEE Spectrum, 61(01):5, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ac

[Ano24c] Anonymous. Advertisement. IEEE Spectrum, 61(01):21, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ad

[Ano24d] Anonymous. Advertisement. IEEE Spectrum, 61(01):59, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ae

[Ano24e] Anonymous. Advertisement. IEEE Spectrum, 61(01):60, January 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Af

[Ano24f] Anonymous. Advertisement.

IEEE Spectrum, 61(01):61, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ag

[Ano24g] Anonymous. Advertisement.

IEEE Spectrum, 61(01):62, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ah

[Ano24h] Anonymous. Advertisement.

IEEE Spectrum, 61(01):63, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ai

[Ano24i]  Anonymous.  Advertisement.

IEEE Spectrum, 61(01):1, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aj

[Ano24j]  Anonymous.  Advertisement.

IEEE Spectrum, 61(01):1, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ak

[Ano24k] Anonymous. Advertisement.

IEEE Spectrum, 61(2):3, February 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Al

[Ano24l]  Anonymous.  Advertisement.

IEEE Spectrum, 61(2):35, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Am

[Ano24m] Anonymous. Advertisement. IEEE Spectrum, 61(2):47, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:An

[Ano24n] Anonymous. Advertisement. IEEE Spectrum, 61(3):3, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ao

[Ano24o] Anonymous. Advertisement. IEEE Spectrum, 61(3):13, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ap

[Ano24p] Anonymous. Advertisement. IEEE Spectrum, 61(3):63, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aq

[Ano24q] Anonymous. Advertisement. IEEE Spectrum, 61(4):3, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ar

[Ano24r] Anonymous. Advertisement.

IEEE Spectrum, 61(4):5, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:As

[Ano24s] Anonymous. Advertisement.

IEEE Spectrum, 61(4):51, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:At

[Ano24t] Anonymous. Advertisement.

IEEE Spectrum, 61(5):3, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Au

[Ano24u] Anonymous. Advertisement.

IEEE Spectrum, 61(5):35, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Av

[Ano24v] Anonymous. Advertisement.

IEEE Spectrum, 61(5):58, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aw

[Ano24w] Anonymous. Advertisement.

IEEE Spectrum, 61(5):59, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ax

[Ano24x] Anonymous. Advertisement. IEEE Spectrum, 61(6):3, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ay

[Ano24y] Anonymous. Advertisement. IEEE Spectrum, 61(6):5, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Az

[Ano24z] Anonymous. Advertisement. IEEE Spectrum, 61(6):67, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aba

[Ano24-27] Anonymous. Advertisement. IEEE Spectrum, 61(7):3, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abb

[Ano24-28] Anonymous. Advertisement. IEEE Spectrum, 61(7):13, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abc

[Ano24-29] Anonymous. Advertisement. IEEE Spectrum, 61(8):3, August 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abd

[Ano24-30] Anonymous. Advertisement.

IEEE Spectrum, 61(8):5, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abe

[Ano24-31] Anonymous. Advertisement.

IEEE Spectrum, 61(8):15, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abf

[Ano24-32] Anonymous. Advertisement.

IEEE Spectrum, 61(9):3, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abg

[Ano24-33] Anonymous. Advertisement.

IEEE Spectrum, 61(9):13, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abh

[Ano24-34] Anonymous. Advertisement.

IEEE Spectrum, 61(9):21, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abi

[Ano24-35] Anonymous. Advertisement.

IEEE Spectrum, 61(9):48, September 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abj

[Ano24-36] Anonymous. Advertisement. IEEE Spectrum, 61(9):65, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abk

[Ano24-37] Anonymous. Advertisement. IEEE Spectrum, 61(9):66, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abl

[Ano24-38] Anonymous. Advertisement. IEEE Spectrum, 61(9):67, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abm

[Ano24-39] Anonymous. Advertisement. IEEE Spectrum, 61(10):3, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abn

[Ano24-40] Anonymous. Advertisement. IEEE Spectrum, 61(10):13, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abo

[Ano24-41] Anonymous. Advertisement. IEEE Spectrum, 61(10):46, October 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abp

[Ano24-42] Anonymous. Advertisement.

IEEE Spectrum, 61(10):47, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abq

[Ano24-43] Anonymous. Advertisement.

IEEE Spectrum, 61(11):3, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abr

[Ano24-44] Anonymous. Advertisement.

IEEE Spectrum, 61(11):13, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abs

[Ano24-45] Anonymous. Advertisement.

IEEE Spectrum, 61(11):17, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abt

[Ano24-46] Anonymous. Advertisement.

IEEE Spectrum, 61(11):23, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abu

[Ano24-47] Anonymous. Advertisement.

IEEE Spectrum, 61(11):71, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abv

[Ano24-48] Anonymous. Advertisement. IEEE Spectrum, 61(11):72, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abw

[Ano24-49] Anonymous. Advertisement. IEEE Spectrum, 61(11):73, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abx

[Ano24-50] Anonymous. Advertisement. IEEE Spectrum, 61(11):74, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aby

[Ano24-51] Anonymous. Advertisement. IEEE Spectrum, 61(11):75, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Abz

[Ano24-52] Anonymous. Advertisement. IEEE Spectrum, 61(12):3, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Aca

[Ano24-53] Anonymous. Advertisement. IEEE Spectrum, 61(12):6, December 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Acb

[Ano24-54] Anonymous. Advertisement.

IEEE Spectrum, 61(12):15, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Acc

[Ano24-55] Anonymous. Advertisement.

IEEE Spectrum, 61(12):31, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Acd

[Ano24-56] Anonymous. Advertisement.

IEEE Spectrum, 61(12):67, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ace

[Ano24-57] Anonymous. Advertisement.

IEEE Spectrum, 61(12):68, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Acf

[Ano24-58] Anonymous. Advertisement.

IEEE Spectrum, 61(12):69, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Acg

[Ano24-59] Anonymous. Advertisement.

IEEE Spectrum, 61(12):70, December 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ach

[Ano24-60] Anonymous. Advertisement. IEEE Spectrum, 61(12):71, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cb

[Ano24-61] Anonymous. Contents. IEEE Spectrum, 61(01):1–2, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ce

[Ano24-62] Anonymous. Contents. IEEE Spectrum, 61(2):1, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cj

[Ano24-63] Anonymous. Contents. IEEE Spectrum, 61(3):1, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Co

[Ano24-64] Anonymous. Contents. IEEE Spectrum, 61(4):1, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ct

[Ano24-65] Anonymous. Contents. IEEE Spectrum, 61(5):1, May 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cy

[Ano24-66] Anonymous. Contents. IEEE Spectrum, 61(6):1, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbd

[Ano24-67] Anonymous. Contents. IEEE Spectrum, 61(7):1, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbi

[Ano24-68] Anonymous. Contents. IEEE Spectrum, 61(8):1, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbn

[Ano24-69] Anonymous. Contents. IEEE Spectrum, 61(9):1, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbs

[Ano24-70] Anonymous. Contents. IEEE Spectrum, 61(10):1, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbx

[Ano24-71] Anonymous. Contents. IEEE Spectrum, 61(11):1, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ccc

[Ano24-72] Anonymous. Contents. IEEE Spectrum, 61(12):1, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cc

[Ano24-73] Anonymous. Contributors. IEEE Spectrum, 61(01):6, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cf

[Ano24-74] Anonymous. Contributors. IEEE Spectrum, 61(2):4, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ck

[Ano24-75] Anonymous. Contributors. IEEE Spectrum, 61(3):4, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cp

[Ano24-76] Anonymous. Contributors. IEEE Spectrum, 61(4):4, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cu

[Ano24-77] Anonymous. Contributors. IEEE Spectrum, 61(5):4, May 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cz

[Ano24-78] Anonymous. Contributors. IEEE Spectrum, 61(6):4, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbe

[Ano24-79] Anonymous. Contributors. IEEE Spectrum, 61(7):4, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbj

[Ano24-80] Anonymous. Contributors. IEEE Spectrum, 61(8):4, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbo

[Ano24-81] Anonymous. Contributors. IEEE Spectrum, 61(9):4, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbt

[Ano24-82] Anonymous. Contributors. IEEE Spectrum, 61(10):4, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cby

[Ano24-83] Anonymous. Contributors. IEEE Spectrum, 61(11):4, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ccd

[Ano24-84] Anonymous. Contributors. IEEE Spectrum, 61(12):5, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cd

[Ano24-85] Anonymous. Cover 2. IEEE Spectrum, 61(2):CV2, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ci

[Ano24-86] Anonymous. Cover 2. IEEE Spectrum, 61(3):CV2, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cn

[Ano24-87] Anonymous. Cover 2. IEEE Spectrum, 61(4):CV2, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cs

[Ano24-88] Anonymous. Cover 2. IEEE Spectrum, 61(5):CV2, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cx

[Ano24-89] Anonymous. Cover 2. IEEE Spectrum, 61(6):CV2, June 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbc

[Ano24-90] Anonymous. Cover 2. IEEE Spectrum, 61(7):CV2, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbh

[Ano24-91] Anonymous. Cover 2. IEEE Spectrum, 61(8):CV2, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbm

[Ano24-92] Anonymous. Cover 2. IEEE Spectrum, 61(9):CV2, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbr

[Ano24-93] Anonymous. Cover 2. IEEE Spectrum, 61(10):CV2, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbw

[Ano24-94] Anonymous. Cover 2. IEEE Spectrum, 61(11):CV2, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ccb

[Ano24-95] Anonymous. Cover 2. IEEE Spectrum, 61(12):CV2, December 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cg

[Ano24-96] Anonymous. Cover 3. IEEE Spectrum, 61(2):CV3, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cl

[Ano24-97] Anonymous. Cover 3. IEEE Spectrum, 61(3):CV3, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cq

[Ano24-98] Anonymous. Cover 3. IEEE Spectrum, 61(4):CV3, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cv

[Ano24-99] Anonymous. Cover 3. IEEE Spectrum, 61(5):CV3, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cba

[Ano24-100] Anonymous. Cover 3. IEEE Spectrum, 61(6):CV3, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbf

[Ano24-101] Anonymous. Cover 3. IEEE Spectrum, 61(7):CV3, July 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbk

[Ano24-102] Anonymous. Cover 3. IEEE Spectrum, 61(8):CV3, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbp

[Ano24-103] Anonymous. Cover 3. IEEE Spectrum, 61(9):CV3, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbu

[Ano24-104] Anonymous. Cover 3. IEEE Spectrum, 61(10):CV3, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbz

[Ano24-105] Anonymous. Cover 3. IEEE Spectrum, 61(11):CV3, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cce

[Ano24-106] Anonymous. Cover 3. IEEE Spectrum, 61(12):CV3, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ch

[Ano24-107] Anonymous. Cover 4. IEEE Spectrum, 61(2):CV4, February 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cm

[Ano24-108] Anonymous. Cover 4. IEEE Spectrum, 61(3):CV4, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cr

[Ano24-109] Anonymous. Cover 4. IEEE Spectrum, 61(4):CV4, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cw

[Ano24-110] Anonymous. Cover 4. IEEE Spectrum, 61(5):CV4, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbb

[Ano24-111] Anonymous. Cover 4. IEEE Spectrum, 61(6):CV4, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbg

[Ano24-112] Anonymous. Cover 4. IEEE Spectrum, 61(7):CV4, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbl

[Ano24-113] Anonymous. Cover 4. IEEE Spectrum, 61(8):CV4, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbq

[Ano24-114] Anonymous. Cover 4. IEEE Spectrum, 61(9):CV4, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cbv

[Ano24-115] Anonymous. Cover 4. IEEE Spectrum, 61(10):CV4, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Cca

[Ano24-116] Anonymous. Cover 4. IEEE Spectrum, 61(11):CV4, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ccf

[Ano24-117] Anonymous. Cover 4. IEEE Spectrum, 61(12):CV4, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ca

[Ano24-118] Anonymous. Cover2. IEEE Spectrum, 61(01):1, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCa

[Ano24-119] Anonymous. Front cover. IEEE Spectrum, 61(01):1, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCb

[Ano24-120] Anonymous. Front cover. IEEE Spectrum, 61(2):1, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCc

[Ano24-121] Anonymous. Front cover. IEEE Spectrum, 61(3):1, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCd

[Ano24-122] Anonymous. Front cover. IEEE Spectrum, 61(4):1, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCe

[Ano24-123] Anonymous. Front cover. IEEE Spectrum, 61(5):1, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCf

[Ano24-124] Anonymous. Front cover. IEEE Spectrum, 61(6):1, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCg

[Ano24-125] Anonymous. Front cover. IEEE Spectrum, 61(7):1, July 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCh

[Ano24-126] Anonymous. Front cover. IEEE Spectrum, 61(8):1, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCi

[Ano24-127] Anonymous. Front cover. IEEE Spectrum, 61(9):1, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCj

[Ano24-128] Anonymous. Front cover. IEEE Spectrum, 61(10):1, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCk

[Ano24-129] Anonymous. Front cover. IEEE Spectrum, 61(11):1, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:FCl

[Ano24-130] Anonymous. Front cover. IEEE Spectrum, 61(12):1, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ma

[Ano24-131] Anonymous. Masthead. IEEE Spectrum, 61(01):6, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mb

[Ano24-132] Anonymous. Masthead. IEEE Spectrum, 61(2):4, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mc

[Ano24-133] Anonymous. Masthead. IEEE Spectrum, 61(3):4, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Md

[Ano24-134] Anonymous. Masthead. IEEE Spectrum, 61(4):4, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Me

[Ano24-135] Anonymous. Masthead. IEEE Spectrum, 61(5):4, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mf

[Ano24-136] Anonymous. Masthead. IEEE Spectrum, 61(6):4, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mg

[Ano24-137] Anonymous. Masthead. IEEE Spectrum, 61(7):4, July 2024.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mh

[Ano24-138] Anonymous. Masthead. IEEE Spectrum, 61(8):4, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mi

[Ano24-139] Anonymous. Masthead. IEEE Spectrum, 61(9):4, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mj

[Ano24-140] Anonymous. Masthead. IEEE Spectrum, 61(10):4, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Mk

[Ano24-141] Anonymous. Masthead. IEEE Spectrum, 61(11):4, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:Ml

[Ano24-142] Anonymous. Masthead. IEEE Spectrum, 61(12):5, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:MM

[Ano24-143] Anonymous. Mechanical megatalent. IEEE Spectrum, 61(4):

14–15, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:MEV

[Ano24-144] Anonymous. A mouse’s-eye view. IEEE Spectrum, 61 (3):14–15, March 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:TVE

[Ano24-145] Anonymous. A taste of Venus on Earth. IEEE Spectrum, 61 (2):14–15, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2024:TT

[Ano24-146] Anonymous. Top tech 2024. IEEE Spectrum, 61(01):23, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Armstrong:2020:VDS

[Arm20] C. M. Armstrong. These vacuum devices stood guard during the Cold War, advanced particle physics, treated cancer patients, and made the Beatles sound better. IEEE Spectrum, 57(11): 30–36, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aronstein:2024:FSF

[Aro24] Jesse Aronstein. The forgotten story of the first automated fab: In 1970, an IBM middle manager had a vision: a finished chip in one day. IEEE Spectrum, 61 (12):44–50, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:Na

[ARPH22] Evan Ackerman, Rahul Rao, Julianne Pepitone, and Michelle Hampson. News. IEEE Spectrum, 59(5):7–13, May 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:YRW

[AS22] Evan Ackerman and Eliza Strickland. Are you ready for workplace brain scanning?: Leveraging brain data will make workers happier and more productive, backers say. IEEE Spectrum, 59 (12):46–52, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ashruf:2023:CTP

[Ash23] Colin Ashruf. Careers: 10 tips for product development: Getting to market requires more than clever engineering. IEEE Spectrum, 60(4):19–20, April 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2022:TWB

[AZPC22] Evan Ackerman, Glenn Zorpette, Julianne Pepitone, and Charles Q. Choi. Transportation: What’s behind the air-taxi craze: a wave of eVTOL startups aim to revolutionize transportation. IEEE Spectrum, 59(3):6–13, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bajema:2022:WAS

[Baj22] Natasha Bajema. Will AI steal submarines’ stealth?: Better detection will make the oceans transparent and perhaps undermine nuclear deterrence. IEEE Spectrum, 59(9):36–41, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bajema:2024:NMH

[Baj24] Natasha Bajema. No more hideand-seek: The scramble to preserve submarine stealth in an age of AI and all-seeing sensors. IEEE Spectrum, 61(12): 24–30, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barger:2020:TRT

[Bar20] T. S. Barger. Turning robots into teacher’s aides: Christian Wanamaker’s robotics software helps students on the autism spectrum — careers. IEEE Spectrum, 57(3):19, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barde:2024:CSS

[Bar24] Henri Barde. Castles in the sky?: a skeptic’s take on beaming power to earth from space. IEEE Spectrum, 61(6):22–29, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baur:2024:URL

[Bau24] Jasper Baur. Ukraine is riddled with land mines: Drones and AI can help. IEEE Spectrum, 61(5):42–49, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baramsai:2022:NNS

[BBFS22] Bayarbadrakh Baramsai, Theresa Benyo, Lawrence Forsley, and Bruce Steinetz. NASA’s new shortcut to fusion power: Lattice confinement fusion eliminates massive magnets and powerful lasers. IEEE Spectrum, 59(3):32–37, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bicknell:2023:HDA

[BBS23] Klinton Bicknell, Claire Brust, and Burr Settles. How Duolingo’s AI learns what you need to learn: The languagelearning app tries to emulate a great human tutor. IEEE Spectrum, 60(3):28–33, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berlinguette:2019:RCC

[BCM^{+}19] Curtis P. Berlinguette, Yet-Ming Chiang, Jeremy N. Munday, Thomas Schenkel, David K. Fork, Ross Koningstein, and Matthew D. Trevithick. Revisiting the cold case of cold fusion. Nature, 570(7759):45–51, May 2019. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Balossino:2022:MND

[BD22] Alessandro Balossino and Faramaz Davarian. The Moon needs decent wireless coverage: Argotec and JPL’s relay satellites could deliver bandwidth for more than 90 missions. IEEE Spectrum, 59(4):32–37, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beane:2022:TRS

[Bea22] Matt Beane. Today’s robotic surgery turns surgical trainees into spectators: Medical training in the robotics age leaves tomorrow’s surgeons short on skills. IEEE Spectrum, 59 (8):32–37, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barth:2024:IMT

[BF24] Phillip W. Barth and Leslie A. Field. Inkjets are for more than just printing: They can build DNA arrays, 3D structures, and much more. IEEE Spectrum, 61(4):32–37, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bajema:2021:N

[BGR^{+}21] Natasha Bajema, Maria Gallucci, Rahul Rao, Philip E. Ross, and Michelle Hampson. News. IEEE Spectrum, 58 (6):6–13, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bajema:2024:N

[BJZM24] Natasha Bajema, Andrew Jones, Glenn Zorpette, and Samuel K. Moore. News. IEEE Spectrum, 61(5):5–13, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bilen:2020:WNR

[BMRW20] S. Bilen, D. Mortensen, R. Reinhart, and A. Wyglinski. Where no radio has gone before: Cognitive radios can keep deep-space missions connected to earth even when faced with alien environments. IEEE Spectrum, 57 (8):44–50, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boris:2024:TCY

[Bor24] Alan Boris. Take control of your home: Escape proprietary smart-home tech with this DIY panel. IEEE Spectrum, 61(9): 16–18, September 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bouman:2020:PBH

[Bou20] K. L. Bouman. Portrait of a black hole: Here’s how the event horizon telescope team pieced together a now-famous image. IEEE Spectrum, 57(2): 22–29, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bouton:2021:BIW

[Bou21] C. Bouton. Brain implants and wearables reroute signals to restore movement and sensation. IEEE Spectrum, 58(2): 28–33, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bowman:2021:CBB

[Bow21] Keith A. Bowman. A circuit to boost battery life: All-digital versions of the low-dropout voltage regulator will save time, money, and power. IEEE Spectrum, 58(8):40–45, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2024:FML

[Boy24] John Boyd. Is the future of Moore’s Law in a particle accelerator?: Wiggling electrons could turbocharge EUV lithography. IEEE Spectrum, 61(7):28– 33, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brand:2021:RCH

[Bra21] J. Brand. The radio that can hear over itself: Self-interference cancellation allows radios to transmit and receive on the same frequency. IEEE Spectrum, 58 (3):30–35, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brewster:2020:PFS

[Bre20]  R. Brewster.  Past forward

[the SCR-536 Handie-Talkie]. IEEE Spectrum, 57(10):72, October 2020. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2021:HLA

[Bro21] Rodney Brooks. A human in the loop: AI won’t surpass human intelligence anytime soon. IEEE Spectrum, 58(10):48–49, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2022:LRO

[Bro22a] Rodney Brooks. The long road to overnight success: Reusable rocket boosters were decades in the making. IEEE Spectrum, 59(4):20–21, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2022:MMN

[Bro22b] Rodney Brooks. Myth and machine: 22 November 1935: The day globalism was born: Juan Trippe’s Pan Am Clippers launched a new era. IEEE Spectrum, 59(6):22–23, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2022:MMB

[Bro22c] Rodney Brooks. Myth and machine: Building boom: Population growth and climate change have an upside. IEEE Spectrum, 59(3):23, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2022:MMO

[Bro22d] Rodney Brooks. Myth and machine: The other side of the Innovator’s Dilemma: Beware the quiet disruptors. IEEE Spectrum, 59(7):20–21, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brock:2023:MTC

[Bro23a] David C. Brock. The machine that transformed computing: The Xerox Alto combined windows, WYSIWYG, menus, and networking to make the modern PC. IEEE Spectrum, 60(3):40–45, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brock:2023:PF

[Bro23b] David C. Brock. Past forward. IEEE Spectrum, 60(3):68, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Broo:2023:HGA

[Bro23c] Didem Gürdür Broo. How generative AI helped me imagine a better robot: It didn’t give me schematics, but it did boost my creativity. IEEE Spectrum, 60 (11):44–50, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Biercuk:2022:QEC

[BS22] Michael J. Biercuk and Thomas M. Stace. Quantum error correction at the threshold: If technologists don’t get beyond it, quantum computers will never be big. IEEE Spectrum, 59(7):28– 46, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Burr:2021:OLK

[BSAH21] Geoffrey W. Burr, Abu Sebastian, Takashi Ando, and Wilfried Haensch. Ohm’s Law + Kirchhoff’s current law = better AI: Neural-network processing done in memory with analog circuits will save energy. IEEE Spectrum, 58(12):44–49, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bajema:2022:ARW

[BSG22] Natasha Bajema, Craig S. Smith, and Dan Garisto. AI’s real worst-case scenarios: Who needs Terminators when you have precision clickbait and ultradeepfakes? IEEE Spectrum, 59 (1):8–14, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baets:2023:WPL

[BVKR23] Roel Baets, Joris Van Campenhout, Bernardette Kunert, and Gunther Roelkens. 4 ways to put lasers on silicon: You can make many things with silicon photonics, but a laser is not one of them. IEEE Spectrum, 60(5):32–37, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caballar:2024:WND

[Cab24] Rina Diane Caballar. We need to decarbonize software: The way we write software has unappreciated environmental impacts. IEEE Spectrum, 61 (4):26–31, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Calaiaro:2022:ATD

[Cal22] Jason Calaiaro. AI takes a dumpster dive: Computer-vision systems sort your recyclables at superhuman speed. IEEE Spectrum, 59(7):22–27, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Calacci:2024:GWW

[Cal24] Dana Calacci. The GIG workers who fought an algorithm: When their pay suddenly dropped, Shipt’s delivery drivers dug into the data. IEEE Spectrum, 61 (8):42–47, August 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:2024:WRD

[Car24] Rob Carlson. The world is running out of data storage. here’s how DNA can save US. IEEE Spectrum, 61(3):28–44, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:GLW

[Cas20a] S. Cass. Geek life: Write a program in a single tweet: What can you code in 280 characters? IEEE Spectrum, 57(5):54, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:LTA

[Cas20b] S. Cass. Let a thousand analog oscillators sing: Sam Battle, champion of analog engineering — [Geek Life]. IEEE Spectrum, 57(10):15, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:LM

[Cas20c] S. Cass. Life on Mars? IEEE Spectrum, 57(1):42–43, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:NMI

[Cas20d] S. Cass. Nvidia makes it easy to embed AI: The Jetson nano packs a lot of machinelearning power into DIY projects — [hands on]. IEEE Spectrum, 57(7):14–16, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:OCB

[Cas20e] S. Cass. Oscilloscope clock:

Build a vector-graphics display using an old cathode-ray tube — [hands on]. IEEE Spectrum, 57 (1):14–16, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:PFP

[Cas20f] S. Cass. Painless FPGA programming: The alchitry AU kit can simplify projects that need a lot of input/output — [hands on]. IEEE Spectrum, 57(12): 18–20, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:PEA

[Cas20g] S. Cass. Plotting experiments: The AxiDraw minikit is the X– Y plotter you didn’t know you wanted — [hands on]. IEEE Spectrum, 57(9):16–18, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:SSR

[Cas20h] S. Cass. The Sinclair Scientific, remade: a classic calculator was reverse engineered from the bare metal — [hands on]. IEEE Spectrum, 57(6):14–16, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:TPL

[Cas20i] S. Cass. The top programming languages: Our latest rankings put Python on top-again — [careers]. IEEE Spectrum, 57(8):22, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2021:Ha

[Cas21a] S. Cass. Hands on. IEEE Spectrum, 58(3):14–16, March 2021.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2021:Hb

[Cas21b] Stephen Cass. Hands on.

IEEE Spectrum, 58(7):16–18, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2021:HCM

[Cas21c] Stephen Cass. Home clustering made easier: Learn Docker and Kubernetes with the Turing Pi. IEEE Spectrum, 58(11): 16–18, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2021:TPL

[Cas21d] Stephen Cass. Top programming languages: Our eighth annual probe into what’s hot and not. IEEE Spectrum, 58(10):17, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:HFR

[Cas22a] Stephen Cass. Hi-fi, radio, retro, and more: The DIY projects Spectrum readers love. IEEE Spectrum, 59(12):18–20, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:PRK

[Cas22b] Stephen Cass. Passive radar with the KrakenSDR: Spot stuff with TV antennas and a softwaredefined radio. IEEE Spectrum, 59 (11):16–18, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:SSY

[Cas22c] Stephen Cass. SQL should be your second language. IEEE Spectrum, 59(10):20–21, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:SAM

[Cas22d] Stephen Cass. Stefany Allaire: a micromanufacturer carves her niche. IEEE Spectrum, 59(3):19, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:UTT

[Cas22e] Stephen Cass. The ultimate transistor timeline: The transistor’s amazing evolution from point contacts to quantum tunnels. IEEE Spectrum, 59(12):29, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:UTM

[Cas22f] Stephen Cass. Upcycling a Tandy Model 100: The tricky part is its weird display. IEEE Spectrum, 59(10):16–18, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2022:WYO

[Cas22g]  Stephen Cass.  Weave your

own memory: a staggering 64 bits of Apollo-era ferrite-core memory can be yours [hands on]. IEEE Spectrum, 59(8): 14–16, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2023:QLF

[Cas23a] Stephen Cass. 5 questions for Limor Fried: What to do when the ghost in the machine is you. IEEE Spectrum, 60(12):21, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2023:HDY

[Cas23b] Stephen Cass. A hacker’s delight: You’ll either love or hate the flipper zero. IEEE Spectrum, 60(5):18–20, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2023:ISB

[Cas23c] Stephen Cass. I^{2}C strikes back:

Surprises can lurk in even a simple comms protocol. IEEE Spectrum, 60(4):16–18, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2023:YHB

[Cas23d] Stephen Cass. Yugoslavia’s home-brewed microcomputer: The Z80-based Galaksija dared to be different. IEEE Spectrum, 60(8):16–18, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2024:QLR

[Cas24a] Stephen Cass. 5 questions: Libi Rose: Letting people play with the past. IEEE Spectrum, 61 (10):21, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

CASS:2024:BNA

[CAS24b] STEPHEN CASS. A brand new analog computer: Go beyond digital with the THAT. IEEE Spectrum, 61(01):16–18, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2024:H

[Cas24c] Stephen Cass. Hands on. IEEE Spectrum, 61(3):16–18, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2024:MLB

[Cas24d] Stephen Cass. It’s meshtastic! LoRa-based tech brings mesh radio to makers. IEEE Spectrum, 61(6):16–18, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2024:MHH

[Cas24e] Stephen Cass. The most hackable handheld ham radio yet: The UV-K5 can be modded at the click of a mouse. IEEE Spectrum, 61(4):16–18, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2024:RCD

[Cas24f] Stephen Cass. A radar cat detector: When you really need to know what’s inside the box. IEEE Spectrum, 61(8):18–20, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Camsari:2021:DCT

[CD21] Kerem Camsari and Supriyo Datta. Dialogue concerning the two chief computing systems: Imagine yourself on a flight talking to an engineer about a scheme that straddles classical and quantum. IEEE Spectrum, 58(4):30–35, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2022:SGQ

[CFPP22] Charles Q. Choi, Peter Fairley, Tekla S. Perry, and Prachi Patel. Sensors: a guide to the quantum-sensor boom: Atomic scale bolsters sensing revolutions in medicine, tech, and engineering. IEEE Spectrum, 59(6):5–13, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2023:HPS

[CG23] Stephen Cass and Harry Goldstein. How Python swallowed the world: Lessons from compiling top programming languages. IEEE Spectrum, 60 (9):2, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2020:MRR

[Cha20a] R. N. Charette. Managing risk and responsibility during a crisis: The COVID-19 pandemic has resurfaced questions about the government’s role when disaster strikes — [spectral lines]. IEEE Spectrum, 57(5):4, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2020:NON

[Cha20b] R. N. Charette. No one notices the creaky software systems that run the world until they fail. IEEE Spectrum, 57 (9):24–30, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chahat:2021:AMI

[Cha21a] Nacer E. Chahat. An antenna made for an icy, radioactive hell: JPL’s all-metal design can withstand Europa’s brutal environment. IEEE Spectrum, 58 (8):34–39, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chance:2021:LDF

[Cha21b] Frances Chance. Lessons from a dragon fly’s brain: Evolution built a small, fast, efficient neural network in a dragonfly. why not copy it for missile defense? IEEE Spectrum, 58 (8):28–33, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2023:ETH

[Cha23] Robert N. Charette. The EV transition is harder than anyone thinks: Clueless policymakers, skeptical consumers, greedy automakers and the tech isn’t ready either. IEEE Spectrum, 60(4):40–45, April 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2024:CFB

[Cha24] Robert N. Charette. Can flow batteries finally beat lithium?: Nanoparticles may boost energy density enough for EVs. IEEE Spectrum, 61(2):22–27, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chivers:2021:HTA

[Chi21] Tom Chivers. How to train an all-purpose robot: DeepMind is tackling one of the hardest problems for AI. IEEE Spectrum, 58 (10):34–41, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Czerwinski:2021:BAF

[CHM21] Mary Czerwinski, Javier Hernandez, and Daniel McDuff. Building an AI that feels: AI systems with emotional intelligence could learn faster and be more helpful. IEEE Spectrum, 58(5):32–38, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2020:PUT

[Cho20] C. Q. Choi. 3 practical uses for topological photonics: an obscure phenomenon could improve chips, lasers, and quantum circuitry. IEEE Spectrum, 57(4):9–10, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2021:RWA

[Cho21] Charles Q. Choi. 7 revealing ways AIs fail: Neural networks can be disastrously brittle, forgetful, and surprisingly bad at math. IEEE Spectrum, 58 (10):42–47, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2023:IQL

[Cho23] Charles Q. Choi. IBM’s quantum leap: The company will take quantum tech past the 1,000qubit mark in 2023. IEEE Spectrum, 60(1):46–47, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caballar:2024:N

[CJGC24] Rina Diane Caballar, Andrew Jones, Dina Genkina, and Charles Q. Choi. News. IEEE Spectrum, 61(7):5–12, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2021:NLD

[CKG^{+}21] Charles Choi, Karen Kwon, Edd Gent, Samuel K. Moore, and Evan Ackerman. News: The latest developments in technology, engineering, and science [6 items]. IEEE Spectrum, 58(7):6– 13, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clynes:2023:N

[CKX^{+}23] Tom Clynes, Gerald Kulcinski, Tammy Xu, Prachi Patel, Junhong Chen, and Samuel K. Moore. News. IEEE Spectrum, 60(11):6–13, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:2024:SDR

[Cla24] Bryan Clark. Sea drones in the Russia–Ukraine War inspire new tactics: Military strategists see a possible means of defending Taiwan. IEEE Spectrum, 61 (10):28–33, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clynes:2020:BIF

[Cly20] T. Clynes. 5 big ideas for fusion power: Startups, universities, and major companies are vying to commercialize a nuclear fusion reactor. IEEE Spectrum, 57(2):30–37, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clynes:2022:PFZ

[Cly22] Tom Clynes. A pinch of fusion: Zap Energy’s new Zpinch reactor will demonstrate a simpler approach to an elusive goal. IEEE Spectrum, 59 (1):54–55, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clynes:2023:CKHb

[Cly23] Tom Clynes. Can 10,000 kilometers of high-temperature superconducting tape transform fusion power? IEEE Spectrum, 60(8):30–36, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clynes:2024:ESS

[Cly24] Tom Clynes. Energy: an offthe-shelf stellarator: Fast prototyping revives a 70-year-old fusion reactor design. IEEE Spectrum, 61(11):28–39, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cohen:2022:GWW

[Coh22] Gregory Cohen. Gooaall!!!: Why we built a neuromorphic robot to play foosball. IEEE Spectrum, 59(3):44–50, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cook:2021:HWG

[Coo21] Jeremy S. Cook. This huge workbench gives you a hand: a little bit of automation and a lot of space makes projects easier. IEEE Spectrum, 58(8):14–16, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cook:2022:MPS

[Coo22] Jeremy S. Cook. My password is:

> security with a high-resolution spectrometer. IEEE Spectrum, 59(1):16–18, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortada:2021:EQ

[Cor21a] James W. Cortada. The essential question. IEEE Spectrum, 58(8):2, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortada:2021:PFT

[Cor21b] James W. Cortada. Past forward: a tool for modern times. IEEE Spectrum, 58(8):48, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherston:2021:SDS

[CP21] Juliana Cherston and Joseph A.

Paradiso. The smartly dressed spacecraft: Wrapped in sensorrich electronic textiles, space structures could double as scientific instruments. IEEE Spectrum, 58(12):30–37, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cline:2021:PBB

[CPBZ21] Brian Cline, Divya Prasad, Eric Beyne, and Odysseas Zografos. Power from below: Buried interconnects will help save Moore’s Law. IEEE Spectrum, 58(9):

46–51, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Crittenden:2020:UBE

[Cri20] M. Crittenden. Ultralight batteries for electric airplanes. IEEE Spectrum, 57(9):44–49, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Crocker:2020:CPH

[Cro20] S. Crocker. The consensus protocol: How a radical process, invented to build the Arpanet, spawned history’s greatest tool for collaborative work. IEEE Spectrum, 57(8):38–43, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choi:2022:N

[CSB22] Charles Q. Choi, Eliza Strickland, and John Boyd. News. IEEE Spectrum, 59(2):6–50, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeHaro:2022:LC

[De 22] Roger De Haro. Living in color. IEEE Spectrum, 59(1):4, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Degertekin:2024:HUB

[Deg24] F. Levent Degertekin. How ultrasound became ultra small. IEEE Spectrum, 61(4):38–43, April 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2020:MGI

[Der20] D. P. Dern. Merryl Gross, information architect: From cockpits to clinics, user experiences can mean life or death — [careers]. IEEE Spectrum, 57 (11):19, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:Ca

[Der21a] D. P. Dern. Careers. IEEE Spectrum, 58(2):17, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:Cb

[Der21b] D. P. Dern. Careers. IEEE Spectrum, 58(3):17, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:Cc

[Der21c] Daniel Dern. Careers. IEEE Spectrum, 58(5):54, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:ESS

[Der21d] Daniel P. Dern. Engineering secrets: Sandia’s Dennisa Thomas ensures systems work as intended [careers]. IEEE Spectrum, 58(7): 19, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:SSF

[Der21e] Daniel P. Dern. From spacecraft to sensor fusion: Disparate jobs let Iverson Bell create a flexible skill set. IEEE Spectrum, 58(8):17, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:MPN

[Der21f] Daniel P. Dern. Marcellus Proctor: This NASA leader oversees major space projects. IEEE Spectrum, 58(11):20, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:BDH

[Der21g] Daniel P. Dern. S. B. Divya: How a working engineer writes some of the best science fiction around. IEEE Spectrum, 58(6):19, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2021:SSG

[Der21h] Daniel P. Dern. Safety at scale: General Motors’ Ankur Ganguli. IEEE Spectrum, 58 (9):21, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:AGV

[Der22a] Daniel P. Dern. Adam Grosser:this VC is big on transportation alternatives [careers]. IEEE Spectrum, 59(7):17, July 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:CPM

[Der22b] Daniel P. Dern. Craig Partridge: In the midst of Internet traffic for 40 years and loving it. IEEE Spectrum, 59 (11):19, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:HPU

[Der22c] Daniel P. Dern. How Purdue University commercializes its research: Yung-Hsiang Lu helps turn research results into revenue. IEEE Spectrum, 59(5):19, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:MSP

[Der22d] Daniel P. Dern. Magaly Sandoval-Pichardo: Her vast range of experience lets her make the smallest circuits. IEEE Spectrum, 59(4):17, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:MMS

[Der22e] Daniel P. Dern. Magdalene Maluta: She’s building a fleet of personal EVs in Kenya. IEEE Spectrum, 59(9):19, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:MRS

[Der22f] Daniel P. Dern. Maria Rerecich: She tests products for Consumer Reports to ensure they work as claimed. IEEE Spectrum, 59(6):19, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:PJA

[Der22g] Daniel P. Dern. Profile: Julie Angus: Her startup builds autonomous ocean explorers [careers]. IEEE Spectrum, 59(8):17, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2022:PMO

[Der22h] Daniel P. Dern. Profile: Melanie Olsen: She leads autonomous marine tech testing at Reef-Works. IEEE Spectrum, 59 (10):19, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2023:CBS

[Der23a] Daniel P. Dern. Careers: Brij Singh: Electrifying John Deere’s heavy-duty vehicles. IEEE Spectrum, 60(9):21–22, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2023:CET

[Der23b] Daniel P. Dern. Careers: Eileen Tanghal: Her journey from EE to VC. IEEE Spectrum, 60 (6):19–20, June 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2023:CEH

[Der23c] Daniel P. Dern. Careers: Eugene H. Spafford: He’s been fighting cybersecurity threats for decades. IEEE Spectrum, 60 (3):19–20, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2023:CJZ

[Der23d] Daniel P. Dern. Careers:

Jonathan L. Zittrain: The Harvard law professor is an expert on digital technology. IEEE Spectrum, 60(5):21–22, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Downey:2020:MLR

[DHOW20] J. Downey, B. Hilburn, T. O’Shea, and N. West. Machine learning remakes radio. IEEE Spectrum, 57(5):35–39, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeReyes:2020:CCD

[DM20] E. De Reyes and J. Macneill. Can cargo drones solve air freight’s logjams? A drone startup says its big vertical-takeoff flier would be quick to land, load, and take off again. IEEE Spectrum, 57(6):30–35, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dolgin:2023:AGJ

[Dol23a] Elie Dolgin. Arthritis gets a jolt: 2023 will see the results of nerve stimulation for autoimmune disorders. IEEE Spectrum, 60(1):48–49, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dolgin:2023:DTB

[Dol23b] Elie Dolgin. Different types of brain stimulation may reduce cravings. IEEE Spectrum, 60 (8):24–29, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2020:DPA

[Dum20] M. Dumiak. Data project aims to organize scientific records: Deeptime Digital Earth will link hundreds of bespoke databases — [news]. IEEE Spectrum, 57 (3):9–10, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2022:RTV

[Dum22] Michael Dumiak. A road test for vehicle-to-grid tech: Utrecht leads the world in using EVs for grid storage. IEEE Spectrum, 59(8):20–25, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2023:ECE

[Dum23] Michael Dumiak. Exascale comes to Europe: Germany will host JUPITER, Europe’s entry into the realm of exascale supercomputing. IEEE Spectrum, 60 (1):50–51, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta:2022:EBE

[Dut22] Barun Dutta. Eavesdropping on the brain: With 10,000 electrodes, this neural implant senses more than ever before.

IEEE Spectrum, 59(6):30–35, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ehnebuske:2023:H

[Ehn23]  David Ehnebuske.  Hands

on. IEEE Spectrum, 60(9): 18–20, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eisler:2023:FSC

[Eis23] Matthew N. Eisler. False starts:

The checkered history of vehicleto-grid power. IEEE Spectrum, 60(4):46–53, April 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2020:BSG

[Fai20a] P. Fairley. Bangladesh scrambles to grow power supply — [news]. IEEE Spectrum, 57 (2):6–7, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2020:IGV

[Fai20b] P. Fairley. India grapples with vast solar park — [news]. IEEE Spectrum, 57(2):8–9, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2021:GSK

[Fai21] Peter Fairley. Off-grid solar’s killer app: Solar pumps, batteries, and microcredit are triggering an African agricultural renaissance. IEEE Spectrum, 58(6):44–49, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2023:BBG

[Fai23a] Peter Fairley. Buoyant behemoths: The global race is on to tap potent winds far offshore. IEEE Spectrum, 60(11): 30–37, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2023:MGH

[Fai23b] Peter Fairley. Making green hydrogen a reality Down Under: Proposals to make hydrogen from renewable electricity in Australia exceed the country’s generating capacity. IEEE Spectrum, 60(1):24–27, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2023:THE

[Fai23c] Peter Fairley. The transformer: How energy modeler Jesse Jenkins helped Congress get serious about the power grid. IEEE Spectrum, 60(12):30–33, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2024:HVD

[Fai24a] Peter Fairley. High-voltage DC power roars ashore in Europe: Scotland’s next-generation transmission network foreshadows a global grid transformation. IEEE Spectrum, 61(01):28–31, January 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2024:PTS

[Fai24b] Peter Fairley. Powering Taiwan’s Silicon Shield. IEEE Spectrum, 61(10):22–45, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farina:2023:RWG

[FBMI23] Dario Farina, Etienne Burdet, Carsten Mehring, and Jaime Ibáñez. Roboticists want to give you a third arm: Unused bandwidth in neurons can be tapped to control extra limbs. IEEE Spectrum, 60(3):22–46, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fishman:2024:CTA

[Fis24] Ted C. Fishman. Climate tech:

The AI boom rests on billions of tonnnes of concrete: Why tech’s climate pledges require the reinvention of concrete. IEEE Spectrum, 61(11):58–69, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:2021:NWW

[Fit21] Alissa M. Fitzgerald. A new wildfire watchdog: Alerts about forest fires shouldn’t depend on pets smelling smoke. we need smart infrastructure, and that needs zero-power sensors. IEEE Spectrum, 58(11):38–43, November 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Fork:2021:HEC

[FK21] David Fork and Ross Koningstein. How engineers can disrupt climate change. IEEE Spectrum, 58(7):24–29, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ferrari:2023:WCM

[FOSMD23] Maximiliano Ferrari, Ben Ollis, Michael Starke, and Arturo Massol-Deya. The wellconnnected microgried: a network of microgrids can withstand severestorms. IEEE Spectrum, 60 (10):36–43, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foust:2021:TWM

[Fou21] J. Foust. Three ways to the Moon: NASA will soon narrow the possible strategies for landing people on the Moon. IEEE Spectrum, 58(1):44–47, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foust:2022:NSL

[Fou22] Jeff Foust. NASA’s space launch system will lift off: But with rival rockets readying for flight, the ultimate value of SLS is murky. IEEE Spectrum, 59 (1):36–39, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2021:SAC

[Fra21] Steven J. Frank. State of the art: This convolutional neural network can tell you whether a painting is a fake. IEEE Spectrum, 58(9):26–31, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frumer:2020:SSL

[Fru20] Y. Frumer. The short, strange life of the first friendly robot: Japan’s Gakutensoku was a giant pneumatic automaton that toured through Asia until it mysteriously disappeared. IEEE Spectrum, 57(6):42–48, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Funk:2020:EER

[Fun20] J. Funk. Expect evolution, not revolution: Despite the hype, artificial intelligence will take years to significantly boost economic productivity. IEEE Spectrum, 57(3):30–35, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2020:WFY

[GA20] E. Guizzo and E. Ackerman.

$74,500 will fetch you a spot: For the price of a luxury car, you can now buy a very smart, very capable, very yellow robot dog. IEEE Spectrum, 57(8):11, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenberg:2020:ASB

[GAG20] J. A. Greenberg, A. Ashok, and M. E. Gehm. The all-seeing baggage scanner. IEEE Spectrum, 57(7):22–27, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallucci:2020:AMR

[Gal20a] M. Gallucci. Altarock melts rock for geothermal wells: Millimeter waves could help us dig deeper and faster than with traditional drills — [news]. IEEE Spectrum, 57(3):8–9, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallucci:2020:WDG

[Gal20b] M. Gallucci. What to do with 177 giant tanks of radioactive sludge. IEEE Spectrum, 57 (5):24–33, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallucci:2021:ASA

[Gal21a] M. Gallucci. The ammonia solution: Ammonia engines and fuel cells in cargo ships could slash their carbon emissions. IEEE Spectrum, 58(3):44–50, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallucci:2021:CSF

[Gal21b] M. Gallucci. The carbon-sucking fans of West Texas: It’s not enough to slash greenhouse gas emissions. Experts say we need direct-air capture. IEEE Spectrum, 58(1):48–49, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ganju:2024:PED

[Gan24] Vineet Ganju. Piezoelectrics enable displays to provide both high-quality audio and touch feedback. IEEE Spectrum, 61 (3):22–27, March 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garisto:2020:NPA

[Gar20]  D. Garisto.  New particle-

accelerating tech passes test: Energy recovery linear accelerators could reduce costs and conserve power — [news]. IEEE Spectrum, 57(3):6–7, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:N

[GAS^{+}24] Edd Gent, Evan Ackerman, Matthew S. Smith, Tom Clynes, and Katherine Bourzac. News. IEEE Spectrum, 61(6):6–13, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2022:N

[GDUF22] Edd Gent, Michael Dumiak, Lawrence Ulrich, and Peter Fairley. News. IEEE Spectrum, 59 (10):6–13, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2022:LSL

[Gen22] Edd Gent. Lessons from a second life > before Meta, Philip Rosedale created an online universe. IEEE Spectrum, 59 (1):19, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2023:CJI

[Gen23a] Edd Gent. Careers: Jim Vanns: The Industrial Light & Magic Engineer wins a technical Oscar. IEEE Spectrum, 60(10): 19–20, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2023:CMP

[Gen23b] Edd Gent. Careers: Morgan Pope: This Disney roboticist brings superheroes to life. IEEE Spectrum, 60(12):19–20, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2023:CPG

[Gen23c] Edd Gent. Careers: Peter Gracey: The former Royal Navy officer advocates for LGBTQ+ engineers. IEEE Spectrum, 60 (8):19–20, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2023:CSC

[Gen23d] Edd Gent. Careers: Sushama Chakraverty: How this UI designer finds the fun factor that makes video games so compelling. IEEE Spectrum, 60 (7):19–20, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2023:CMU

[Gen23e] Edd Gent. A cryptocurrency for the masses or a universal ID?: Worldcoin aims to scan all the world’s eyeballs. IEEE Spectrum, 60(1):42–57, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Genkina:2024:DSC

[Gen24a] Dina Genkina. Don’t start a career as an AI prompt engineer: AI will take your job. IEEE Spectrum, 61(5):30–34, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Genkina:2024:NOS

[Gen24b] Dina Genkina. News: Opensource hardware secures a win: sprung from RISC-V’s roots, the first open security chip hits the market. IEEE Spectrum, 61(4):6–13, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Genkina:2024:QTS

[Gen24c] Dina Genkina. Quantum tech:

Squeezing an optical atomic clock into a briefcase: a mindset shift propelled the leap from basic science to invention. IEEE Spectrum, 61(11):44–70, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CAS

[Gen24d] Edd Gent. Careers: Aakhilesh Singhania: This Bosch engineer speeds hybrid race cars to the finish line. IEEE Spectrum, 61(7):19–20, July 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CAE

[Gen24e] Edd Gent. Careers: Arthur Erickson: This aerospace engineer builds crop-spraying drones for farmers. IEEE Spectrum, 61(6):19–20, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CEC

[Gen24f] Edd Gent. Careers: Erika Cruz: She keeps Whirlpool’s machines spinning smoothly. IEEE Spectrum, 61(9):19–20, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

GENT:2024:CGB

[GEN24g] EDD GENT. Careers: Greta Bekeryte: Her journey from cleaning offices to being one of the top women in the EV industry. IEEE Spectrum, 61 (01):19–20, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CMS

[Gen24h] Edd Gent. Careers: Manan Suri: His brain-inspired computers run AI on sensors and drones. IEEE Spectrum, 61(2):19–20, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CPB

[Gen24i] Edd Gent. Careers: Patryk Borkowski: He has a cool gig at Unilever’s Ice Cream Innovation Center. IEEE Spectrum, 61(3):19–20, March 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CSC

[Gen24j] Edd Gent. Careers: Susana Contrera: a network engineer who keeps Meta’s AI infrastructure humming. IEEE Spectrum, 61(5):19–20, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:CZR

[Gen24k] Edd Gent. Careers: Zach Rattner: His startup’s AI tool makes moving day easier. IEEE Spectrum, 61(4):20–21, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:IDE

[Gen24l] Edd Gent. Irene Degl’Innocenti:

This engineer keeps CERN’s particle accelerators on track. IEEE Spectrum, 61(8):21–22, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gent:2024:LEN

[Gen24m] Edd Gent. Laser embeds nanoscale device in silicon. IEEE Spectrum, 61(9):14–15, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldman:2023:AHH

[GH23] Josh Goldman and Kurt House. AI hunts for hidden minerals: Machine learning is uncovering hoards of vital EV battery metals. IEEE Spectrum, 60(6):22–27, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garisto:2022:N

[GHG^{+}22] Dan Garisto, Mark Harris, Dina Genkina, Rahul Rao, and Prachi Patel. News. IEEE Spectrum, 59 (12):6–13, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibbs:2021:HVW

[Gib21] W. W. Gibbs. This is how to vaccinate the world: We can manufacture and distribute enough doses to protect humanity from COVID-19. IEEE Spectrum, 58 (1):32–37, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2024:WMR

[GK24] Erico Guizzo and Randi Klett. Where’s my robot butler?: Here’s how we could finally build humanoid robots that do all our domestic chores. IEEE Spectrum, 61(11):40–43, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Genkina:2024:N

[GMB^{+}24] Dina Genkina, Andrew Moseman, Katherine Bourzac, Laura Hautala, and Matthew Hutson. News. IEEE Spectrum, 61 (10):5–12, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goebel:2022:MMW

[GO22] Dan M. Goebel and David Oh.

Mission to a metal world: In August, NASA will launch a probe to study a strange metallic asteroid called Psyche. IEEE Spectrum, 59(2):24–29, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gobron:2023:NTT

[Gob23] Robert Gobron. Nanostructures transform tiny cameras: Thin semiconductor metalenses are finally moving into consumers’ hands. IEEE Spectrum, 60 (7):22–27, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2021:YNI

[Gol21] Harry Goldstein. Your new IEEE Spectrum website: An enhanced member experience is central to the new Spectrum website. IEEE Spectrum, 58 (9):6, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:ACH

[Gol22a] Harry Goldstein. AI can help make recycling better: But only humans can solve the plastics problem. IEEE Spectrum, 59(7): 2, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:AHH

[Gol22b] Harry Goldstein. AI helps humans level up: Chess players and programmers now have access to tools that can improve skills and unleash creativity. IEEE Spectrum, 59(11):2, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:DCE

[Gol22c] Harry Goldstein. The device that changed everything: Transistors are civilization’s invisible infrastructure. IEEE Spectrum, 59(12):2, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:HWC

[Gol22d] Harry Goldstein. How we celebrate engineers: The Institute focuses on fascinating people and the technology they create. IEEE Spectrum, 59 (9):2, September 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:NAU

[Gol22e] Harry Goldstein. Nothing about us without us: Put the user at the center of assistive technology design. IEEE Spectrum, 59(10):2, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2022:WVT

[Gol22f] Harry Goldstein. What V2G tells us about EVs and the grid: Vehicle-to-grid technology adds another layer of complexity to the electric-vehicle transition. IEEE Spectrum, 59(8):2, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:AEA

[Gol23a] Harry Goldstein. AI everywhere, all at once: It’s time to get serious about regulating artificial intelligence. IEEE Spectrum, 60(6):2, June 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:FHM

[Gol23b] Harry Goldstein. Fusion is having a moment. IEEE Spectrum, 60(8):2, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:KIR

[Gol23c] Harry Goldstein. Kicking it with robots: RoboCup and the new Robots Guide are gateways to the wonderful world of robots. IEEE Spectrum, 60(7):

2, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:LBS

[Gol23d] Harry Goldstein. Looking back to see ahead: Technologies decades in the making finally get their star turn in 2023. IEEE Spectrum, 60(1):4, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:OHC

[Gol23e] Harry Goldstein. Our hands-on citizen scientist: David Schneider’s DIY projects unveil nature’s secrets. IEEE Spectrum, 60(10):2, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:PAH

[Gol23f] Harry Goldstein. The portrait artist: How Tekla S. Perry chronicles consequential careers. IEEE Spectrum, 60(5):2, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:RHW

[Gol23g] Harry Goldstein. Robots and the humans who make them: Our chatbot podcast features roboticists in conversation with each other. IEEE Spectrum, 60 (11):2, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2023:SSE

[Gol23h] Harry Goldstein. The staggering scale of the EV transition: One engineer grapples with the challenges ahead. IEEE Spectrum, 60(4):2, April 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:CBI

[Gol24a] Harry Goldstein. The chasm between imagination and feasibility: Space-based solar power started as sci-fi. And it still is. IEEE Spectrum, 61(6):2, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:DTC

[Gol24b] Harry Goldstein. Defending Taiwan with chips and drones: Taiwan doubles down on its Silicon Shield strategy while the U.S. promises a storm of drones. IEEE Spectrum, 61(10):2, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:DVB

[Gol24c] Harry Goldstein. The doyen of the valley bids adieu: Tekla S. Perry blazed a trail for women tech journalists. IEEE Spectrum, 61(8):2, August 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

GOLDSTEIN:2024:FWS

[GOL24d] HARRY GOLDSTEIN. The future we saw coming is now:

Technologies we’ve been following for years will be huge in 2024. IEEE Spectrum, 61(01):4, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:GTW

[Gol24e] Harry Goldstein. The greening of transportation: We’ll need new inventions and novel adaptations to decarbonize the sector. IEEE Spectrum, 61(2):2, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:MCT

[Gol24f] Harry Goldstein. Making connections: Travels with AI made possible by Bob Kahn’s historic contributions. IEEE Spectrum, 61(5):2, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:MMY

[Gol24g] Harry Goldstein. A match made in Yorktown Heights: How one artist fostered a visually stunning collaboration. IEEE Spectrum, 61(9):2, September 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:MCS

[Gol24h] Harry Goldstein. Moore on chip scaling: Scaling compute to satiate AI’s appetite will take extreme measures. IEEE Spectrum, 61(7):2, July 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:RPC

[Gol24i] Harry Goldstein. Reinventing a patent case: Cob Burandt is still down and out in Ham Lake, but his attorneys have just begun to fight — again. IEEE Spectrum, 61(11):2, November 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2024:SFH

[Gol24j] Harry Goldstein. Sci-fi and hifi: The microstory of a storyteller who delivers both. IEEE Spectrum, 61(3):2, March 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goodrich:2023:APE

[Goo23a] Joanna Goodrich. Arjun Pillai: The entrepreneur shares lessons learned on his road to success. IEEE Spectrum, 60 (1):18–19, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goodrich:2023:PF

[Goo23b] Joanna Goodrich. Past forward.

IEEE Spectrum, 60(12):68, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guedes-Pinto:2022:AFM

[GP22] Paulo Guedes-Pinto. An axialflux motor for an electrified world: It combines the best of two motor designs to save weight and energy. IEEE Spectrum, 59(4):38–43, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2020:YNS

[Gui20] E. Guizzo. Your next salad could be grown by a robot. IEEE Spectrum, 57(1):34–35, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hampson:2020:CDR

[Ham20a] M. Hampson. COVID-19 disrupts research plans: The coronavirus pandemic has interrupted fieldwork and lab experiments — [news]. IEEE Spectrum, 57(6):9–10, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hampson:2020:MMJ

[Ham20b] M. Hampson. Messages from a MicroLED — [journal watch]. IEEE Spectrum, 57(6):7, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hampson:2020:MMM

[Ham20c] M. Hampson. Mobile malware mutates — [news]. IEEE Spectrum, 57(7):08, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hampson:2020:MBL

[Ham20d] M. Hampson. Motorized backpack lightens the load — [journal watch]. IEEE Spectrum, 57(10):7, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamerly:2021:FDL

[Ham21] Ryan Hamerly. The future of deep learning is photonic: Reducing the energy needs of neural networks might require computing with light. IEEE Spectrum, 58(7):30–47, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harris:2021:MTJ

[Har21] M. Harris. Military tests that jam and spoof GPS signals are an accident waiting to happen. IEEE Spectrum, 58(2): 22–27, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harris:2022:RST

[Har22] Mark Harris. The radical scope of Tesla’s data hoard: Every Tesla is providing reams of sensitive data about its driver’s life. IEEE Spectrum, 59(10): 40–45, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2020:ECP

[Has20a] S. Hassler. Ending the COVID-19 pandemic: Embracing notech solutions will make hightech solutions possible — [spectral lines]. IEEE Spectrum, 57 (12):6, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2020:GNA

[Has20b] S. Hassler. Google has Nest.

Amazon has Echo and Alexa. Savant Systems has GE lighting!

The luxury home-automation company has its sights set on high-quality mass-market home networking — [spectral lines]. IEEE Spectrum, 57(7):04, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hutson:2023:N

[HCP^{+}23] Matthew Hutson, Charles Q. Choi, Prachi Patel, Lawrence Ulrich, and Michelle Hampson. News. IEEE Spectrum, 60 (5):7–15, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hecht:2020:SOF

[Hec20] J. Hecht. Same old fibers, recordbreaking speeds: Equipment in the lab can carry 100 million Zoom videos — [news]. IEEE Spectrum, 57(10):9–10, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbst:2020:FO

[Her20] L. Herbst. Fishing for oil. IEEE Spectrum, 57(6):36–50, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huitema:2022:ITM

[HF22] Edzer Huitema and Ian French. E ink’s technicolor moment: The road to color E-paper took two decades. IEEE Spectrum, 59 (2):30–35, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herr:2024:DCS

[HH24] Anna Herr and Quentin Herr.

A data center in a shoebox: IMEC’s plan to use superconductors to shrink computers. IEEE Spectrum, 61(6):37–41, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:ADC

[Hig20a] S. Higginbotham. Anonymous by design — [CrossTalk]. IEEE Spectrum, 57(9):22–23, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:DDP

[Hig20b] S. Higginbotham. Data-driven power — [Internet of Everything]. IEEE Spectrum, 57(8):20, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:ICP

[Hig20c] S. Higginbotham. The IoT’s consent problem — [CrossTalk]. IEEE Spectrum, 57(7):20, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:MLE

[Hig20d] S. Higginbotham. Machine learning on the edge — [Internet of Everything]. IEEE Spectrum, 57(1):20, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:MSV

[Hig20e] S. Higginbotham. Making sure votes count — [Internet of Everything]. IEEE Spectrum, 57 (10):18, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:NII

[Hig20f] S. Higginbotham. Network included — [Internet of Everything]. IEEE Spectrum, 57(11): 22–23, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:PVP

[Hig20g] S. Higginbotham. Pandemic vs. privacy — [CrossTalk]. IEEE Spectrum, 57(6):20, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:WAD

[Hig20h] S. Higginbotham. We all deserve broadband — [Internet of Everything]. IEEE Spectrum, 57(5):22, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:WHG

[Hig20i] S. Higginbotham. What 5G hype gets wrong — [Internet of Everything]. IEEE Spectrum, 57(3):22, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:WFL

[Hig20j] S. Higginbotham. Wi-Fi’s long goodbye: Crosstalk. IEEE Spectrum, 57(2):19, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:C

[Hig21] S. Higginbotham. CrossTalk.

IEEE Spectrum, 58(3):20–21, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hochman:2022:NYW

[Hoc22] Brian Hochman. 1950s New York was a wiretapper’s dream: Eavesdropping flourished after WW II, aided by legal loopholes, clever hacks, and private ears. IEEE Spectrum, 59 (8):38–45, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2020:WBR

[HP20] S. Higginbotham and M. Pesce.

Who’s behind that robot? — [CrossTalk]. IEEE Spectrum, 57 (12):24–25, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMa

[HP21a] S. Higginbotham and M. Pesce.

Internet of Everything: Macro micro. IEEE Spectrum, 58(1): 18–19, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMb

[HP21b] S. Higginbotham and M. Pesce. Internet of Everything: Macro micro. IEEE Spectrum, 58(2): 20–21, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMe

[HP21c] Stacey Higginbotham and Mark Pesce. Internet of Everything: Macro & micro. IEEE Spectrum, 58(7):22–23, July 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMf

[HP21d] Stacey Higginbotham and Mark Pesce. Internet of Everything: Macro & micro. IEEE Spectrum, 58(8):20–21, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMc

[HP21e] Stacey Higginbotham and Mark Pesce. Internet of Everything: Macro micro. IEEE Spectrum, 58(4):20–21, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higginbotham:2021:IEMd

[HP21f] Stacey Higginbotham and Mark Pesce. Internet of Everything: Macro micro. IEEE Spectrum, 58(6):22–23, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horowitz:2021:N

[HPHU21] Brian T. Horowitz, Prachi Patel, Jeremy Hsu, and Lawrence Ulrich. News. IEEE Spectrum, 58(5):6–13, May 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hecht:2022:IUM

[HPK22] Jeff Hecht, Ned Potter, and Michael Koziol. Inside the Universe Machine: IEEE Spectrum explores the Webb Telescope’s groundbreaking engineering. IEEE Spectrum, 59(9): 22–30, September 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Holden:2023:GRC

[HS23] Richard Holden and Elias Stein.

Get ready for corporate digital currencies: Facebook failed, but another tech giant might soon pull it of. IEEE Spectrum, 60 (12):34–39, December 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsu:2020:CAH

[Hsu20a] J. Hsu. Can AI hiring systems be made antiracist? makers and users of AI-assisted recruiting software reexamine the tools’ development and how they’re used — [news]. IEEE Spectrum, 57 (9):9–11, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsu:2020:DCT

[Hsu20b] J. Hsu. The dilemma of contacttracing apps: Can this crucial technology be both effective and private? IEEE Spectrum, 57 (10):56–59, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsu:2020:SHW

[Hsu20c] J. Hsu. Smart headphones warn of nearby cars — [journal watch]. IEEE Spectrum, 57(2):8, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hubert:2024:WBS

[Hub24] Bert Hubert. Why bloat is still software’s biggest vulnerability: a 2024 plea for lean software. IEEE Spectrum, 61 (4):22–50, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hussain:2024:SSI

[Hus24] Asim Hussain. Software sucks, but it doesn’t have to: How to make leaner, greener software. IEEE Spectrum, 61(4):2, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hutson:2020:MBM

[Hut20] M. Hutson. The mess behind the models: Too many of the COVID-19 models led policymakers astray. Here’s how tomorrow’s models will get it right. IEEE Spectrum, 57(10): 30–35, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hutson:2021:OAI

[Hut21] M. Hutson. The opacity of artificial intelligence makes it hard to tell when decision-making is biased. IEEE Spectrum, 58(2): 40–45, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haber:2020:MET

[HVW20] N. Haber, C. Voss, and D. Wall.

Making emotions transparent: Google Glass helps autistic kids understand facial expressions through augmented-reality therapy. IEEE Spectrum, 57(4): 46–52, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jaffe:2022:SPD

[Jaf22] Paul Jaffe. Spooky power at a distance: Researchers have beamed substantial amounts of energy at distances over 1 kilometer. IEEE Spectrum, 59 (6):36–41, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jagula:2022:BVS

[Jag22] Dexter Jagula. A boom with a view: The satellite-imaging industry is exploding. Here’s how to take advantage of it. IEEE Spectrum, 59(3):38–43, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jankovsky:2024:FHS

[JAR24] Amy Jankovsky, Christine Andrews, and Bill Rogers. Fly the hybrid skies: NASA, GE Aerospace, and Boeing are collaborating on a hybrid-electric airliner. IEEE Spectrum, 61 (2):28–34, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jessen:2021:SSG

[Jes21]  Gregg H. Jessen.  The su-

percharged semiconductor: Gallium oxide could make powerful radios and switch thousands of volts. IEEE Spectrum, 58 (4):36–42, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jhunjhunwala:2022:WIC

[JJS22] Ashok Jhunjhunwala, Kaushal Kumar Jha, and Anson Sando. Weaning India from coal: How to hasten the country’s transition to solar and wind power. IEEE Spectrum, 59(7):40–45, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:FF

[Jon22a] Willie D. Jones. Figure from fiction. IEEE Spectrum, 59 (6):14–15, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:HW

[Jon22b] Willie D. Jones. Help not wanted. IEEE Spectrum, 59 (8):12–13, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:MRL

[Jon22c] Willie D. Jones. Megatruck runs on the lightest gas. IEEE Spectrum, 59(7):12–13, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:RAD

[Jon22d] Willie D. Jones. Revamping an art deco masterpiece. IEEE Spectrum, 59(9):14–15, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:RHD

[Jon22e] Willie D. Jones. A rich harvest in the desert. IEEE Spectrum, 59 (11):14–15, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:RCP

[Jon22f] Willie D. Jones. RoboCup class picture. IEEE Spectrum, 59 (10):14–15, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:RR

[Jon22g] Willie D. Jones. A Roomba for rivers. IEEE Spectrum, 59(5):14–15, May 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2022:SNF

[Jon22h] Willie D. Jones. Shot of nuclear fusion. IEEE Spectrum, 59 (12):14–15, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:QAL

[Jon23a] Willie D. Jones. 5 questions for Andrew Leland: The history of assistive technology for the blind. IEEE Spectrum, 60(10):21, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:APT

[Jon23b] Willie D. Jones. Amazon plans to take home delivery to new heights. IEEE Spectrum, 60 (1):16–17, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:ABB

[Jon23c] Willie D. Jones. Autodidacts build brain-controlled prosthetic. IEEE Spectrum, 60(5):16–17, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:EFQ

[Jon23d] Willie D. Jones. England’s fission quest. IEEE Spectrum, 60 (9):16–17, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:GG

[Jon23e] Willie D. Jones. Guided gliders. IEEE Spectrum, 60(12): 14–15, December 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:ISC

[Jon23f] Willie D. Jones. India’s solar center. IEEE Spectrum, 60 (10):14–15, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:IG

[Jon23g] Willie D. Jones. Intentional grounding. IEEE Spectrum, 60 (3):14–15, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:JB

[Jon23h] Willie D. Jones. Juice box.

IEEE Spectrum, 60(4):14–15, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:NDS

[Jon23i] Willie D. Jones. Neutrino detector to solve Universe’s mysteries. IEEE Spectrum, 60(7):14–15, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:SF

[Jon23j] Willie D. Jones. Space flight.

IEEE Spectrum, 60(11):14–15, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2023:WNW

[Jon23k] Willie D. Jones. Where no wind turbine has gone before. IEEE Spectrum, 60(6):14–15, June 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2024:BMA

[Jon24a] Andrew Jones. Back to the moon, almost: NASA gears up for the world’s first crewed deepspace mission in 50 years. IEEE Spectrum, 61(01):50–53, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2024:MMM

[Jon24b] Willie D. Jones. Municipal model maker. IEEE Spectrum, 61(6):14–15, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2024:RBS

[Jon24c] Willie D. Jones. Robotic bees’ swarm autonomy. IEEE Spectrum, 61(7):14–15, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jordahn:2021:SWS

[Jor21] Henrik Jordahn. A skydiver who sews. IEEE Spectrum, 58 (12):2, December 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2021:LDT

[JPWB21] Andrew Jones, Prachi Patel, Emily Waltz, and Natasha Bajema. The latest developments in technology, engineering, and science: News. IEEE Spectrum, 58(10):6–13, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kastrup:2023:SDA

[Kas23] Bernardo Kastrup. Softwaredefined architecture: Learn system-level design with this dual-CPU computer. IEEE Spectrum, 60(11):16–18, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kroposki:2024:PPR

[KH24] Benjamin Kroposki and Andy Hoke. A path to 100 percent renewable energy: Grid-forming inverters will give us the grid we need now. IEEE Spectrum, 61(5):50–57, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kim:2022:GSK

[Kim22] Dong-Won Kim. The godfather of South Korea’s chip industry: Kim Choong-Ki’s engineer’s mind helped make the country a semiconductor superpower. IEEE Spectrum, 59 (10):32–38, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kovatchev:2021:CAP

[KK21] Boris Kovatchev and Anna Kovatcheva. Creating the artificial pancreas: This wearable device senses blood glucose and administers insulin accordingly. IEEE Spectrum, 58(12):38–43, December 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Koningstein:2024:GQN

[Kon24] Ross Koningstein. Google’s quiet nuclear quest: For a decade, Google has advanced fusion R&D and other frontiers. IEEE Spectrum, 61(12):32–37, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). See a research summary of the cold fusion review, see [BCM^{+}19].

Koziol:2020:GSS

[Koz20a] M. Koziol. Geostationary satellites for small jobs. IEEE Spectrum, 57(1):48–58, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2020:HRH

[Koz20b] M. Koziol. Helping radios hear themselves — [news]. IEEE Spectrum, 57(1):12–13, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2020:NES

[Koz20c] M. Koziol. New encryption strategy passes early test: Ghost polarization harnesses ultrafast fluctuations that occur in a light wave. IEEE Spectrum, 57(7):11, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2021:SIW

[Koz21a] M. Koziol. A small island waits on big data rates: Satellite operators salivate over access to Google’s Equiano cable, but St. Helena’s telecom monopoly stands in the way. IEEE Spectrum, 58(1):50–51, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2021:CFM

[Koz21b] Michael Koziol. The clash over 5G’s first mile: The wireless industry is divided on Open RAN’s goal to make network components interoperable. IEEE Spectrum, 58(5):40–46, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2023:QAT

[Koz23a] Michael Koziol. 5 questions for Anton Troynikov: His company’s creation identifies the art behind AI-generated images. IEEE Spectrum, 60(5):23, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2023:QCT

[Koz23b] Michael Koziol. 5 questions for Chris Taylor: Aalyria’s CEO on using lasers to close the digital divide. IEEE Spectrum, 60(3):21, March 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2023:QMW

[Koz23c] Michael Koziol. 5 questions for Molly White: The “Web3 is going just great” creator on why it isn’t. IEEE Spectrum, 60(4):21, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2023:RAL

[Koz23d] Michael Koziol. The return of the airship: LTA Research’s Pathfinder 1 will begin test flights in 2023. IEEE Spectrum, 60(1):28–29, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2024:QDF

[Koz24a] Michael Koziol. 5 questions for David Frankel: One man’s quixotic quest to stop robocalls. IEEE Spectrum, 61(5):21, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2024:QSF

[Koz24b] Michael Koziol. 5 questions for Seth Fraden: How to boot up a new engineering program. IEEE Spectrum, 61(4):19, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2024:QVC

[Koz24c] Michael Koziol. 5 questions for Vince Cate: How the AI boom has been a windfall for the island of Anguilla. IEEE Spectrum, 61(2):21, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2024:WFB

[Koz24d] Michael Koziol. Wi-fi’s big bet on reliability: The next generation of the wireless tech retackles a common complaint. IEEE Spectrum, 61(01):48–49, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahn:2024:GI

[KP24] Robert Kahn and Tekla S.

Perry. The great interconnector. IEEE Spectrum, 61(5):36–41, May 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2023:N

[KPCA23] Michael Koziol, Prachi Patel, Charles Q. Choi, and Evan Ackerman. News. IEEE Spectrum, 60(10):5–13, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koziol:2021:N

[KPG^{+}21] Michael Koziol, Ned Potter, Edd Gent, Ludovico Rella, and Prachi Patel. News. IEEE Spectrum, 58 (12):6–50, December 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kuhlbusch:2024:NFP

[Kuh24] Tim Kuhlbusch. A no-fuss particle detector: an inexpensive and robust way to see the invisible. IEEE Spectrum, 61 (10):16–18, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2020:RLM

[Kum20] Jean Kumagai. The return of the lithium-metal battery. IEEE Spectrum, 57(1):30–31, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2021:MBL

[Kum21] J. Kumagai. Momentum builds for lithium-ion battery recycling: The goal is to prevent thousands of tons of spent batteries from going to waste. IEEE Spectrum, 58(1):52–53, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levi:2022:NCN

[LA22] Anthony F. J. Levi and Gabriel Aeppli. The naked chip: No trade secret or hardware trojan can hide from ptychographic Xray laminography. IEEE Spectrum, 59(5):38–43, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laursen:2023:N

[Lau23a] Lucas Laursen. News. IEEE Spectrum, 60(8):5–13, August 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laursen:2023:NMN

[Lau23b] Lucas Laursen. No more no service: Cellphones will increasingly text via satellite. IEEE Spectrum, 60(1):52–55, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laursen:2024:AIC

[Lau24a] Lucas Laursen. Artificial intelligence: Can we automate fureka moments?: That’s the goal of this Swiss company’s AIpowered invention tech. IEEE Spectrum, 61(11):24–27, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laursen:2024:POH

[Lau24b] Lucas Laursen. Paris Olympics host a new event: Algorithmic video surveillance: Security olympics spin-offs are coming for you, and you, and you. IEEE Spectrum, 61(01): 34–37, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazarev:2021:H

[Laz21] A. Lazarev. Hands on. IEEE Spectrum, 58(1):12–14, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazarus:2022:RST

[Laz22]  Mitchell Lazarus.  Radio-

spectrum turf wars. IEEE Spectrum, 59(7):34–39, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lau:2020:DSC

[LCBH20] J. H. Lau, T. Cohn, T. Baldwin, and A. Hammond. “Deepspeare” crafted Shakespearean verse that few readers could distinguish from the real thing. IEEE Spectrum, 57(5):40–53, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Li:2022:HAG

[LDO22] Qi Li, Yin Ding, and Jorel Olan. How audio is getting its groove back: Deep learning is delivering the century-old promise of truly realistic sound reproduction. IEEE Spectrum, 59 (10):46–52, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lee:2020:SDP

[Lee20] P. T. Lee. The software-defined power grid: How software and sensors are bringing century-old grid technology into the modern age. IEEE Spectrum, 57(7):40– 46, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levick:2020:RMW

[Lev20] E. Levick. A robot is my Wingman. IEEE Spectrum, 57 (1):32–56, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levick:2021:SST

[Lev21] Ewen Levick. A supercold sapphire ticks more regularly than anything else on the planet: The most precise timekeeper in the world. IEEE Spectrum, 58(7):40– 46, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:2020:MHP

[Lew20] P. Lewis. Make a hack-proof garage door opener: a new breakout board offers cryptographic security — [hands on]. IEEE Spectrum, 57(3):16–18, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:2023:HFD

[Lew23] Pete G. Lewis. Hi-fi DIY ear defense: Lower the volume without muffling sounds. IEEE Spectrum, 60(12):16–18, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liu:2020:AVL

[LG20] S. Liu and J. Gaudiot. Autonomous vehicles lite selfdriving technologies should start small, go slow. IEEE Spectrum, 57(3):36–49, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levine:2024:UBD

[LH24] Sergey Levine and Karol Hausman. Using big data to make robots more capable. IEEE Spectrum, 61(2):42–46, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Landreneau:2023:PMS

[LSM23] John Landreneau, Lauren Schricker, and John MacNeill. Plugging in the Moon: To supply outposts sunk in the lunar night, solar power must come from the South Pole. IEEE Spectrum, 60 (8):46–49, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2020:BEF

[Luc20a] R. W. Lucky. Back to the elusive future — [spectral lines]. IEEE Spectrum, 57(1):4, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2020:ISR

[Luc20b] R. W. Lucky. The ingenuity of spies — [reflections]. IEEE Spectrum, 57(1):21, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liu:2024:PTT

[LW24] Mark Liu and H.-S. Philip Wong. The path to a 1-trillion-transistor GPU: AI’s boom demands new chip technology. IEEE Spectrum, 61(7):22–27, July 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:BDP

[Mar20a] A. Marsh. Born digital — [past forward]. IEEE Spectrum, 57(7): 48, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:CCP

[Mar20b] A. Marsh. Calling from the cockpit — [past forward]. IEEE Spectrum, 57(4):56, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:FLO

[Mar20c] A. Marsh. The first laptop in orbit — [past forward]. IEEE Spectrum, 57(6):51, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:FR

[Mar20d] A. Marsh. The first radio. IEEE Spectrum, 57(5):56, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:FU

[Mar20e] A. Marsh. Fun with uranium!

IEEE Spectrum, 57(2):52, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:IXR

[Mar20f] A. Marsh. If the X-ray fits — [past forward]. IEEE Spectrum, 57(11):52, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:PRP

[Mar20g] A. Marsh. The Proto-Roomba — [past forward]. IEEE Spectrum, 57(3):50, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:SDS

[Mar20h] A. Marsh. Special delivery by satellite. IEEE Spectrum, 57 (12):50, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:WLP

[Mar20i] A. Marsh. Weaving light — [past forward]. IEEE Spectrum, 57(8):52, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2020:HFC

[Mar20j] Allison Marsh. The hidden figures of Colossus. IEEE Spectrum, 57(1):64, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

URL https://spectrum.ieee.
org/tech-history/dawn-of-electronics/
the-hidden-figures-behind-
bletchley-parks-codebreaking-
colossus.

Marsh:2021:CRS

[Mar21a] A. Marsh. Catfish, robot, swimmer, spy. IEEE Spectrum, 58(3):51, March 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:CC

[Mar21b] A. Marsh. Counting calories. IEEE Spectrum, 58(1):56, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFa

[Mar21c] A. Marsh. Past forward. IEEE Spectrum, 58(2):48, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:BBT

[Mar21d] Allison Marsh. The black box that’s round and red [past forward]. IEEE Spectrum, 58(7):48, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:CHP

[Mar21e] Allison Marsh. Crystal history [past forward]. IEEE Spectrum, 58(6):68, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:NRP

[Mar21f] Allison Marsh. The Nazi radio [past forward]. IEEE Spectrum, 58(4):56, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFb

[Mar21g] Allison Marsh. Past forward.

IEEE Spectrum, 58(5):56, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFAa

[Mar21h] Allison Marsh. Past forward:

Anthrax by mail. IEEE Spectrum, 58(10):64, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFEb

[Mar21i] Allison Marsh. Past forward: Electricity’s perilous narrative arc. IEEE Spectrum, 58(11): 56, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFW

[Mar21j] Allison Marsh. Past forward: It’s a wonderful light. IEEE Spectrum, 58(12):76, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2021:PFEa

[Mar21k] Allison Marsh. Past forward: The electric motor at 200. IEEE Spectrum, 58(9):72, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFa

[Mar22a] Allison Marsh. Past forward. IEEE Spectrum, 59(5):52, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFb

[Mar22b] Allison Marsh. Past forward. IEEE Spectrum, 59(6):64, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFc

[Mar22c] Allison Marsh. Past forward. IEEE Spectrum, 59(7):48, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFd

[Mar22d] Allison Marsh. Past forward.

IEEE Spectrum, 59(8):48, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFf

[Mar22e] Allison Marsh. Past forward. IEEE Spectrum, 59 (9):68, September 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFg

[Mar22f] Allison Marsh. Past forward.

IEEE Spectrum, 59(10):56, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFh

[Mar22g] Allison Marsh. Past forward. IEEE Spectrum, 59 (11):48, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFi

[Mar22h] Allison Marsh. Past forward.

IEEE Spectrum, 59(12):76, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFE

[Mar22i] Allison Marsh. Past forward:

Electric hearing. IEEE Spectrum, 59(2):48, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFL

[Mar22j] Allison Marsh. Past forward: Lessons from the Little Professor. IEEE Spectrum, 59(4):52, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFO

[Mar22k] Allison Marsh. Past forward: Ode to the Envoy. IEEE Spectrum, 59(1):64, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2022:PFN

[Mar22l] Allison Marsh. Past forward: The Nobelist’s first laser. IEEE Spectrum, 59(3):68, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:CCC

[Mar23a] Allison Marsh. Computer chess, circa 1920. IEEE Spectrum, 60(7):52, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:HOP

[Mar23b] Allison Marsh. History in an object: Past forward. IEEE Spectrum, 60(10):52, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:JWE

[Mar23c] Allison Marsh. James Wimshurst’s electrostatic immortality: Little is known about the man, but his iconic generators live on. IEEE Spectrum, 60(6):68–70, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFa

[Mar23d] Allison Marsh. Past forward.

IEEE Spectrum, 60(1):64, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFb

[Mar23e] Allison Marsh. Past forward.

IEEE Spectrum, 60(4):56, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFc

[Mar23f] Allison Marsh. Past forward.

IEEE Spectrum, 60(5):56, May 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFd

[Mar23g] Allison Marsh. Past forward.

IEEE Spectrum, 60(6):68, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFe

[Mar23h] Allison Marsh. Past forward.

IEEE Spectrum, 60(8):53–57, August 2023. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:PFf

[Mar23i]  Allison Marsh.  Past for-

ward. IEEE Spectrum, 60 (11):52, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2023:TME

[Mar23j] Allison Marsh. Taking the measure of the earthquake that destroyed Tokyo: a century on, we’re good at detecting earthquakes, not so good at protecting against them. IEEE Spectrum, 60 (9):69–72, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mary:2023:CAE

[Mar23k] L. Mary. 5 conclusion from an automation expert with firsthand knowledge of highway regulation what self-driving cars tell us about AI risks: Missy Cummings. IEEE Spectrum, 60 (10):30–35, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:BES

[Mar24a] Allison Marsh. Behold the electronic sackbut. IEEE Spectrum, 61(2):48, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PF

[Mar24b] Allison Marsh. Past forward. IEEE Spectrum, 61(3):64, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFP

[Mar24c] Allison Marsh. Past forward: a planetarium is born. IEEE Spectrum, 61(5):60, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFB

[Mar24d] Allison Marsh. Past forward:

Brass for brains. IEEE Spectrum, 61(6):68, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFFb

[Mar24e] Allison Marsh. Past forward: Farewell, my animatronic friends. IEEE Spectrum, 61 (12):72, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFAb

[Mar24f] Allison Marsh. Past forward:

The automatic rice cooker’s unlikely inventor. IEEE Spectrum, 61(11):76, November 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFFa

[Mar24g] Allison Marsh. Past forward:

The first transistor radio. IEEE Spectrum, 61(10):48, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFM

[Mar24h] Allison Marsh. Past forward: The mechanical TV. IEEE Spectrum, 61(9):68, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFS

[Mar24i] Allison Marsh. Past forward: The Scientology machine. IEEE Spectrum, 61(4):52, April 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFW

[Mar24j] Allison Marsh. Past forward: The wearable computer as bling. IEEE Spectrum, 61(7):48, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:PFAa

[Mar24k] Allison Marsh. Past forword: The AD-X2 affair. IEEE Spectrum, 61(8):48, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marsh:2024:STH

[Mar24l] Allison Marsh. The sound of two hands clapping. IEEE Spectrum, 61(01):64, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McCray:2020:BJH

[McC20] W. P. McCray. Big in Japan: How art, tech, and Pepsico collaborated, then clashed at the 1970 World’s Fair. IEEE Spectrum, 57(3):40–47, March 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandelbaum:2024:IBB

[MCG24] Ryan Mandelbaum, Antonio D.

Córcoles, and Jay Gambetta. IBM’s big bet on the quantumcentric supercomputer: Recent advances point the way to useful classical-quantum hybrids. IEEE Spectrum, 61(9):24–33, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mueller:2023:WHT

[MD23] Matthias Mueller and Roland Dittmeyer. Wind-to-hydrogen tech goes to sea: Is electrolysis cheaper offshore? A new project will find out. IEEE Spectrum, 60 (9):24–29, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2022:N

[MDC^{+}22] Samuel K. Moore, Dave Ditzel, Charles Q. Choi, Mark Harris, Philip E. Ross, and Evan Ackerman. News. IEEE Spectrum, 59(4):5–11, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2023:N

[MHLC23] Samuel K. Moore, Michelle Hampson, Lucas Laursen, and Rina Diane Caballar. News. IEEE Spectrum, 60(6):5–12, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mierse:2022:EMT

[Mie22] Markus Mierse. EZ Mechanical TV: a new twist on the oldest type of television display. IEEE Spectrum, 59(6):16–18, June 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mierse:2024:SFA

[Mie24] Markus Mierse. A sci-fi aerial display: This touchscreen floats in midair. IEEE Spectrum, 61 (11):14–16, November 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mishra:2023:WWW

[Mis23] Umesh K. Mishra. What will win the wide-bandgap wars? IEEE Spectrum, 60(4):32–39, April 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mitry:2020:RSK

[Mit20] M. Mitry. Routers in space: Kepler communications’ CubeSats will create an Internet for other satellites. IEEE Spectrum, 57 (2):38–43, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2021:LDT

[MJH^{+}21] Samuel K. Moore, Dexter Johnson, Mark Harris, Emily Waltz, Prachi Patel, and Michelle Hampson. The latest developments in technology, engineering, and science: News. IEEE Spectrum, 58(8):5–12, August 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mckeever:2020:MSM

[MJS20] J. Mckeever, D. Jervis, and M. Strupler. Microsatellites spot mystery methane leaks. IEEE Spectrum, 57(11):38–43, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mendelsohn:2024:N

[MLF^{+}24] Kohava Mendelsohn, Lucas Laursen, Peter Fairley, Michelle Hampson, Sally Adee, and Greg Uyeno. News. IEEE Spectrum, 61(12):7–14, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Majumdar:2024:AAI

[MMM24] Somdeb Majumdar, Uday Mallappa, and Hesham Mostafa. AI alone isn’t ready for chip design: a combination of classical search and machine learning may be the way forward. IEEE Spectrum, 61 (12):38–43, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maclean:2022:EVE

[MMP22] Heather L. Maclean, Alexandre Milovanoff, and I. Daniel Posen. The electric vehicle is not enough: To decarbonize road transport we need to complement EVs with bikes, rail, city planning, and alternative energy. IEEE Spectrum, 59(11): 28–33, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2020:WHM

[Moo20a] S. K. Moore. 4 ways to handle more qubits — [news]. IEEE Spectrum, 57(2):9–10, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2020:CFP

[Moo20b] S. K. Moore. Chiplets are the future of processors: Three advances boost performance, cut costs, and save power. IEEE Spectrum, 57(5):11–12, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2020:HSR

[Moo20c] S. K. Moore. How systemic racism destroyed black innovation in the U.S.: Violence and segregation undermined African American inventors throughout the 20th century — [news]. IEEE Spectrum, 57(9):6–7, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2020:NN

[Moo20d] S. K. Moore. The node is nonsense. IEEE Spectrum, 57 (8):24–30, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2020:HCS

[Moo20e] Samuel K. Moore. Huge chip smashes deep learning’s speed barrier. IEEE Spectrum, 57 (1):24–27, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2021:UDG

[Moo21] S. K. Moore. The ups and downs of gravity energy storage: Startups are pioneering a radical new alternative to batteries for grid storage. IEEE Spectrum, 58(1):38–39, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2022:PP

[Moo22a] Samuel K. Moore. 3 paths to 3D processors. IEEE Spectrum, 59(6):24–29, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2022:ACC

[Moo22b] Samuel K. Moore. AI computing comes to memory chips: Samsung will double performance of neural nets with processing-in-memory. IEEE Spectrum, 59 (1):40–41, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2022:TWW

[Moo22c] Samuel K. Moore. The transistor of 2047: What will the device be like on its 100th anniversary? IEEE Spectrum, 59 (12):38–39, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2023:QSN

[Moo23a] Samuel K. Moore. 5 questions for Sam Naffziger: How he drove AMD’s pioneering move to chiplets. IEEE Spectrum, 60(7): 21, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2023:DBH

[Moo23b] Samuel K. Moore. A dark (blue) horse emerges to speed up computing: Avicena’s blue microLEDs are in a race with Ayar Lab’s laser-based system. IEEE Spectrum, 60(1):30–31, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2024:QRT

[Moo24a] Samuel K. Moore. 5 qestions: Rao Tummala: How to start a chip industry from scratch. IEEE Spectrum, 61(8):23, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2024:QJR

[Moo24b] Samuel K. Moore. 5 questions for Julien Ryckaert: Why CMOS 2.0 is the next phase of Moore’s Law. IEEE Spectrum, 61(3):21, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2024:CCE

[Moo24c] Samuel K. Moore. Chips to compute with encrypted data are coming: Fully homomorphic encryption could make data unhackable. IEEE Spectrum, 61 (01):38–40, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2024:CCH

[Moo24d] Samuel K. Moore. The copper connection: Hybrid bonding is the 3D-chip tech that’s saving Moore’s Law. IEEE Spectrum, 61 (9):34–39, September 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2024:SWU

[Moo24e] Samuel K. Moore. Semiconductors: Will the U.S. CHIPS Act speed the lab-to-fab transition?: a semiconductor startup’s journey shows how it could work. IEEE Spectrum, 61(11): 52–57, November 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moser:2022:RHB

[Mos22] Tobias Moser. Restoring hearing with beams of light. IEEE Spectrum, 59(9):30–35, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2022:STY

[MS22] Samuel K. Moore and David Schneider. The state of the transistor: In 75 years, it’s become tiny, mighty, ubiquitous, and just plain weird. IEEE Spectrum, 59 (12):30–31, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2021:HDL

[MSS21] Samuel K. Moore, David Schneider, and Eliza Strickland. How deep learning works: Inside the neural networks that power today’s AI. IEEE Spectrum, 58 (10):32–33, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullaney:2024:BQO

[Mul24] Thomas S. Mullaney. Before QWERTY, the options were ingenious and strange: The forgotten history of Chinese keyboards. IEEE Spectrum, 61 (6):30–48, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murray:2023:LIB

[Mur23] Charles J. Murray. The lithiumion battery’s long and winding road. IEEE Spectrum, 60 (8):40–45, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mercier:2020:PST

[MW20] P. Mercier and J. Wang. Powered by sweat: Throw out the batteries: Biofuels will change the future of wearable devices. IEEE Spectrum, 57(7):28–33, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mantooth:2021:VCS

[MZR21] Alan Mantooth, Carl-Mikael Zetterling, and Ana Rusu. Venus calling silicon carbide radio: Circuits can take the heat needed to phone home from our hellish sister planet. IEEE Spectrum, 58(5):24–30, May 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Namuduri:2023:CSB

[NG23] Kamesh Namuduri and Keven Gambold. 3 challenges to solve before air taxis can soar: a Texas consortium is working out a traffic system for eV-TOLs. IEEE Spectrum, 60 (8):58–63, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nordrum:2020:SDG

[Nor20a] A. Nordrum. Storing data in glass: Optical storage media is advancing from CDs and DVDs to small glass squares — [news]. IEEE Spectrum, 57 (6):10–11, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nordrum:2020:TSH

[Nor20b] A. Nordrum. Transgenic salmon hits U.S. shelves. IEEE Spectrum, 57(1):44–47, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nordrum:2020:EDP

[NS20] A. Nordrum and E. Strickland.

Engineering during a pandemic:

7 CEOs, engineers, and scientists describe how their work has changed in response to COVID-19 — [news]. IEEE Spectrum, 57(5):6–7, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Noda:2024:BSL

[NYI24] Susumu Noda, Masahiro Yoshida, and Takuya Inoue. The brightest semiconductor laser ever: Photonic crystals enable tiny lasers to melt steel. IEEE Spectrum, 61(5):22–57, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ong:2020:UCP

[Ong20]  S. Ong.  U.S. cities pilot

connected-vehicle tech: Enabling conventional vehicles to communicate could save lives right now — [news]. IEEE Spectrum, 57(7):07–09, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Palomaki:2022:QDO

[Pal22] Peter Palomaki. Quantum dots + OLED = your next TV: Formerly rival technologies will come together in new Samsung displays. IEEE Spectrum, 59 (1):52–53, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Panwar:2020:BMB

[Pan20] S. Panwar. Breaking the millisecond barrier: Robots and selfdriving cars will need completely reengineered networks. IEEE Spectrum, 57(11):44–49, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2020:PBS

[Pat20] P. Patel. Potassium batteries show promise: Hurdles remain, but potassium could someday make sense for grid storage — [news]. IEEE Spectrum, 57(3):11–12, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2022:CGW

[Pat22] Prachi Patel. China’s Green Winter Olympics: a variety of climate-friendly strategies will be on show, along with the athletes. IEEE Spectrum, 59(1): 44–45, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2023:BUS

[Pat23a]  Prachi Patel.  Building a

U.S. semiconductor workforce: CHIPS Act-funded new fabs are spawning university programs. IEEE Spectrum, 60(6):28–35, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2023:CNB

[Pat23b]  Prachi Patel.  China’s new

breeder reactors may produce more than just watts: They could also make weapons-grade plutonium. IEEE Spectrum, 60 (1):38–39, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2024:KFM

[Pat24a] Prachi Patel. Kelp farms make carbon capture a growth industry: Solar-powered systems are already sequestering carbon in the briny deep. IEEE Spectrum, 61(01):40–41, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2024:MSN

[Pat24b] Prachi Patel. Merchant shipping’s nuclear option: Nuclearpowered cargo ships failed decades ago, but small modular reactors could make all the difference. IEEE Spectrum, 61 (2):36–41, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paul:2021:TEE

[Pau21] Jon D. Paul. In the twilight of electromechanical encryption, an exceptional machine figured in a major spy scandal. IEEE Spectrum, 58(9):32–52, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paulson:2022:NGD

[Pau22] Matthew Paulson. A next-gen DIY audio amp: a Web-enabled integrated amp that won’t break the bank. IEEE Spectrum, 59 (3):16–18, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

PATEL:2024:N

[PCRC24] PRACHI PATEL, RINA DIANE CABALLAR, GWENDOLYN RAK, and CHARLES Q. CHOI. News. IEEE Spectrum, 61 (01):7–15, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:BST

[Per20a] T. S. Perry. Biosensors for on-the-spot tests: Two companies plan to debut new tech for COVID-19 testing by year’s end — [news]. IEEE Spectrum, 57(7):09–10, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:BTP

[Per20b] T. S. Perry. Black tech professionals are still paid less than their white colleagues: And women make less than their male colleagues, regardless of racial identity — [spectral lines]. IEEE Spectrum, 57(8):4, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:FFC

[Per20c] T. S. Perry. The father of Fin-Fets: Chenming Hu took transistors into the third dimension to save Moore’s Law. IEEE Spectrum, 57(5):46–51, May 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:PPT

[Per20d] T. S. Perry. The post-pandemic tech workplace: Hopes and expectations for the new normal — [departments]. IEEE Spectrum, 57(9):19, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:RUA

[Per20e] T. S. Perry. Researchers are using algorithms to tackle the coronavirus test shortage: The scramble to develop new test kits that deliver faster results — [spectral lines]. IEEE Spectrum, 57(6):4, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:RRU

[Per20f] T. S. Perry. Rocket recruiting is using AI to get you a job. For free: Its candidate pool is attracting recruiters from small and large companies. IEEE Spectrum, 57(9):4, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:SHT

[Per20g] T. S. Perry. Stay-at-home tech’s star turn: Even technophobes rely on consumer tech now. IEEE Spectrum, 57(9):11–12, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:TPC

[Per20h] T. S. Perry. Take a pay cut and work wherever you want: But should the value of your work depend on where it gets done? — [spectral lines]. IEEE Spectrum, 57(11):4, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:WDI

[Per20i] T. S. Perry. Wage discrimination increases in tech: Women lose gains made in 2018 — [careers]. IEEE Spectrum, 57(6):17, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:AVY

[Per20j] Tekla S. Perry. AR/VR is this year’s hot ticket for jobs: But growth in demand for blockchain developers stutters — [careers]. IEEE Spectrum, 57(4):19, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:GAY

[Per20k] Tekla S. Perry. A guardian angel for your car. IEEE Spectrum, 57(1):54–55, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:HCD

[Per20l] Tekla S. Perry. Here comes driverless ride sharing: Cruise unveils the origin, a fully autonomous SUV designed for appcontrolled urban transportation — [spectral lines]. IEEE Spectrum, 57(3):4, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:ITD

[Per20m] Tekla S. Perry. Is it time to drag tech jobs out of Silicon Valley?

— [careers]. IEEE Spectrum, 57(2):50, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:JDQ

[Per20n] Tekla S. Perry. John Deere’s quest to solve agricultures deeplearning problems — [spectral lines]. IEEE Spectrum, 57 (2):4, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2020:WYC

[Per20o] Tekla S. Perry. When you cut, copy, or paste today, thank Larry Tesler: Remembering the computer scientist who revolutionized the user interface — [spectral lines]. IEEE Spectrum, 57(4): 4, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2021:LAA

[Per21a] T. S. Perry. Look out for Apple’s AR glasses: With head-up displays, cameras, inertial sensors, and lidar on board, Apple’s augmented-reality glasses could redefine wearables. IEEE Spectrum, 58(1):26–54, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2021:CCW

[Per21b] Tekla S. Perry. Conjurer of compression: From WinZips to cat GIFs, Jacob Ziv’s algorithms have been making data disappear and reappear for decades. IEEE Spectrum, 58(5):48–53, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2021:TPR

[Per21c] Tekla S. Perry. Tech pay rises (almost) everywhere: The great resignation is pushing salaries up. IEEE Spectrum, 58(12):17, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2022:AML

[Per22a] Tekla S. Perry. Asad Madni and the lifesaving sensor: His pivot away from defense led to a tiny tuning fork that helped prevent SUV rollovers and plane crashes. IEEE Spectrum, 59 (5):24–29, May 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2022:BWM

[Per22b] Tekla S. Perry. A bitcoin wallet for the masses: Square simplified credit-card transactions. now it wants to build cryptocurrency hardware. IEEE Spectrum, 59 (1):42–43, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2022:IAI

[Per22c] Tekla S. Perry. Inflation-adjusted income for U.S. engineers drops: Insights from IEEE–USA’s annual salary survey. IEEE Spectrum, 59(12):16–17, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2023:MIV

[Per23] Tekla S. Perry. Mr. Internet: Vint Cerf’s 1973 sketch kicked off five decades of improving and evangelizing what we now know as the Internet. IEEE Spectrum, 60(5):24–31, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2024:CIY

[Per24a] Jeff Perry. Careers: Improve your workplace communication: Pro tips for talking to leaders, clients, and teams. IEEE Spectrum, 61(12):21–22, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2024:EFF

[Per24b] Tekla S. Perry. Engineering the first Fitbit: The fitness tracker’s creators had a billion-dollar idea and a lot to learn about hardware. IEEE Spectrum, 61(9): 40–47, September 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:BTV

[Pes20a] M. Pesce. Being there, virtually — [macro micro]. IEEE Spectrum, 57(8):21, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:CSI

[Pes20b] M. Pesce. Certifiably secure IoT — [macro micro]. IEEE Spectrum, 57(10):19, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:CTM

[Pes20c] M. Pesce. The complacency trap — [macro micro]. IEEE Spectrum, 57(3):23, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:BUC

[Pes20d] M. Pesce. Not business as usual — [CrossTalk]. IEEE Spectrum, 57(6):21, June 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:PTC

[Pes20e] M. Pesce. Privacy in the time of COVID-19. IEEE Spectrum, 57(5):23, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2020:SSC

[Pes20f] M. Pesce. Spidey senses — [CrossTalk]. IEEE Spectrum, 57(7):21, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2021:PPE

[Pes21a] M. Pesce. Peering into the pandemic end game: Before COVID 19 fades, we’ll see a flurry of advances in contact tracing, cloud computing, surveillance, and online gaming. IEEE Spectrum, 58 (1):22–25, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2021:BAM

[Pes21b] Mark Pesce. Bricked by age: Manufacturers should supply the software to keep things working indefinitely. IEEE Spectrum, 58(10):22, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pesce:2022:MM

[Pes22] Mark Pesce. Macro & micro. IEEE Spectrum, 59(1):24, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Palomaki:2020:SQD

[PK20] P. Palomaki and S. Keuleyan. Snapshots by quantum dots: Look out, CMOS image sensors. There’s a new kid in cameratown. IEEE Spectrum, 57 (3):24–29, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pepitone:2023:N

[PLUM23] Julianne Pepitone, Lucas Laursen, Lawrence Ulrich, and Samuel K. Moore. News. IEEE Spectrum, 60(12):5–12, December 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2023:N

[PMPG23] Prachi Patel, Samuel K. Moore, Ned Potter, and Edd Gent. News. IEEE Spectrum, 60 (4):5–13, April 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parisa:2020:ACS

[PN20] Z. Parisa and M. Nova. This AI can see the forest and the trees. IEEE Spectrum, 57(8): 32–37, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pomazov:2021:H

[Pom21] E. Pomazov. Hands on. IEEE Spectrum, 58(2):14–16, February 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poor:2024:GNB

[Poo24a] Alfred Poor. Grid “Neurons”

boost power-line performance. IEEE Spectrum, 61(8):16–17, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poor:2024:HMT

[Poo24b] Alfred Poor. How to make TV screens disappear: See-through TVs could banish the big black rectangle — at a cost. IEEE Spectrum, 61(10):34–39, October 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Potter:2020:MHP

[Pot20a] N. Potter. A Mars helicopter preps for launch: The first drone to fly on another planet will hitch a ride on NASA’s Perseverance rover — [news]. IEEE Spectrum, 57(7):06–07, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Potter:2020:MBP

[Pot20b] N. Potter. May the best personal flyer win: The $1 million GoFly Prize has yet to be claimed — [news]. IEEE Spectrum, 57(4):6–7, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Potter:2024:ANM

[Pot24] Ned Potter. Aerospace: NASA made the Hubble Telescope to be remade: Spacewalk repairs and upgrades were always part of the plan. IEEE Spectrum, 61 (11):50–51, November 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poulakis:2022:MST

[Pou22] Marios Poulakis. 6G’s metamaterials solution: There’s plenty of bandwidth available if we use reconfigurable intelligent surfaces. IEEE Spectrum, 59(11): 40–45, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Powell:2024:HRP

[Pow24] Corey S. Powell. The hunt for rogue planets: The galaxy may hold a trillion wandering, sunless worlds. so how do we find them? IEEE Spectrum, 61 (8):24–35, August 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2022:N

[PPUS22] Prachi Patel, Ned Potter, Lawrence Ulrich, and Matthew S. Smith. News. IEEE Spectrum, 59(7):5–11, July 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pretz:2020:UMU

[Pre20] K. Pretz. Using modeling to understand how COVID-19 preventive measures work: Washing hands, wearing masks: Does it help? — [spectral lines]. IEEE Spectrum, 57(10):4, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Price:2021:BLS

[Pri21] D. Price. Breakthrough Listen is scanning more of space than ever to try to find intelligent life. IEEE Spectrum, 58 (2):34–39, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pelkey:2023:EID

[PRR23] James L. Pelkey, Loring G.

Robbins, and Andrew L. Russell. The early Internet’s do-or-die moments: a handful of demos shaped today’s networked world. IEEE Spectrum, 60 (5):46–52, May 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Psaras:2022:IFS

[PSD22] Yiannis Psaras, Jorge M. Soares, and David Dias. To the Inter-Planetary File System and beyond!: Peer-to-peer file sharing would make the Internet far more efficient. IEEE Spectrum, 59 (11):34–39, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rabault:2022:TAI

[Rab22] Jean Rabault. Tracking Arctic ice: an inexpensive sensor package gets data back despite harsh conditions. IEEE Spectrum, 59(4):14–16, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Radu:2020:ATT

[Rad20] I. Radu. Atom-thick transistors. IEEE Spectrum, 57(2): 44–49, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rae:2020:BFB

[Rae20] A. Rae. This blimp flies on buoyancy alone: the Phoenix turns up-and-down motion into thrust. IEEE Spectrum, 57 (7):34–39, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rafter:2020:TIR

[Raf20a] M. V. Rafter. Traffic Institute to review yellow-light timing — [news]. IEEE Spectrum, 57(1):8–10, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rafter:2020:WFG

[Raf20b] M. V. Rafter. Wanted: a fallback for GPS: The U.S. government seeks tech that could step in if its global satellite fleet were to fail — [news]. IEEE Spectrum, 57(5):9–10, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rak:2024:IHL

[Rak24a] Gwendolyn Rak. Intel hopes to leapfrog its competitors: The chipmaker bets on new transistors, power-delivery tech, and a nascent foundry business. IEEE Spectrum, 61(01):32–56, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rak:2024:SYP

[Rak24b] Gwendolyn Rak. The shrink in your pocket. IEEE Spectrum, 61(7):42–47, July 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rao:2024:QSB

[Rao24] Rahul Rao. 5 questions: Scott Best: Prepping today’s systems for tomorrow’s post-quantum cryptography. IEEE Spectrum, 61(7):21, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rashid:2020:FED

[Ras20a] F. Y. Rashid. The fight over encrypted DNS — [news]. IEEE Spectrum, 57(1):11–12, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rashid:2020:RCC

[Ras20b] F. Y. Rashid. The rise of confidential computing: Big tech companies are adopting a new security model to protect data while it’s in use — [news]. IEEE Spectrum, 57(6):8–9, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rao:2023:N

[RCAP23] Rahul Rao, Charles Q Choi, Evan Ackerman, and Prachi Patel. News. IEEE Spectrum, 60(3):5–12, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Redd:2020:BSB

[Red20] N. T. Redd. Bringing satellites back from the dead: Mission extension vehicles give defunct spacecraft a new lease on life — [news]. IEEE Spectrum, 57(8):6–7, August 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Regan:2023:WPM

[Reg23] Matthew Regan. When Pac-Man met Turing: The game runs on an emulated 6502 CPU. IEEE Spectrum, 60(6):16–18, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Russ:2020:WHP

[RG20] S. H. Russ and J. Gatlin. Ways to hack a printed circuit board: PCB production is an underappreciated vulnerability in the global supply chain. IEEE Spectrum, 57(9):38–43, September 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rezvani:2022:LRS

[RHBM22] Behrooz Rezvani, Babak Hassibi, Fredrik Brännström, and Majid Manteghi. Letting robocars see around corners: Using several bands of radar at once can give cars a kind of second sight. IEEE Spectrum, 59 (2):36–41, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Radosavljevic:2022:TML

[RK22] Marko Radosavljevic and Jack Kavalieros. Taking Moore’s Law to new heights: When transistors can’t get any smaller, the only direction is up. IEEE Spectrum, 59 (12):32–37, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rallabandi:2024:HEC

[ROK24] Vandana Rallabandi, Burak Ozpineci, and Praveen Kumar. How EVs can escape the rare earth trap: Promising experimental motors are using exotic materials and ingenious configurations. IEEE Spectrum, 61 (8):36–41, August 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2020:FBM

[Ros20a] P. E. Ross. Flying beyond Mach 5 is back, decades after the original need-for-speed arms race ended: Going hypersonic. IEEE Spectrum, 57(12):32–37, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2020:PSM

[Ros20b] P. E. Ross. Putting sirens on mute: Noise-canceling smart windows filter out street sounds — [news]. IEEE Spectrum, 57 (9):8–9, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2021:STR

[Ros21] P. E. Ross. Supersonic travel returns: Boom’s XB-1 test aircraft may usher in faster-than-sound commercial flight. IEEE Spectrum, 58(1):40–41, January 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenberg:2022:HPA

[Ros22a]  Louis Rosenberg.  How a

parachute accident helped jumpstart augmented reality: In 1992, hardware for the first interactive AR system literally fell from the skies. IEEE Spectrum, 59 (6):42–47, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosezky:2022:DGR

[Ros22b] Matthias Rosezky. DIY gammaray spectroscopy: But beware a glitch in the popular Pi Pico microcontroller. IEEE Spectrum, 59(7):14–16, July 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2022:BST

[Ros22c] Philip E. Ross. Brakes that slam themselves: Automatic emergency braking will become standard in Europe. IEEE Spectrum, 59(1):32–33, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2022:FPP

[Ros22d] Philip E. Ross. Flying pallets without pilots: a drone startup will test a radical new vision of long-range cargo transport in Europe. IEEE Spectrum, 59 (1):30–31, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2022:STR

[Ros22e] Philip E. Ross. Schrödinger’s tardigrade: Have researchers quantum-entangled hardy critters? IEEE Spectrum, 59 (2):19, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2023:QMC

[Ros23a] Philip E. Ross. 5 questions for Missy Cummings: The former fighter pilot on why autonomous vehicles are so risky. IEEE Spectrum, 60(6):21, June 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2023:EDR

[Ros23b] Philip E. Ross. Economics drives a ray-gun resurgence: Lasers, cheaper by the shot, should work well against drones and cruise missiles. IEEE Spectrum, 60 (1):40–41, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2023:N

[RPC^{+}23] Philip E. Ross, Prachi Patel, Charles Q. Choi, Michelle Hampson, Edd Gent, Margo Anderson, and Tekla S. Perry. News. IEEE Spectrum, 60(7):6–13, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raghavan:2023:BNP

[RS23] Barath Raghavan and Bruce Schneier. A bold new plan for preserving online privacy and security: Decoupling our identities from our data and actions could safeguard our secrets in the cloud. IEEE Spectrum, 60 (12):22–29, December 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rao:2021:AMI

[RZ21] S. Rao and L. Zhang. The algorithms that make Instacart roll: How machine learning and other tech tools guide your groceries from store to doorstep. IEEE Spectrum, 58(3):36–42, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sachan:2020:PBV

[Sac20] D. Sachan. Profile: Blessin Varkey: Creating tech for students with disabilities — [careers]. IEEE Spectrum, 57(7):17, July 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sarnoff:2020:BCA

[Sar20] M. Sarnoff. 8 bits, 6 chips: The Amethyst home computer is surprisingly capable — [hands on]. IEEE Spectrum, 57(4):16–18, April 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:LDT

[SAS^{+}21] David Schneider, Evan Ackerman, Eliza Strickland, Lawrence Ulrich, and Rahul Rao. The latest developments in technology, engineering, and science: News. IEEE Spectrum, 58(9): 7–15, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:CEM

[SB22a] Vaclav Smil and Rodney Brooks.

COVID: Excess mortalities two years later: The death toll is increasingly comparable to that of the 1918–1920 flu. IEEE Spectrum, 59(5):20–22, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2022:GMMa

[SB22b] Matthew S. Smith and Rodney Brooks. Gizmo: Myth and machine. IEEE Spectrum, 59 (1):22–23, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2022:GMMb

[SB22c] Matthew S. Smith and Rodney Brooks. Gizmo: Myth and machine. IEEE Spectrum, 59 (2):22–23, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schloemer:2023:SPC

[SC23] Tracy H. Schloemer and Daniel N. Congreve. Smashing photons: Changing the color of light supercharges solar energy, 3D printing, and night vision. IEEE Spectrum, 60(10):44–49, October 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scalfani:2022:NWS

[Sca22] Charles Scalfani. A new way to squash bugs: Functional programming is hard to learn but yields fewer nasty surprises. IEEE Spectrum, 59(12): 40–45, December 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2020:DDC

[Sch20a] D. Schneider. The delivery drones are coming. IEEE Spectrum, 57(1):28–29, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2020:DAI

[Sch20b] D. Schneider. A DIY audio induction loop: Here’s a simple way to transmit audio to people who wear T-coil-equipped hearing aids — [hands on]. IEEE Spectrum, 57(2):14–16, February 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2020:PDB

[Sch20c] D. Schneider. Pedaling out of the dark: Build a double-backup home power supply — [hands on]. IEEE Spectrum, 57(11): 16–18, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2020:ZBL

[Sch20d] D. Schneider. The Zoom Box:

Look people in the eye while videoconferencing — [hands on]. IEEE Spectrum, 57(8):14–16, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:DLS

[Sch21a] D. Schneider. Deep learning at the speed of light: Lightmatter bets that optical computing can solve AI’s efficiency problem. IEEE Spectrum, 58 (1):28–29, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:WNO

[Sch21b] D. Schneider. Where no one has seen before: The James Webb Space Telescope will let us see back almost to the Big Bang. IEEE Spectrum, 58(1): 30–31, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:H

[Sch21c] David Schneider. Hands on. IEEE Spectrum, 58(4):14–16, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:IBE

[Sch21d] David Schneider. An instant e-bike: Electrifying a bike can be electrifyingly easy. IEEE Spectrum, 58(9):18–20, September 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2021:MTG

[Sch21e] David Schneider. Mach 2, take 2: Grappling again with the promise and peril of supersonic airliners. IEEE Spectrum, 58 (11):26–31, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2022:BSW

[Sch22a] David Schneider. A barometer for space weather: Detect solar flares and gamma-ray bursts for less than $100. IEEE Spectrum, 59(2):16–18, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2022:CDW

[Sch22b]  David Schneider.  Chasing

downed weather balloons: Hunt them with a cheap SDR receiver and a homemade antenna. IEEE Spectrum, 59(9):16–18, September 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2022:DSM

[Sch22c] David Schneider. Deep-sea mining stirs up muddy questions: A controversial pilot program will collect metal-rich nodules from the ocean floor. IEEE Spectrum, 59(1):56–57, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2022:DMS

[Sch22d] David Schneider. DIY magnetometer: Search for buried treasure for less than $100. IEEE Spectrum, 59(5):16–18, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2022:EEU

[Sch22e] David Schneider. The Exascale Era is upon us: The Frontier supercomputer may be the first to reach 1,000,000,000,000,000,000 operations per second. IEEE Spectrum, 59(1):34–35, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2023:DQR

[Sch23a] David Schneider. Detect quakes with raspberry shakes: a raspberry pi-based seismometer can detect even distant earthquakes. IEEE Spectrum, 60 (10):16–18, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2023:RFE

[Sch23b]  David Schneider.  A retro

fuel-economy computer: This Arduino-based project outputs MPGs on a vintage dashboard gauge. IEEE Spectrum, 60 (3):16–18, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2023:RYO

[Sch23c] David Schneider. Roll your own all-sky camera: Use Raspberry Pi hardware to capture mesmerizing time-lapse images of the heavens. IEEE Spectrum, 60(7):16–18, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2024:BND

[Sch24a] David Schneider. Birding at night: Detect migrating birds with a plastic dish and a cheap microphone. IEEE Spectrum, 61(7):16–18, July 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2024:EAA

[Sch24b] David Schneider. Electronically assisted astronomy: Take detailed images of the heavens on the cheap. IEEE Spectrum, 61(5):16–18, May 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schroeder:2024:WIW

[Sch24c] Karl Schroeder. What if we built a really big computer? IEEE Spectrum, 61(3):34–43, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:N

[SCMP22] Eliza Strickland, Charles Q.

Choi, Samuel K. Moore, and Prachi Patel. News. IEEE Spectrum, 59(8):5–12, August 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2023:N

[SCPC23] Matthew S. Smith, Rina Diane Caballar, Prachi Patel, and Charles Q. Choi. News. IEEE Spectrum, 60(1):8–15, January 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2022:N

[SCPG22] Craig S. Smith, Charles Q. Choi, Prachi Patel, and Edd Gent. News. IEEE Spectrum, 59(11):

5–12, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scudellari:2020:QDT

[Scu20a] M. Scudellari. Quantum-dot tattoos store vaccine history — [news]. IEEE Spectrum, 57 (2):11, February 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scudellari:2020:RBA

[Scu20b] M. Scudellari. Researchers are betting that AI and automation can cut drug discovery from five years to six months: Automating antivirals. IEEE Spectrum, 57(10):44–49, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:FWN

[SG22] Eliza Strickland and Maria Gallucci. First win for the neurorights campaign: Chile plans to regulate all neurotech and ban the sale of brain data. IEEE Spectrum, 59(1):26–58, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Samad:2024:CDY

[SG24] Tariq Samad and Gus Gaynor. Careers: Do you actually want to be a manager?: Understand the pros and cons before you decide to switch tracks. IEEE Spectrum, 61(10):19–20, October 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2023:N

[SGS^{+}23] Eliza Strickland, Edd Gent, Zoltán Spakovszky, Mark Harris, and Michael Koziol. News. IEEE Spectrum, 60(9):1–8, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:DD

[SH22a] Eliza Strickland and Mark Harris. Dynamic duo. IEEE Spectrum, 59(3):2, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:WHW

[SH22b] Eliza Strickland and Mark Harris. What happens when a bionic body part becomes obsolete?: Blind people with second sight’s retinal implants found out. IEEE Spectrum, 59(3):24–31, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:2023:PMI

[Sha23] Scott J. Shapiro. Patch me if you can: How a New Jersey teenager’s malware threatened the entire Internet. IEEE Spectrum, 60(6):42–48, June 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2024:Na

[SHC^{+}24] Eliza Strickland, Michelle Hampson, Charles Q. Choi, Margo Anderson, Matthew S. Smith, and Evan Ackerman. News. IEEE Spectrum, 61(9):5–12, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stewart:2020:HCT

[SHD20] G. Stewart, K. Heusden, and G. A. Dumont. How control theory can help us control Covid-19. IEEE Spectrum, 57(6):22–29, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shestakofsky:2024:MRB

[She24] Benjamin Shestakofsky. The messy reality behind a Silicon Valley unicorn. IEEE Spectrum, 61(4):44–49, April 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shirriff:2022:VTF

[Shi22] Ken Shirriff. The vacuum tube’s forgotten rival: Magnetic amplifiers, the alt-tech of the Third Reich, lasted into the Internet era. IEEE Spectrum, 59 (4):44–49, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sackett:2024:DWD

[SHP24a] Casey Sackett, Devin Harper, and Aaron Pavez. Do we dare use generative AI for mental health? IEEE Spectrum, 61 (6):42–47, June 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2024:Nb

[SHP^{+}24b] Eliza Strickland, Tim Hornyak, Ned Potter, Kohava Mendelsohn, and Dina Genkina. News. IEEE Spectrum, 61(11):5–12, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Slade:2021:DAT

[Sla21] Patrick Slade. DIY activity tracker: Count your calories more accurately than a smartphone. IEEE Spectrum, 58 (10):14–16, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2024:IWA

[SM24] Eliza Strickland and Peter B.

Meyer. Introduction: Why the art of invention is always being reinvented: As IEEE Spectrum turns 60, we celebrate creativity and chutzpah in tech. IEEE Spectrum, 61(11):18–22, November 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sen:2020:BNS

[SMD20] S. Sen, S. Maity, and D. Das. The body is the network: To safeguard sensitive data, turn flesh and tissue into a secure wireless channel. IEEE Spectrum, 57 (12):44–49, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:BGP

[Smi20a] V. Smil. Building the Great Pyramid — CrossTalk. IEEE Spectrum, 57(6):18–19, June 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:CNC

[Smi20b] V. Smil. Claiming the Nile — [CrossTalk]. IEEE Spectrum, 57 (11):20–21, November 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:CMB

[Smi20c] V. Smil. Cranes (the machines, not birds) — [CrossTalk]. IEEE Spectrum, 57(8):18–19, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:EYL

[Smi20d] V. Smil. Energiewende, 20 years later — [CrossTalk]. IEEE Spectrum, 57(12):22–23, December 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:DWC

[Smi20e] V. Smil. (Not) drinking wine — [CrossTalk]. IEEE Spectrum, 57(4):20–21, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:POR

[Smi20f] V. Smil. Peak oil: a retrospective — [CrossTalk]. IEEE Spectrum, 57(5):20–21, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:WCR

[Smi20g] V. Smil. Why chicken rules — [CrossTalk]. IEEE Spectrum, 57 (1):18–19, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2020:YCA

[Smi20h] V. Smil. The year is 118 A.A.C. (after air conditioning)

— [CrossTalk]. IEEE Spectrum, 57(7):18–19, July 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:C

[Smi21a] V. Smil. CrossTalk. IEEE Spectrum, 58(3):18–19, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:NDLa

[Smi21b] V. Smil. Numbers don’t LIE. IEEE Spectrum, 58(1): 16–17, January 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:NDLb

[Smi21c]  V. Smil.  Numbers don’t

lie. IEEE Spectrum, 58(2): 18–19, February 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:BCI

[Smi21d] Vaclav Smil. A boat can indeed be too long and too skinny: The length-to-beam ratio still has practical limits: Numbers don’t lie. IEEE Spectrum, 58 (8):18–19, August 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:EFB

[Smi21e] Vaclav Smil. Electric flight: Batteries are nowhere near able to sustain wide-body airliners over flights measuring in the thousands of kilometers: Numbers don’t lie. IEEE Spectrum, 58 (11):22–23, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:HHO

[Smi21f] Vaclav Smil. The hyperloop is hyper old: Elon Musk merely renamed a 200-year-old dream: Numbers don’t lie. IEEE Spectrum, 58(12):18–19, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:NDLc

[Smi21g] Vaclav Smil. Numbers don’t lie. IEEE Spectrum, 58(4):18–19, April 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:NDLd

[Smi21h] Vaclav Smil. Numbers don’t lie. IEEE Spectrum, 58(6):20–21, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:NDLe

[Smi21i] Vaclav Smil. Numbers don’t lie. IEEE Spectrum, 58(7):20–21, July 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:SAG

[Smi21j] Vaclav Smil. SUVs ascendant:

Growth in SUV use could more than offset carbon savings from electric vehicles: Numbers don’t lie. IEEE Spectrum, 58(9): 22–23, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2021:WGW

[Smi21k] Vaclav Smil. What goes up:

World population growth has slowed, and some countries are actually declining: Numbers don’t lie. IEEE Spectrum, 58 (10):18–19, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:RSP

[Smi21l] Alvy Ray Smith. The real story of Pixar: How a bad hardware company turned itself into a great movie studio. IEEE Spectrum, 58(11):44–49, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:WLS

[Smi21m] Alvy Ray Smith. What a long, strange trip it’s been. IEEE Spectrum, 58(11):2, November 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:IAQ

[Smi21n] Matthew S. Smith. The indoor air-quality paradox: Easy to measure tough to fix. IEEE Spectrum, 58(10):20, October 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:AAE

[Smi22a] Vaclav Smil. Africa’s access to electricity: In an era of plenty, one continent still lags: Numbers don’t lie. IEEE Spectrum, 59(3):20–21, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:CFP

[Smi22b] Vaclav Smil. The carbon footprint of plate glass: The energy required is not transparent: Numbers don’t lie. IEEE Spectrum, 59(4):18–19, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:DAC

[Smi22c] Vaclav Smil. Decarbonization algebra: The COP26 calls for impossibly steep cuts in carbon emissions: Numbers don’t lie. IEEE Spectrum, 59(2): 20–21, February 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:DOC

[Smi22d] Vaclav Smil. Decarbonization is our costliest challenge: It has no clear beginning or end, and it affects every aspect of life. IEEE Spectrum, 59(10):22–23, October 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:EDP

[Smi22e] Vaclav Smil. Extreme designs:

From the pyramids to the Hummer, more is often less: Numbers don’t lie. IEEE Spectrum, 59(1):20–21, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:FFW

[Smi22f] Vaclav Smil. Fuel, food, and war: Energy and grain are part and parcel. IEEE Spectrum, 59 (9):20–21, September 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:HFR

[Smi22g] Vaclav Smil. Hydropower, the forgotten renewable: This stillmighty source of electricity has been unfairly relegated to the scrap heap. IEEE Spectrum, 59 (11):20–21, November 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:MEH

[Smi22h] Vaclav Smil. Meat-eating is as human as apple pie: We evolved to eat meat; the world will not turn vegan. IEEE Spectrum, 59 (8):18–19, August 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:NDLb

[Smi22i] Vaclav Smil. Numbers don’t lie: a Moore’s Law-for Bombs: the rising power of destructiveness is, unfortunately, the most impressive metric of modern technology. IEEE Spectrum, 59(7):18, July 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:NDLa

[Smi22j] Vaclav Smil. Numbers don’t lie: Electricity’s slow rollout: The spread of electricity has followed a protracted and characteristic pattern — slow, then fast, then slow again. IEEE Spectrum, 59(6):20–21, June 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smil:2022:WST

[Smi22k] Vaclav Smil. Waiting for superbatteries: They still can’t match the energy density of liquid fuel. IEEE Spectrum, 59(12):21, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2022:GLT

[Smi22l] Matthew S. Smith. Gizmo: a laptop that’s fit to be fixed: Dell’s concept responds to the right to repair movement. IEEE Spectrum, 59(3):22, March 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2022:SCC

[Smi22m] Matthew S. Smith. The spectacular collapse of cryptokitties. IEEE Spectrum, 59(9): 42–47, September 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2023:BGW

[Smi23a] Craig S. Smith. Baidu and Geely will mass-produce an autonomous EV: The Chinese tech giants aim for a fully self-driving car. IEEE Spectrum, 60(1): 36–37, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2023:MFS

[Smi23b] Matthew S. Smith. The miracle of flight simulator: It was a wonder that the original video game worked at all. Decades later, it still enthralls millions. IEEE Spectrum, 60 (6):36–41, June 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2023:GVD

[Smi23c]  Peter Rey Smith.  Guided

by voices: Digital voiceprinting may not be ready for the courts. IEEE Spectrum, 60 (7):28–33, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2024:CCN

[Smi24] Matthew S. Smith. Challengers are coming for Nvidia’s crown: In AI’s Game of Thrones, don’t count out the upstarts. IEEE Spectrum, 61(10):40–44, October 2024. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Someya:2021:PSD

[Som21] Takao Someya. Put down that smartphone: The display is on your skin: The screen is the last frontier in stretchable, bendable, body-conforming electronics. IEEE Spectrum, 58(6):38–43, June 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:GMMa

[SP21a] Matthew S. Smith and Mark Pesce. Gizmo: Macro & micro. IEEE Spectrum, 58(9): 24–25, September 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:GMMb

[SP21b] Matthew S. Smith and Mark Pesce. Gizmo: Macro & micro. IEEE Spectrum, 58(11): 24–25, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2021:GMMc

[SP21c] Matthew S. Smith and Mark Pesce. Gizmo: Macro & micro. IEEE Spectrum, 58(12): 20–21, December 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stenfatt:2021:FCA

[Ste21] Mads Stenfatt. The first crowdfunded astronaut. IEEE Spectrum, 58(12):22–29, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stockton:2020:CTB

[Sto20]  N. Stockton.  China takes

blockchain national: The statesponsored platform will launch in 100 cities. IEEE Spectrum, 57(4):11–12, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stone:2023:DRB

[Sto23] Peter Stone. Designing robots to best World Cup winners has inspired generations of roboticists. IEEE Spectrum, 60(7):40– 49, July 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2020:FTD

[Str20a] E. Strickland. Facebook takes on deepfakes. IEEE Spectrum, 57 (1):40–57, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2020:HCA

[Str20b] E. Strickland. Health care algorithms show racial bias — [news]. IEEE Spectrum, 57(1):6–7, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2021:TPU

[Str21] Eliza Strickland. The turbulent past and uncertain future of AI: Is there a way out of AI’s boomand-bust cycle? IEEE Spectrum, 58(10):26–31, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:ANA

[Str22a] Eliza Strickland. Andrew Ng, AI minimalist: The machinelearning pioneer says small is the new big. IEEE Spectrum, 59(4):22–50, April 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2022:BNS

[Str22b] Eliza Strickland. A bionic nose to smell the roses again: Covid survivors drive demand for a neuroprosthetic nose. IEEE Spectrum, 59(11):22–27, November 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2023:QLT

[Str23] Eliza Strickland. 5 questions for Luke Tan: Why his startup is producing green hydrogen for a distillery. IEEE Spectrum, 60 (11):21, November 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2024:QJB

[Str24a] Eliza Strickland. 5 questions for Joy Buolamwini: Why AI should move slow and fix things. IEEE Spectrum, 61(01):22, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2024:EYL

[Str24b] Eliza Strickland. This election year, look for content credentials: Media organizations combat deepfakes and disinformation with digital manifests. IEEE Spectrum, 61(01):24–27, January 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svrcek:2023:OGM

[Svr23] Matt Svrcek. The omnivorous generator: Mainspring’s linear generator can run on almost any fuel. IEEE Spectrum, 60 (3):34–46, March 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2020:CTU

[SZ20] E. Strickland and G. Zorpette.

COVID-19 has taught US that foresight and tech are a winning combination. IEEE Spectrum, 57 (10):20–23, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2023:AAM

[SZ23] Eliza Strickland and Glenn Zorpette. The AI apocalypse matrix. IEEE Spectrum, 60(8): 38–39, August 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2024:N

[SZCB24] Matthew Smith, Glenn Zorpette, Charles Q. Choi, and John Boyd. News. IEEE Spectrum, 61(2): 5–13, February 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Szkandera:2021:BYO

[Szk21] Filip Szkandera. Build your own RISC-V CPU: Even homebrew processors can use hot new tech. IEEE Spectrum, 58 (6):16–18, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tabrizian:2021:HAN

[TB21] Roozbeh Tabrizian and Swarup Bhunia. The hidden authenticators: Nanometer-scale electromechanical tags could thwart counterfeiters. IEEE Spectrum, 58(6):32–37, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teisberg:2023:SSI

[Tei23] Thomas Teisberg. Seeing secrets in the ice: a drone-based radar looks for the future of climate change. IEEE Spectrum, 60 (10):22–29, October 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Temiz:2020:LMV

[Tem20]  Y. Temiz.  The Lego mi-

croscope: a valuable lab tool began as a DIY project — [hands on]. IEEE Spectrum, 57(5):16–18, May 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thompson:2021:DLD

[TGLM21] Neil C. Thompson, Kristjan Greenewald, Keeheon Lee, and Gabriel F. Manso. Deep learning’s diminishing returns: The cost of improvement is becoming unsustainable. IEEE Spectrum, 58(10):50–55, October 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thomas:2021:PFT

[Tho21a] Dale Thomas. Phone-free texting: HamMessenger makes it easy to SMS over VHF. IEEE Spectrum, 58(12):14–16, December 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thomas:2021:TLT

[Tho21b] Edd Thomas. Turning letters into tones: a century ago, the optophone allowed blind people to hear the printed word. IEEE Spectrum, 58(7):34–39, July 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Towfiq:2021:ICS

[Tow21] Bayan Towfiq. The Internet’s coming sunny day: Subspace’s network provides a weather map and dedicated backbone for faster traffic. IEEE Spectrum, 58 (11):32–37, November 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2020:CCC

[Ulr20a] L. Ulrich. Curbside cab charging:

Wireless power tech keeps EVs on the go — [news]. IEEE Spectrum, 57(10):8–9, October 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2020:GBB

[Ulr20b] L. Ulrich. GM bets big on batteries: a new $ 2.3 billion plant cranks out Ultium cells to power a future line of electric vehicles. IEEE Spectrum, 57(12): 26–31, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2020:TTC

[Ulr20c] L. Ulrich. Top 10 tech cars: The scramble for electric dominance has begun. IEEE Spectrum, 57(4):30–39, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2021:TTA

[Ulr21] Lawrence Ulrich. The trend toward all-electric is accelerating. IEEE Spectrum, 58 (4):44–53, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2022:TIL

[Ulr22] Lawrence Ulrich. Tesla inside: The Land Rover Defender gets an electric makeover. IEEE Spectrum, 59(5):44–49, May 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Uskova:2021:RFR

[Usk21] Olga Uskova. On Russian farms, the robotic revolution has begun: Hundreds of aftermarket AIs are harvesting grain. IEEE Spectrum, 58(9):40–45, September 2021. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Vaidya:2023:CLA

[VA23] Jaideep Vaidya and Hafiz Asif.

A critical look at AI-generate software: Coding with the new AI tools is both irresistible and dangerous. IEEE Spectrum, 60 (7):34–39, July 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanSchoot:2023:MLM

[Van23] Jan Van Schoot. The Moore’s Law machine: The next trick to tinier transistors is highnumerical-aperture EUV lithography. IEEE Spectrum, 60(9): 44–48, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Verhoeven:2024:USR

[Ver24] Sjef Verhoeven. The ultimate SDR receiver: This open-source, low-cost radio has ultrahigh performance. IEEE Spectrum, 61 (12):18–20, December 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Versluis:2020:QCS

[VH20] R. Versluis and C. Hagen. Quantum computers scale up: Constructing a universal quantum computer with a large number of qubits will be hard but not impossible. IEEE Spectrum, 57 (4):24–29, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2024:STI

[WA24] Emily Waltz and Luigi Avantaggiato. Spent but not trashed: an Italian startup recovers valuable materials from old solar panels. IEEE Spectrum, 61(7):34–41, July 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2020:TTI

[WAC^{+}20] E. Waltz, E. Ackerman, C. Q. Choi, P. Dhar, and M. Dumiak. Take a tour inside a cell: Advances in microscopy let researchers give immersive VR trips through brain cells — [news]. IEEE Spectrum, 57(12): 8–14, December 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2020:ATB

[Wal20a] E. Waltz. AI takes its best shot: What AI can and can’t do in the race for a coronavirus vaccine — [vaccine]. IEEE Spectrum, 57 (10):24–67, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2020:BDJ

[Wal20b] E. Waltz. The boredom detector: Journal watch. IEEE Spectrum, 57(3):8, March 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2021:VGE

[Wal21] Emily Waltz. Vaccines go electric: a handheld gadget could usher in a new era of vaccines. IEEE Spectrum, 58 (6):24–31, June 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2022:AMO

[Wal22] Emily Waltz. The algorithm that mapped omicron: With antigenic maps, vaccines can evolve with COVID-19 variants. IEEE Spectrum, 59(4):26–31, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2024:QYL

[Wal24a] Emily Waltz. 5 questions for Yilu Liu: How to EMP-proof a building. IEEE Spectrum, 61(6):21, June 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2024:BIC

[Wal24b] Emily Waltz. The brain-implant company going for neuralink’s jugular: Synchron’s electrodes are delivered via blood vessel. IEEE Spectrum, 61(01): 42–43, January 2024. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watkins:2024:BDQ

[Wat24] Gavin Watkins. The best DIY quartz clock yet: Vastly superior timing tech comes home. IEEE Spectrum, 61(2):16–18, February 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wright:2021:APS

[WB21] J. Wright and C. Blake. Atomically precise sensors could detect a planet much like our own: Earth 2.0: This artist’s rendition shows how an Earth-like exoplanet might appear. IEEE Spectrum, 58(3):22–28, March 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wong:2023:CND

[WC23] Wendy H. Wong and Harry Campbell. The creepy new digital afterlife industry: These companies could bring you back without your consent. IEEE Spectrum, 60(11):38–43, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weaver:2020:NDE

[Wea20] C. Weaver. The new driver’s ED: Game developers teach Cruise’s autonomous vehicles to understand gestures made by people on the street. IEEE Spectrum, 57 (9):32–37, September 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watts:2023:LCE

[WPBS23] Michael R. Watts, Christopher Poulton, Matthew Byrd, and Greg Smolka. Lidar on a chip enters the fast lane: Sensors for self-driving cars and robots will be tiny, reliable, and affordable. IEEE Spectrum, 60 (9):38–43, September 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Young:2022:BHA

[You22] Britt H. Young. The bionichand arms race: High-tech hands are complicated, costly, and often impractical. IEEE Spectrum, 59 (10):24–30, October 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yan:2020:RHN

[YS20] W. Yan and D. Schneider.

The race for: a here-and-now COVID-19 test: Fast, high-tech, use-anywhere diagnostic tests are key to battling pandemics — [testing]. IEEE Spectrum, 57 (10):60–65, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zeloof:2020:HPT

[Zel20] S. Zeloof. Hack a Polaroid with a thermal printer: Go from film to digital — and add Wi-Fi — [hands on]. IEEE Spectrum, 57 (10):12–14, October 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zimmer:2021:AWS

[Zim21]  Katarina Zimmer.  Artificial

wombs are science fiction: But artificial placentas are on the horizon. IEEE Spectrum, 58 (4):22–29, April 2021. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2024:Na

[ZJL^{+}24] Glenn Zorpette, Debdeep Jena, Lucas Laursen, Michelle Hampson, Greg Uyeno, Dina Genkina, and Nikolai Kiselev. News. IEEE Spectrum, 61(3):5–12, March 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zahedi:2020:WGT

[ZL20a] S. M. Zahedi and B. C. Lee. A win for game theory in the data center. IEEE Spectrum, 57(4):40–44, April 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zhou:2020:DBS

[ZL20b] H. Zhou and T. W. Lee. Display that bend and stretch: Some smartphones can now fold like a wallet. In a few years, you may wear one on your skin. IEEE Spectrum, 57(11):24–29, November 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2020:RMD

[Zor20] G. Zorpette. RF MEMS deliver the “ideal switch”: After two decades of development, MEMS-based RF switches are finally finding real-world uses. IEEE Spectrum, 57(8):8–9, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2022:FTH

[Zor22a] Glenn Zorpette. The first transistor and how it worked: It emerged from a combustible mix of brilliance and ambition. IEEE Spectrum, 59(12):24–28, December 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2022:OSS

[Zor22b] Glenn Zorpette. One small step for radio: Next: One giant leap for Earth–Moon communications. IEEE Spectrum, 59(4): 4, April 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2022:REH

[Zor22c] Glenn Zorpette. The ripples from Edwin Hall’s effect: How a 143year-old insight is shaping spacecraft propulsion. IEEE Spectrum, 59(2):4, February 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2022:SSA

[Zor22d] Glenn Zorpette. The seabed solution: After 150 years, is the time finally right for deepocean mining? IEEE Spectrum, 59(1):7, January 2022. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2022:YES

[Zor22e] Glenn Zorpette. Your eye in the sky: Satellite reconnaissance comes in from the cold. IEEE Spectrum, 59(3):4, March 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2023:CKC

[Zor23a] Glenn Zorpette. Careers: Kyle Clark: From hockey enforcer to high-flying eVTOL CEO. IEEE Spectrum, 60(11):19–20, November 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2023:FEW

[Zor23b] Glenn Zorpette. Farewell, editorial wizard: Philip E. Ross retires after 17 years at IEEE Spectrum. IEEE Spectrum, 60 (12):2, December 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2023:FEY

[Zor23c] Glenn Zorpette. Finally, an eVTOL you can buy soonish: Opener’s BlackFly is the first of a radical new class of automated ultralight fliers. IEEE Spectrum, 60(1):32–57, January 2023. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2024:QRG

[Zor24a] Glenn Zorpette. 5 questions: Richard Garwin: The designer of the first hydrogen bomb on creating “the sausage”. IEEE Spectrum, 61(9):22–23, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2024:RDC

[Zor24b] Glenn Zorpette. Remembering Don Christiansen: The former Editor in Chief of IEEE Spectrum left an indelible mark. IEEE Spectrum, 61(12):2–4, December 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2024:Nb

[ZSG^{+}24] Glenn Zorpette, Eliza Strickland, Edd Gent, Gwendolyn Rak, and Katherine Bourzac. News. IEEE Spectrum, 61(8):6–14, August 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).